An Incomplete Bibliography of Publications in Historia Scientiarum (International Journal of the History of Science Society of Japan)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 December 2018 Version 1.03

Title word cross-reference

$124.00 [Hij90]. $29.00 [Bur93a]. 3[+]1/8 [Hol03]. $49.95 [Yaj07]. 7 × 7 [Yam10]. 9 [Mur94]. ^{e} [Gui86, Pen04, Yos81a]. ème [DP88]. ^{ieme} [Ano94b]. ^{th} [Pen04]. π [HKY89, Nak94a, Vol94]. π ≈ 3(1/8) [Mur92b]. ∑_{i=1}^{n} x_{i}^{2} = x^{2} [Ras94a]. × [Har87b].

0 [Hig01a, Izu05, Miu03, Mor04b, Sat05, She06a, She06b]. 0-19-860665-6 [Sat05]. 0-19-927016-3 [Izu05]. 0-520-24607-1 [Yaj07]. 0-691-11445-5 [She06b]. 0-8018-8235-4 [She06a]. 0-86078-668-4 [Hig01a]. 000 [Sas81b]. 00FF [Yos82]. 02/06/2000 [Has01].

1 [Kaw93a, Oka98, Yos98]. 10 [Høy03, Yos81c]. 10th [Suz81]. 11 [Hay94]. 12/02/1906 [Has01]. 1475/76 [Hig01b]. 1500 [Ito83]. 15073 [Mur05b].

16th [Maa91]. 17 [Sat86, Sat87]. 1700 [Nak83]. 17th [Maa91, Oh14, Yin13]. 18.5cm [Har87b]. 1843 [Ito16]. 1847 [Nak00a]. 1847/48 [Nak00a]. 1880s [Kim08a]. 18th [Ano94b, Kaw11, Kob02, Lor86, Nag80, Oh14, Nak98, THI17]. 18th-century [Kob02]. 19 [Nis92]. 190F [Yos98]. 1920s [Bro07, Kan13, Kim08b]. 1930 [Yaj07]. 1930s [Bro89, Kan13]. 1940 [Fur97]. 1940s [Mat98, YW05]. 1950s [HR15, Yam09]. 1955 [Nis92]. 1960s [FH12]. 1970s [Kim09b, LH12, NNSA09]. 1979 [Har82]. 1981 [Iya85]. 1984 [Nak85]. 1990 [Has91]. 1993 [Ano95a]. 1994 [Ano95c, Ano95d, Ano95e, Ano95b]. 1995 [Ano95i]. 1996 [Ano96]. 1997 [Ano97c]. 1998 [Ano98, Ano99]. 19th [Chu11a, Hor00, Hor02, SW15, Yin13, Gui86, Kim10, Kon10, Moo08, She01c, Uns84]. 19th-century [Hor00]. 1ère [Kaw93a].

2 [Kaw93b, Yos98]. 2-251-42005-3 [Yos98]. 2-251-42006-1 [Yos98]. 200 [Tom81]. 2006 [Nak07a]. 2007 [Has07a, Has08]. 2011 [Oka12]. 20th [Ito04, Kaw83, Miy08a, Sat06a, Bay18a, Bro83]. 20th-Century [Miy08a, Bay18a]. 215F [Yos98]. 21st [Ish06]. 23 [Ano85a]. 23.5cm [Har87b]. 295.00 [Jam90]. 2ème [Kaw93b].

3 [Izu05]. 3-528-06692-X [Nak00a]. 300 [Nag80]. 3rd [Che94].

4 [Ano95h, Has98a, Hig01a, Miu03, Mor04b, Oka98, Sas03, She06a]. 4-13-026210-6 [Sat06a]. 4-13-061351-0 [Miu03]. 4-13-061355-3 [Hor06]. 4-326-15366-0 [Mor04b]. 4-535-78243-1 [Oka98]. 4-7872-3236-3 [Kan06]. 4-938424-73-8 [Has98a]. 400 [Ito94, Yat83]. 42nd [Ano94i, Ano95i]. 43rd [Ano96]. 44th [Ano97c]. 45th [Ano99, Ano98]. 460 [Kaw84]. 47th [Ano00]. 48 [Nak00a]. 4800 [Suz82]. 49th [Ano02]. 4th [Ima12].

5 [She06b]. 50th [Ano03a]. 51st [Ano04b]. 52301 [Yus06]. 52nd [Ano05b]. 53rd [Ano06a]. 54th [Ano07]. 57th [Ano11a]. 5800 [Kon83a]. 5th [Ima12, Sza81].

6 [Sat05]. 670p [Har87b].

700 [Sas81a]. 70th [Ano94b]. 7537 [Mur91].

8 [Has98a]. 800 [Sas80, Tom81, Yos81b]. 83 [Tak84a]. 8393 [Yus09]. 84 [Mur84]. 8521 [Mur90]. 8528 [Mur90].

90-288-2284-6 [Hij90]. 978-81-70304-809-8 [Ike10]. 995 [Ike10]. 99F [Yos84a]. 9th [Qu96, Suz81].

A-Bomb [Mal09]. a-t-il [Har85]. Abaci [Miu81]. absorption [Kon00]. Academic [Hij90, Kik00, Nak80]. Accademia [Hak90]. Acceptance [Kaw81, Koj00, Sas85, Yam10]. access [KIF96]. according [Bie84, Mar81].

Accounting [Low89a]. Accounts [Kaw07]. Accumulated [Kik08]. Achaemenid [Toy82]. acides [Yos81a]. acids [Yos81a]. act [Ito16]. Actes [Ano94a]. Action [Yam83a]. Active [NM02]. Activities [Ano93c, Ano01b, Ano05d, Ich95, Yam95, Ano94a, Toc18]. Activity [Tak97]. adjustment [She01a]. Administrateurs [Ano94b]. Advance [Has80]. Advancement [Tuc14]. advisors [Nak18]. Aepinus [DP88]. AERI [Dim09]. Aerodynamic [Has07b]. Aeronautical [Has02b, Has17a, Nis03, Ano18, Has94]. Affirmation [Has92b]. after [Fuj16, Fum02, Has98b, KMT89, MS03, Sai16, Tak09]. Age [Bad10, Koj10, NKY15, Pye05, She06b]. Ages [Kaw88, Ras96, Sat85, Tak08]. Agricultural [Chu11a, De 03, Maa07, Moo12]. Agriculture [Bad10]. Aikitu [YS97]. Air [Has17a, Ano18]. Aircraft [Nis03]. Airplane [Has07b]. Akira [Sas81b, Set09, Nag80, Suz99]. al [Bur93a, Far95, Miu87, Hog91, Far95, Ras94a, Suz82]. al-Andalusī [Bur93a]. al-Dīn [Far95]. al-Hiyal [Miu87]. Al-Khwārizmī [Hog91]. al-Qanun [Suz82]. al-Tūsī [Far95]. al-Yāzdī [Ras94a]. Albany [Ano94b]. Albin [Ano94b]. Aldershot [Hig01a]. Alexander [Bie84, Bie84]. Alexandrian [Ima11]. Alfred [Tan90]. Algebra [Miu81, Pen04, Sas99a, Tia99]. algebraic [Ras01]. algèbre [Pen04]. algébriques [Hir95b]. algorithm [Luo15, Xu99]. Algorithmic [Che91]. Algorithms [Yus08]. Alienation [Wen16]. Aligning [Kus14]. Alleged [She08, Mur92b]. Allen [Yos81c]. allgemeinen [Kim82]. Almagest [Ano94a]. Alok [Bur93a]. along [Ito11]. also [Ano94a]. alumni [Kis06]. Amane [Kim10]. Amano [Tak92a, Tak00]. ambition [Kaw04]. Ambitions [PO05]. American [Bar15, Has98b, Hin89, Kar10, Nak18, Pye97, Sat01, Wan06]. ami [Loc00]. Ampère [Has83]. Amrita [Hay02a]. Amsterdam [Ano94b, Har82]. analyse [Ras99]. Analysis [Kik08, YO10, Luo15, Ras99, Tak98, Hay98b]. Analytic [Kno85, Sas85, Tak85]. Analytical [Yag81, Pan91, Pan92, Far95]. analytiques [Far95, Tak85, Pan91, Pan92]. Analytische [Nak00a]. anatomical [Ima15]. Ancient [HJ82, Lap08, Mor03, Siu95, Vol94, Gup91, Izu05, Oha95]. Andalusī [Bur93a]. André [Yad00]. Andrew [Ano00, Sat88]. angled [Hay94]. Angles [Tsu00]. Anglo [Bar15]. Anglo-American [Bar15]. animal [Mat83]. Anne [Kaw11]. Annibale [Ano94b]. Annick [iS95a]. annotated [Miu87]. annotating [Has17b]. annotée [Har87b]. Announcement [Ano91a, Ano92a, Ano97a, Ano09, Sas92]. Annual [Ano93b, Ano94i, Ano95i, Ano96, Ano97c, Ano98, Ano99, Ano00, Ano02, Ano03a, Ano04b, Ano05b, Ano06a, Ano07, Ano11a, Ano01a]. anti [Yam09]. anti-nuclear [Yam09]. Antiquité [Ano97a]. Antoine [Lam09]. any [Hay94]. Apercu [Mar85]. Apollonius [Sai86]. Apparatus [Nag11]. appearing [Sat98b]. Appendix [Ano94b, Yos01]. application [Rie12]. Applications [Miu86, Miu89]. Applied [Kik08]. apport [Has91]. Approach [Sat07, Tod88a, Yag81, Nav91, Tod88a]. Approximation [Hay91]. Arab [Has80]. arabe [Ras94b]. Arabic [Ras94b, Ses91]. Archaismes [Dho87].

Archaisms [Dho87]. Archimedean [Ras96]. Archimedes [Hay94, Sat81, Sat86, Sat87]. Archives [YMN97]. Area [Sat85, Mor04a]. argument [Kat98]. Aristotelian [Tan83]. Aristotle [Hig01a, Ima12]. Arithmetic [Ras94b]. Arithmetica [Mur80]. arithmetical [Sla95]. Arithmétiques [Ras94b]. Arm [Nak11b]. Armand [Har87b]. arms [HY15]. Arnold [Oza05]. arrangements [Yam14]. Arrhenius [HY83]. Art [Chu11b, She06a, Tsu81, Wat90a]. article [Mor03]. arts [Cro80]. Aryabhata [HKY89, Yan80]. Asahi [Kus81a, Suz82, Yos81b]. Asahi-Shuppan [Yos81b, Kus81a, Suz82]. Ashgate [Hig01a]. Asia [Sat06a, YY18, Ano92a, Bay18b, Has12a, Juž07, Kim08c, Kim10, Mak15, Nag83, Sat06b, Tak08]. Asian [Has10a, Sat95, Chi94, Mak15]. Aspect [Fum02, Ano94b]. Aspects [Boa06, Che91, Far95, Mak89, Mur87a, Sat89, Sat90, Sat92a, Sat92b, Sat96, Ano94a]. Assessment [Sat17]. Assistance [Sat01]. Association [Tuc14]. asthma [FH12]. astral [Mak15]. Astrolabe [Miy08b]. Astronomer [Juž09]. Astronomical [Has92a, Kus81a, Yos81b, Shi03, Sza81]. Astronomische [Sza81]. Astronomy [Has92a, Pye84, THI17, Yab80, Yab97, Ano94a]. Astrophysics [Has92a]. Atlanta [Ano94b]. Atom [LY89]. Atomic [Fuj82, Hin89, Ito94, Kis06, Koj10, Pye97, Sas09, Viz05, Yam89, Ich16b, Nak09, Tak09, Yam14, Dim09]. Atoms [Dim09]. Attempt [Iya85, Kon98]. attention [Tuc16]. Attitude [Fum02]. attitudes [Han15]. Aufenthalt [Oza05]. August [Har82]. Augustin [Juž10]. Austin [Bur93a]. author [Ano94b, Tod88a]. Authoritarian [Bot89, Moo12]. Authorities [Ich13a]. authorship [Hay95]. Autobiography [Sat05]. Autonomy [Wen16]. aux [Pen04]. avant [Hig01b]. azia [Sat06a].

B [Ano94b, Hij90, Tak93c]. B.V [Har82]. Baark [Sat88]. Babylonian [Yus09, Hol03, Mur88, Mur90, Mur91, Mur92c, Mur96, Mur97, Mur98, Mur99, Mur02, Mur03a, Mur03b, Mur05b, Yus04, Yus06]. Background [Hak90, Kaj03]. bacteriologist [IE18]. bacteriology [Obe04]. Badash [Yag10]. Bairstow [Has07b]. Bakteriologen [IE18]. bamboo [Luo15]. Banū [Ras96, Sat85]. Banarsidass [Hay02a]. Band [Fuj83a, Fuj83b, Nak00a]. Banu [Miu87]. Bao [Hor00]. Barrow [Sas85]. Based [Yam95, Hay02a]. Basel [Hir95a]. Basic [Ara95]. BC [Ima12]. BCE [Sza81]. BCW [iT84]. Beaulieu [Har87b]. Beckmann [Tom81, Kim82]. Beckmanns [Kim82]. Become [PO05]. Bedeutung [Kim82]. Beeckman [Hon96]. before [Ano18, Fum02, Has17a, Hig01b, Lia06, Nak09, YKK05, Fur97]. Beginn [Lor86]. Beginning [De 03, Ish06, Mot98, YS97, Hay04, Tak98, Yos81a]. beginnings [Ras99]. Begriffswandlung [Kim82]. Behind [Sat17, Kim09b]. Behring [IE18]. Behrings [IE18]. Beijing [Juž10]. being [Ima15]. Beitrag [Lor86]. belles [Yos98]. Bells [HJ82]. Beriberi [Obe04]. Berichtigungen [Sug86]. Berkeley [Yaj07, Sug13]. Berlin [Nak00a]. Bernoulli [She06a, Sla95]. Bertrand [She03a]. Bessel [She00, She01a]. Better [Kat00, Kim08b]. Between [Kan13, Bie85, Boa06, Duc07, Hay02b, Juž10, LXF99, Maa91, Mat04b, Nak94b, Ras97, iS95b, Sat12b, Shi08, Wen16, Yat83]. Beyond [HR15, Smi18, Wen16, Ich16a]. Beyträge [Tom81]. Beziehungen [Bie85]. Bézout [Dho85]. Bhāskarācārya [Hay02a]. Bhāskara [Hay91]. Bian [Qu96]. bibl [Hig01a]. bibliography [Hay96]. Bibliometrics [Juž15]. bicentenaire [Ano94a]. Big [Bro07, Fur97]. Bijutsukan [Yos01]. bin [Miu87]. Biography [Dar88, Oka12, Has17b]. Biological [Miz07, iT86]. Biologists [Miz98, Miz07]. Biology [Oba92, Sai96, Sai16]. biophysics [NKY15]. Biot [Has83]. Birkhäuser [Hir95a]. Birth [Kaw04, Nak83, Ano94b, Mur86]. bis [Bie84]. Bisekibungaku [Sas81a]. Blackwell [Hig01a]. Blay [Yos84a]. Blume [Fum89]. Board [Yam95]. Bockstaele [Ano94b]. Bodily [Kan06]. body [Ima12, Nak91]. Bohr [Kat98, Kon83b, Kon00]. Bomb [Hin89, Koj10, Mal09, Pye97, Sas09, Yam89, Kim09a, Nak09, Tak09]. Book [Bar15, Bur93b, Bur93a, Fuj88, Fum89, Fur97, Fur98, Got94, Har82, Har87b, Has92b, Has92a, Has98a, Has99, Has02a, Has07a, Has08, Has10a, Hig01a, Hij90, Hir95a, Hor06, Ike10, Ito83, Ito94, Iya85, Izu05, Jam90, Jid91, Jim98, Kan06, Kat04, Kaw84, Kaw86, Kaw88, Kaw82, Kno85, Koj10, Kon83a, Kur00, Kus81a, Mat89, Miu87, Miu03, Mor04b, Mur84, Nag80, Nak83, Nak00a, Nak85, Nak07a, Oka98, Oka12, Sas80, Sas81a, Sas81b, Sas03, Sas07, iS95a, Sat09, Sat88, Sat91, Sat93, Sat94, Sat95, Sat05, Sat06a, Set09, She06a, She06b, Suz82, Tak84a, Tak92b, Tod88a, Tod88b, Wäc92, Yaj07, Yam83b, Yat83, Yos81b, Yos82, Yos84a, Yos98, Yos01, Yos81c, Ano94b, Low11, Luo15, Mur92a, Sai85, Sai93]. Books [Ano94a, Ano94b, Ano03b, Ano04c, Ano05c, Ano06b, Ras94b]. Bordered [Ses91]. Born [Fre01]. borrowings [Lam09]. Bortkiewicz [She08]. Bos [Ano94b, Har82]. Boston [Hij90]. botany [Mét02]. bottom [Bay12]. boundaries [Ich16a, Jul99]. Bovelles [Miu88]. Bradwardine [Tak84b]. Bragg [Yam14]. Brahe [Has92a]. Brahmagupta [Kus81b]. Braunschweig [Nak00a]. Braunschweig/Wiesbaden [Nak00a]. Bray [Has08]. Brazil [Mot98, Mot01, Nag98]. Brazilian [Bot89, Mot01]. Breadth [Kon98]. Breite [Sza85]. Brésil [Nag98]. Brewing [Kam95]. Brian [Ano94b]. Bridge [Miu90]. Bridgman [Oka93, Oka03, Oka04, Tak98]. Brief [Oit82]. Bringing [Yaj98]. Britain [Hig01a, Has07b, Has10b, Has16, Ito16, Nis03]. Britain/Brookfield [Hig01a]. British [Tuc14, RS07, Tuc14]. broad [Ano94a]. Bronze [Lap08]. Brookfield [Hig01a]. brothers [Sat98b]. bubble [Yam14]. Buddhist [Mak15]. Buffon [Ano94a, Ano94a]. Building [Has07a, FH12]. Bun [Ano83b]. Bun-ichi [Ano83b]. Bunders [Fum89]. Bunka [Yat83]. Busard [Ano94b, Ano94b]. Butsuri [Has92b]. Buying [LY89].

C [Ano94b, Kon97, Pye84, Sug13]. C. [iT84, Yos81c]. c.1890 [RS07]. c.1950 [RS07]. C0036 [Kan06]. calcul [Bla99]. calculating [Nak94a]. Calculation [Bla99, HJ82, Mur90, Vol94, Has94, Sat98b]. calculations [Sug86]. Calculus [Sas81a]. Calendar [Sza88]. calendérique [Hor87]. calendrical [Hor87]. Calibration [Nha88]. California [Yaj07]. Call [Ano93a]. Callomon [Set09].

Cambridge [Yam83b, Sas81c]. canal [Mur92c]. candidates [SV99]. Canon [Ama82]. Capital [Yos83a]. Career [Bar94, Hig15, IE18]. Carl [Nak00a]. Carnot [Yos82, KIF96]. Carothers [Has99]. Cartesian [Jul99]. Cartesianism [Nak94b]. cartésiennes [Jul99]. cas [DP88]. Case [Ito04, Iya83, Kaw05, Lee06, Par03, Ste08, THI17, Yaj06, Yam82, DP88, Hor00, Hor02, Kaw11, Mea17, Nak13, Zha18]. cases [Shi04]. Casting [HJ82]. Casualties [Nak15]. Categories [Bur93a]. Catherine [Has10a, Has91]. Catholic [Ano94b]. Catoptrica [Bur93b]. Cauchy [Dho85]. causation [Bay18a]. causes [Obe04]. Celestial [Ait81]. Cell [Hay98a]. Center [Lim08, Hay94, Wac82]. Centers [Sat81]. central [Ima12, Mak15]. Centre [Har87b]. Centuries [Hig01a, Suz81, Ima12, Maa91, Pen04, Sas85]. Century [Ara95, Dai18, IL17, Ito04, Joh18, Kim10, Lim08, Low11, Mat02, Miy08a, Moo08, Nag80, Nak98, Sat06a, Sza81, Yin13, Ano94a, Ano94b, Che94, Chu11a, Hig01b, Hor00, Hor02, Kaw11, Kob02, Loh17, Lor86, Mat83, Nak13, Oh14, SW15, Sas99a, Sat12a, Sat13, Shi04, Bay18a, Bro83, Che91, DP88, Gui86, Hay99b, Ish06, Kaw83, Kaw93a, Kaw93b, Kim97, Kon10, Nak11b, Qu96, Sas81c, She01c, Sho90, THI17, Uch90, Uns84, Yam03, Yos81a, You18]. Ceos [Ima15]. cercle [Has91]. Certain [Mur87a]. certainty [Cro80]. certification [FH12]. Chūō [Iya85, Kaw84]. Chūōkōron [Ito83, Yat83]. Chūgoku [Kur00]. Châtelet [Kaw90, Kaw93a, Kaw93b, Kaw04, Kaw05, Kaw07, Kaw15]. Chaia [Ano94a]. Chalcedon [Ima11]. Chamarik [Sat95]. Change [Fum02, Sas92]. changes [Mat83]. Changing [Hin89]. chapître [Nag98]. Chapter [Nag98, Bie85]. Chapters [Ano94b, Che94]. characteristic [Nak91]. Characteristics [Lap08, NM02, Yab80]. Characters [HJ97, Che94]. Charan [Hay96]. Charles [Miu88, Yaj98]. Chemical [Fur07, He07, Kaj03, Kaj08, RS07, Tsu94, Uch90]. Chemist [Shi00]. Chemistry [Abe81, Bar15, Bro04, Bro07, Fur96, Fur97, Kaj03, Kik00, Kik08, Sak93, Shi00, Tsu00, Has99]. Chemists [Fuj88]. chez [Hig01b]. Chicago [Ano94a, Oga80]. Chidanken [Toc18]. Chikara [Ano04a, Got94, Hir95a, Miu03]. Chikatoshi [Mat89]. Chime [HJ82]. China [Ano94b, Has08, Lim08, Sat88, Uns84, Che07, HJ97, Hor00, Juž15, Kob02, Lap08, LXF99, Lia06, Lim08, Liu94, Mar85, Mét02, Oh14, SW15, iS95b, Siu95, Tia99, Uns84, Vol94, Yab97, Yaj06, Zha07, pZ04]. China. [Ano94a]. Chine [Mar85]. Chinese [Ano94b, Izu05, Kur00, Yat83, Yos81b, Che91, Che94, Han15, HJ97, He07, HJ82, IL17, Juž09, KY96, Mar85, Mor03, Nak80, SV99, Smi18, Wen16, Xia99, Zha18]. Ching [Xia99]. chinois [Has91]. Chinoises [Has91, Mar85, Jam90]. chitujyo [Sat06a]. Choe [Kim10]. Chofu [MV18]. Chokkan [Has92b]. Cholera [Joh18]. Chongzhi [Vol94]. Chosôn [Shi03]. Chosakushū [Got94]. Choson [Shi04]. Chr [Sza81]. Christiaan [Har82]. Christian [Lor86, Kaw81, Lor86, MV18, Uns84]. Christianities [GG01]. Christianity [Low08]. christliche [Uns84]. Christopher [Ano94b]. Chrysanthemum [Sas01]. Chugoku [Yos81b]. Church [Ano94b, Ano94b]. Chusei [Kaw88]. Cimatti [MV18]. CIMPA [Ano97a]. circles [Ait81].

Circulation [Moo08]. City [Ano94b, Luo15]. civil [Hor82]. civils [Hor82]. Clairaut [Nak02]. Clancey [Yaj07]. clarification [Hir84]. class [Chu11b]. Classical [Oka98, Sas07]. Classification [Sho90]. Claude [Jam90]. Claudine [Ano94a]. Claudio [Tak92b]. Clausius [KIF96, Yag81]. clocks [Has12b]. Co [Koj13]. Co-Producer [Koj13]. Cobo [Jim98]. Cohen [Ano94a]. cold [Oh14, Ich06, Les06, Wan06]. Collaboration [Low89b]. Collaborative [Low89b]. Collected [Got94, Kon83a, Tak84a]. Collection [Ano94a, Kus81a, Nag11, Yos81b, MV18]. Collège [Has91, Yos81c]. collision [Nak94b]. Colloque [Ano94a]. Colloquium [Iya99]. Colonial [Kaj08, Lee06, Low10, Miy08a, Nak05a, Sat06a, Hom05, Maa07]. Colonial/Imperial [Sat06a]. Colonialism [Bab05, Kar10]. Colonies [Par03]. Colonization [Hon06, Sat06b]. color [Tak13]. combustion [Yos84b]. Commemorate [Ano04a]. commencée [Har87b]. Comment [Har85]. commentary [Che94, Bur93b]. Commentatio [Kot01]. commented [Kon83a]. Comments [Gav86]. Committee [Ano93c, Ano97b, Ano01b, Ano05d]. communication [Toc18]. Community [Fum02, Fur96, Kim97]. Company [Fum89]. Comparative [Iya85, Kim10, Mét02, Sas99a, YW05]. compared [Pan91, Pan92]. Comparison [Lap08, MT82, Nak80, Oh14, Tak87b, Ano94a, Nak00b, SW15, iS95b]. Competes [Bla88]. compiled [Ano94a]. complementarity [Kat98]. Completing [Has92b, Mur03b]. Completion [Ano04a, Ito94]. complex [Tak85]. complexes [Tak85]. Composition [Uch90]. Compounded [Sai86]. Computation [YV90]. computers [Sug13]. Concept [Fur82, Hor95, Mat83, Nak94a]. Conception [Sug92]. concepts [Che94]. conceptualisation [Yos84a]. concernent [Hir95b]. concerning [GP91]. conditions [Got00b]. Conduction [Ano91i, YT91]. Conference [Ano92a]. Conflict [Ich13a]. Conic [Sai85]. Conjecturing [She06a]. Connection [Mat97, Mak15, Sat86, Sat87]. Connections [Bar15, Juž07]. Connexion [Sho90]. conoid [Hay94]. consequences [Bie85]. Conservation [Yam83a]. consideration [Hay02b]. Constance [Hig01a]. Constant [She07]. constitution [Dai18]. constitutional [Dai18]. Constructing [Yam03]. Construction [Has17a, Ses91, Ano18, Hay02b, Nak91]. Contact [Bar15, Kot06]. containing [Ano94b]. Contemporaries [Tsu00]. Contemporary [Mal09, Oit82, Bie84]. Contents [Ano83a, Ano83f, Ano83g, Ano84a, Ano84f, Ano84e, Ano85a, Ano85e, Ano85f, Ano86a, Ano86f, Ano86e, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano92b, Ano92c, Ano92d, Ano92e, Ano92j, Ano92k, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95f, Ano95g, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano05a, Ano13, Jam90]. Contest [Has02b]. Contested [Sat01]. Context [Boa06, Che07, Has92a, Kaj08, Kaw86, iS95b]. Continental [Tan90]. continuous [Kot06]. contradictory [Maa91]. Contribution [Kat00, Tak92b, Toy82, Lor86]. Contributions [KY96, Hay02b]. Contributors [Ano83c, Ano83d, Ano84b, Ano84c, Ano85b, Ano85c, Ano86c, Ano86b, Ano91j]. controllability [Oka03]. Controlled [Nis92]. Controversies [Fum02]. Controversy [Fre97, Ste08, Bie85]. Conversations [Hig01a]. Conversion [Kim82]. Cooperation [Ich13a]. Copernicantsm [Ano94b]. copy [KIF96]. Cornelis [Har87b]. corps [LH12]. Corrections [Sug86]. Correspondance [Har87b]. Correspondence [Juž10, Kon92]. Cosine [GP91]. cosmological [Ano94b]. cosmologie [Lam09]. cosmology [Izu97, Lam09]. Cotton [Ham11]. couleur [Yos84a]. Countries [De 03, Has80, Par03, Ano94a]. country [SR87]. Court [Ano94a, She06a]. coverage [Sat12a]. Coyne [Ano94b]. cracks [Has14b]. Critical [Bur93b, Oka12]. Criticism [Yan90, Mor16]. Cross [Sat88]. Cross-cultural [Sat88]. crossing [Ota17]. cruris} [Tak13]. crystals [Has14b, Yam14]. Cultural [Hak90, Has98a, Sat17, Yaj07, Yat83, Sat88]. Culture [Kan84, Mot01, Bur12, Joh18, Kaw93a, Kaw93b, Oki13b, Tuc14, Has02a, Izu05]. Cultures [Cla02, Has07a]. Curie [Kaw17]. Current [Wan06]. curriculum [SV99]. Curtis [Has92a]. curves [Ada99]. Cutting [Ish06]. Cutting-Edge [Ish06]. Cyclotron [Koj10, Smi02].

D [Fum89, Miu87]. Dīn [Far95]. d’Aepinus [DP88]. Daigaku [Hor06, Tod88b]. Dainamikusu [Nak07a]. Dairy [Suz81]. daishu [Tia99]. Daiyamondo [Tom81]. Daiyamondo-sha [Tom81]. Dajka [Hij90]. d’Alembert [Nak02, Nak00b]. damage [Oh14]. dans [Far95, Jul99, Kaw93a, Kaw93b, Nag98, Tak85]. Dantei [Has92b]. Darquier [Lam09]. Darwin [Kaw86, Kaw86, Yaj98]. Darwinism [Kaw86, Low08, Nag83]. Dauben [Hir95a]. David [Tod88a]. Davy [Yos81a]. Dawn [Tsu81, Ada95]. dealt [Sat86, Sat87]. death [Oga93]. Debate [Sat90, Sat92a, Sat92b, Sat96, Ima12]. debut [Yos81a]. débuts [Ras99]. Decide [Kat00]. Decision [Mal09, Yam89]. Deck [Nis03]. Deck-Landing [Nis03]. decline [Oki13a]. découvert [Har85]. d’écriture [Dho87]. d’Edo [iS95a]. define [Ano94a]. Definition [Yan90]. Definitive [Sas81b]. degree [Sat98b]. Dekaruto [Miu03]. Delhi [Hay02a]. demonstration [DP88, Tak13, DP88]. Demystifying [Tak98]. Dento [Nak07a]. depiction [Tuc16]. depuis [Ano95h, Iya95]. Derek [Yag84]. Derivation [GP91]. Derived [HKY89]. Descartes [Miu03, Cos85, Har85, Hee06]. Descending [Ito11]. Design [HJ82]. Deslandres [Fuj83a, Fuj83b]. destin [Ano94a]. Detecting [Kim11]. detonation [Tak09]. d’Euclide [Mur92a]. d’Euler [Dho87]. Deutsche [Nak00a]. Deutschland [Lor86]. developing [SR87]. Development [Abe81, Ara95, Ara89a, Ara89b, Bad05, De 03, Fuj83a, Fuj83b, Fuj91, Hay98a, Hin89, Iha95, Kim95, Kim03, Kon10, LY89, Mak89, MS03, Nag98, Nak84, Nak86, Nug88, Sat86, Sat87, Sat88, Sat94, Toy82, Yab80, Yam82, Got00a, Luo03, Nak18, Tak08]. développement [Nag98]. Devices [Miu87]. Dewey [Oka03]. Dhombres [Jam90]. diagrams [Yin13]. Dialectical [Fre01]. Dialectics [Kan13]. dialogical [Ota17]. Did [HY83, Ita82, Miy99, Har85, Hor95, Sat86, Sat87, SV99]. Diet [Smi18].

Diet-induced [Smi18]. Different [Che94, Kim97, Kon10, Tsu00, Shi08]. Differential [Sas81a, Kot06]. Diffusion [Mot98, Nag83, Hir84, Hom05, Nak97]. diffusions [Hir84]. Digression [Mat83]. dimensional [Tak98]. Dimensions [RS07, HY15]. Dingler [Oka03]. Diophante [Ras94b]. diophantien [Chr96]. diophantienne [Ras99]. diophantine [Chr96, Ras99]. Diophantus [Ras94b]. Dioptrics [Has89, Sas07]. diplomacy [Ich16b]. Diplomatic [Hak90, Kim09b]. Direction [Fum89]. Directions [Wan06, Bay18b]. disaffirmance [Nak97]. Disaster [Ogu06]. discord [Ich13b]. discourse [HR15]. discover [Har85]. Discovery [Fur98, Got00a, Hee06, Iya83, Kaj03, Kat04, Miy08b, She08, Yam83a, Yos00, YKK05, She99]. discrete [Pan91, Pan92]. discuss [Ano94a]. discusses [Ano94b]. Disease [Bay12, Bay18a, Hir18]. Disguise [Hee06]. dislocations [Yam14]. disorders [Sat12a]. disparu [Loc00]. Dispersion [Fuj82, Kon97, Kon00]. Dispute [Has10b, Ich13b, Kan13]. Disseminating [Hor00]. Dissertation [Kaw15, Kaw05]. dissimilarities [Mur80]. dissolution [Maa91]. distances [Lam09]. Diversity [Fur07]. Divination [IL17]. ’Division [Ama82, Ama82, Ich95]. DNA [NNSA09]. Do [Has10b, Siu95]. Doctor [PO05]. Documents [Yos01]. Dojin [Fuj88]. Dokumente [Nak00a]. Domesticating [You18]. Donald [Miu87]. Done [Sat07]. Dordrecht [Hij90, Miu87]. Dorofeeva [Has08]. Dorofeeva-Lichtmann [Has08]. Dotei [Has92b]. Double [Sat05]. downfall [Fuj16, Sai16]. Dr [Has01]. Dr. [Ano92i, Hig15, Loc00, iS94, Shi92, Suz99]. Dragon [HR15]. Drake [Oga89]. dream [Kaw93a, Kaw93b]. Drift [Tan90]. driving [Obe04]. Drude [Fuj82]. dual [Moo12]. Dudley [She06a]. Dulong [Yos81a]. d’un [Loc00]. d’une [DP88]. Dunhuang [IL17]. Duplicate [Sai93]. during [Ara95, Bro83, Bro07, Hir18, Ich06, Ich13b, Jam91, Mai06, Mur84, NRGY05, Yab97, Yam89, Yam95]. Duseo [Chu11b]. Dutch [Yat83, Shi00]. Dynamics [Nak07a]. dynasties [Yab97]. Dynasty [Ham11, Chu11b, Shi03, Luo15].

E.H.F. [iT86]. earlier [Pan91, Pan92]. Early [Ano94b, Boa06, Has07b, Hin89, Hor06, Mur98, Nak84, Sas81a, Toy82, Yam82, Got00a, Got00b, Has94, Ima11, Kat98, Loh17, Lor86, Maa91, Mea17, Nak18, Nav91, Sai16, Sat12b, Sza85, Yad00, Dai18, Ham11, Ito04, Jam91, Kaw83, Kim97, Kim08c, Mat02, Miy08a, Moo08, PO05, Sat85, Sho90, Uns84]. early-Meiji [Mea17]. Earth [Ano94b, Ano94b, Lim08, Yan80]. Earthquake [Kim11, Ogu06, Yaj07]. earthwork [Mur05b]. East [Ano92a, Has10a, Kim10, Sat06a, YY18, Yat83, Bay18b, Has12a, Jim98, Juž07, Kim08c, Mak15, Nag83, Pye84, Sat06b, Tak08, Yab80]. Eastern [Juž10, NM02]. École [Ano97a]. Economic [Ich95, Sat17]. Écrits [Har84a, Har84b]. ed [Ano94a, Ito94, Kaw86, Kaw88, Tak92b]. Edge [Ish06]. Edison [Has17b]. Edited [Fum89, Har82, Nak00a, Ano94b, Bur93a, Got94, Kon83a, Sas81b, Yos01]. Edith [She06a]. Edition [Bur93b, Sas81b]. Editions [Tak87b, Ano94a, Har87b]. Editor [Ano91k, Mur87b, Sas92]. Editors [iT92]. Edmund [Sug13]. Edo [NM02, Kam95, Nak94a]. eds [Ano94b, Has92a, Has02a, Has08, Has10a, Hig01a, Hij90, Hir95a, Kan06, Mor04b, Sat88, Sat95, Tod88a]. Education [Kaw84, Oka11, Sai16, Siu95]. Educational [De 03]. educator [Sug13]. Edward [Kus14]. effects [Tak09]. Egidi [Tak92b]. Égypte [Ano97a]. Eighteenth [Lim08, Low11, Nak13, DP88, Nak11b, Uch90]. Eighteenth-Century [Lim08, Nak13]. Einfluss [IE18]. Einstein [HY15, Abi00, Bla88, Got00b, Hom05, Ita82, Kan84, Mor16, Nav91, Tan87]. Electric [Iha95]. Electrical [Tak92b]. Electricity [Juž10]. Electrification [Dim09]. Electrodynamics [Ara87, Ara89a, Ara89b]. Electromagnetic [Has83, Nak84]. Electronics [Ish06]. electrons [Kon00]. Element [Kaj03, YKK05]. Elementary [Low89a, Mak89]. Elements [Dar88, He07, Nak98, Mur92a, Sai85, Sai93, Mur92a]. Elimination [Tsu94, Pen04, Pen04]. Elliptic [Ada99]. Emergence [Fur82, Has99, Hon06, Kaj03]. Emeritus [Ano92i]. Emil [IE18, IE18]. Émilie [Kaw07, Kaw15]. Emission [Kon92, Kon00]. empire [Low10]. emprunts [Lam09]. En-gi [Suz81]. Enactment [Mat98]. End [Bad10, Kon10]. endeavor [Mea17]. Endo [Sas81b]. Energy [Dim09, Got00a, Yam09, Yam83a]. Engineer [Kaw84, Miu90]. Engineering [Has07a, Kik08, Les06, Oka11]. engineers [Has94, Hor82]. English [Bur93b, Jam90, SW15]. enigmatic [Mur97]. Enlarged [Sas81b]. Enlightenment [Sas01, Kaw93a, Kaw93b]. eno [Ito94]. ensembles [Mur86]. Entangled [IL17]. Enterprise [Kar08]. enthalten [Sug86]. entitled [Fre01]. entre [Ras97]. Entropy [YO10]. Entwicklung [Wäc92]. environment [Bay18a]. Epidemiological [Hir18]. Épistémologie [Ras97]. Epistemology [Ras97]. époque [iS95a]. Equation [HY83, Dho87, Ras94a, Dho87, Ras94a]. equations [Che94, Hir95b]. equinoxes [Nak00b]. Era [Fuj88, Kaw84, Mat89, Iya95, Sug04, De 03, Kaw81]. eradication [Bay18a]. erased [Tak09]. Erasistratus [Ima15]. Eratosthenes [Ano94b]. ère [Ano95h, Iya95]. Erfindungen [Tom81]. Eric [Ano94b]. Erik [Sat88]. Erm [Mur05b]. Errata [Ano91i, Ano91h, Ano92f, Ano95h, Ano99]. Erratum [Ano18]. erroneous [Tak93b, Tak93a]. Errors [She07]. Esprit [DP88]. Essay [Miu86, Oga80, Pye97, Sat98a, Ano94b]. Essays [Sat09, Ano94a]. Essence [Ike10]. Essential [Oga80]. est [Mar85]. Establishment [He07, Kaw81, Oga96, Oka04, Oki13a]. État [Ano94b]. Ethical [Ano94a]. Études [Has91, Ano95h, Iya95]. Euclid [Bur93b, Kno85, Mur92a, Sai85, Sai86, Sai93]. Euclidean [Mar81, SW15]. euclidienne [Mar81]. Eugenic [Mat98, Yok18]. Euler [Dho87, Nak00b, Nak02]. Europe [Ano94a, CK90, Juž10, Kaw83]. European [Kim03]. evaluation [Yos00]. Evidence [Gav86, Kon83b]. Evolution [Ste08, Yaj98, Yam83b, Kaw13, Oga93, Iya95, Ano95h, Iya95]. Evolutionism [Mat02]. Examination [Has10b, Tsu00, Yan90]. examinations [SV99]. example [Hol03]. examples [GG01]. Exchange [Has98a, Yat83, Zha07]. exchanges [LXF99]. Expanding [Kim11]. Experience [Sat91, Nak13, Sat95]. expériences [Yos84b]. Experiment [Wit08]. Experimental [Ano91i, Juž10, Kat00, Kon10, Nak11b, Suz81, YT91]. Experiments [Has10b, Has10c, Kon81, Yos84b]. experimentum [Tak13]. Experts [Has12a]. explication [Maa91]. exponents [Mur97]. Expressions [Mur02, Mur03a, Mur91]. Extraction [Mur99, Hay99b].

F.R.S. [Juž07]. Fabricated [Iya83]. Factors [Low89a, Sat17]. Faint [Kon81]. Fall [Oga89, Tak93b, Tak93a]. fallout [HR15, Hig15]. familial [You18]. Family [Kaw17, YKK05]. Fangmatan [Luo15]. Fantoli [Ano94b]. FAPESP [Mot98, Nag98]. Far [Has10b, Jim98, Juž10]. Farmers [NM02]. Farming [NM02]. Favor [Bla88]. fear [Hig15]. Features [Ara95]. Feher [Hij90]. femme [Kaw93a, Kaw93b]. Fermat [Iya99, Ada99, Bla99, Kat99, Miy99, Ras99, Ras01, Sas99b]. Fermentation [Yam83a]. feu [Kaw15, Kaw05]. FF [Jam90]. fi [Suz82]. field [Hir18, Tak08]. Fieldwork [Kik08]. fifties [Fre01]. figs [Hig01a]. figure [Xia99]. Figures [Sat81]. Files [Yam95]. finding [Tak93b, Tak93a]. First [Hir84, Oka11, Yag81, Abi00, Ich16b, JN12, Mur94, Ras94b, Che91]. Firth [Miu90]. Fish [Hig15]. Five [Ano94b, pZ04]. Flat [Ano94b]. Flight [Has02b]. Flora [Yat83]. Floris [Ano94a]. fluctuation [Abi00]. Fluid [Nak11b, Nak13]. fluid-resistance [Nak13]. Focus [Par03]. Focused [YO10]. Focusing [Sat12a]. Folgen [Bie85]. Folkerts [Ano94b]. fonctionnelle [Dho87]. fonctions [Pan91, Pan92, Tak85]. foot [Has12b]. foot-ruler [Has12b]. Force [Kon92, Obe04]. Forces [Kaw90]. Foreign [Cla02, Wit08]. Forest [Chu11a]. form [Has14b]. Formal [Ger13]. Formation [Ara87, Ito94, Kik08, Mak89, Oka98, Sas81a, Yam83a, Yat83, Dai18, Hor82, Nak07a, Tak08]. formulae [Luo03]. Forth [Miu90]. Forty [Ano94a]. Forum [Sat92a, Sat92b, Sat96, Ano92f]. Foucquet [HJ97]. Found [Ōmo84]. foundation [Pan91, Pan92]. Foundations [Ano94a, Tak84b]. Four [Ano94b, Che94, Shi08]. Fourier [Ano91i, Hay99a, Tak93c, YT91]. fox [Bur12]. Fragestellung [Oit82]. Fragment [IL17]. Framework [Iha95]. francais [Iya99, Sas99b]. francaise [Yos82]. France [Ano94b, Has91, Hor82, Koj00, Mur86, Nak98]. Francesca [Has08]. Francois [HJ97]. Fraser [Yam83b]. Fray [Jim98]. Free [Oga89, Mor04a]. Freiberg [Wac82]. French [Ama82, Bla99, Chr96, Cos85, Dho85, Dho87, DP88, Eul87, Far95, Gui86, Har84a, Har84b, Har85, Har87a, Har92b, Har92a, Has91, Hig01b, Hir84, Hor82, Hor87, Iya95, Jul99, Kaw93a, Kaw93b, Lam09, Loc00, Mar81, Mar85, Mur86, Mur92a, Nag98, Nak98, Pen04, Ras94a, Ras94b, Ras97, Ras99, Sas99a, Sas99b, Tak85, Yos81a, Yos84b]. Friars [Jim98]. Friedr [Nak00a]. Friend [She06a, Loc00]. Frontier [Nak18]. frontières [Jul99]. Frühe [Uns84]. Frühgeschichte [Sza85]. fu [Nee80]. Fuel [Kim09b]. Fujikawa [MHM98]. Fukasaku [Sat94]. Fukui [Bar94]. Fukuoka [Bur93b]. Full [Has10b]. Fumihiko [Has02a, Nak11a]. function [Nak91]. functional [Dho87]. functions [Tak85]. fundamental [Lor86]. Further [Gav86].

Furukawa [Has99]. Fushimi [THI17]. Fusion [Nis92, UTN00]. Future [Wan06].

G [Ano94b]. G. [Izu05]. gaku [Kaw81]. Gakuin [Ano05b]. Galileo [Ano94b, Ano94b, Duc07, Ito11, Oga89, Tak93b, Tak93a]. Galois [Hir84]. Gang [Qu96]. Ganitapañcavimśī [Hay95]. Ganitasarakaumudi [Ike10]. Ganot [Tak87a]. Garden [Chu11a]. Gas [Kat00]. Gauss [RV03, She01a, Sug86]. Gauss’schen [Sug86]. Gayon [Ano94a]. Genbaku [Koj10]. Gender [Ito04, Kaw05, Kaw07, Chu11b, Kaw13]. General [Has92a, Kim82, Nak02, Tuc16]. genèse [Yos84b]. Genesis [Kon83b, Yam83b, Yus08, Ama00, Ama01, Yos84b]. Geneticist [Ste08]. Genetics [Sai13, Sai13]. Gensen [Yat83]. Genshi [Ito94]. Genshogaku [Kaw82]. Gentle [Bro89]. Geo [Yam03]. Geo-History [Yam03]. Geographie [Sza85]. geographische [Sza85]. geography [Sza85]. Geological [Fum02, MV18, Toc18, YY18, Zha18]. Geology [Yaj06, Zha18]. geomètres [Ano94b]. Geometric [She03a]. geometrical [Kot01, Yin13]. géometrie [Mar81, Har92b, Har92a, Kot06]. Geometry [Yus06, Yus08, Yus09, Har92b, Har92a, Hay98b, Kot06, Mar81, Ras01, SW15, Yus04, Sas07]. Geophysical [Kim08a]. George [Ano94b, Has08]. Geosciences [YY18]. German [Bie84, Bie85, Kim82, Lor86, Oit82, Oza05, Oza10, Sug86, Sza81, Sza85, Sza88, Uns84, Wac82, Wal05]. Germany [YW05, Kon10, She03b]. Gernet [Jam90]. Geschichte [Nak00a, Tom81]. gesellschaftlichen [Wäc92]. GHQ [Ich95]. GHQ/SCAP [Ich95]. gi [Kim10, Suz81]. Gift [Sat01]. Gijutsu [Nak07a, Has07a, Mat89]. Gijutu [Jid91]. gil [Hor02]. Ginza [Ito94]. Giorgi [Tak92b]. Giovanni [Tak92b]. Girisha [Sas80]. Globalization [Fur07]. Globe [Lim08]. Gnomon [Sza88]. God [Kan88]. Gokan [Kaw81]. Goonatilake [Sat95]. Got [Low10]. governance [Lee08]. government [Shi04, Wen16]. Grand [Has07a]. Graphical [Has94]. Graphics [Has08]. Gratama [Shi00]. Gravitation [Kan88, Kat04]. Gravity [Sat81, Hay94]. Great [Hig01a, MS95, Nis03, Ogu06]. Greece [Toy82]. Greek [Izu05, Sas80, Izu97]. Greeks [Sza81]. green [Moo12]. Greeting [Iya99]. Gregory [Yaj07]. Griechen [Sza81]. group [Mor03]. groupes [Hir95b]. groups [Kot06]. Growth [Ano91i, YT91]. Grundlagen [Kot06]. Grundprinzipien [Lor86]. Gu [Che07]. Guarded [iT86]. Guide [Tod88a, Tod88b]. Gullen [Ano94b]. Guns [Nee80]. Gupta [Hay96]. Gustav [Nak00a]. Gyakuenkinhō [Ito83]. gyeongjeji [Chu11a].

H [Ano94a, Ano94b, Har82, Hay02a]. Hajime [Tod88a]. Hakken [Fur98, Kat04]. Half [Ara95, Ich15]. Hall [Fur97]. Hallerstein [Juž09, Juž10]. Hamilton [Nak91]. Han [Ano94b, Kim10]. Han-gi [Kim10]. Hansen [Hir18]. Hartogs [Tak85]. Harvard [Oka04]. harvest [Mur96]. Hashimoto [Sat09, Yos81b]. Hasso [Nak85]. Hatsumei [Mea17]. haute [Kaw93a, Kaw93b]. Hautes [Has91]. Hayakawa [Ano92f, Nis92]. Hayashi [Kus81a, Mor04b, Sat91]. Health [iT84]. Heat [Ano91i, Yag81, Nak97, YT91]. Heaven [IL17, Ano94b]. Heibonsha [Mat89]. Heinrich [Lam09]. Helix [Sat05]. Helmholtz [Yam83a]. Helmut [Nak00a]. her [Kaw93a, Kaw93b, YMN97]. Herbal [Yat83]. Hermann [Fur82, Oza10]. Herophilus [Ima11]. Heuristics [Mau85]. Heurlinger [Fuj91]. Hidden [Sat06a]. Hideki [Ano92f, Yos81b, Nis92]. Hideo [Sat12b]. Higashi [Sat06a]. High [MS03, Kaw93a, Kaw93b]. Higher [Oka11]. Hikaku [Iya85]. Hikari [Kaw86]. Hilgendorf [Yaj98]. HILL [Miu87]. Hindering [Has80]. Hindu [Kus81a]. Hinokawa [Koj10]. Hippocratic [Ima11]. Hirase [Set09]. Hirayama [Sas81b, Suz99]. Hirono [Mor04b]. Hiroshi [Tod88b]. Hiroshima [Nak09]. Hirosige [Kon83a, Kon83a]. Hirota [Fuj88]. Hiroto [iI82]. Histoire [Ano97a, Nag98, Ras97, Mar85, Jam90]. Historia [Ano18, Høy03, Ano83f, Ano83g, Ano84f, Ano84e, Ano85e, Ano85f, Ano86f, Ano86e, Ano91b, Ano91c, Ano91d, Ano91h, Ano92b, Ano92c, Ano93d, Ano93e, Ano94c, Ano94d, Ano95f, Ano95g]. Historian [Low89b, Nak07b, Sak93, Hay96, Tak92a]. Historians [Fur96, Pye05, Sas96, Uem96, Liu94]. Historical [Ano04a, Ano05b, Has98b, Has10c, Ich06, Kaw86, Kim82, Kim95, Nag11, Oga80, Yam95, Tak13, Sat09]. Historicity [Fum02]. Historiographical [Ano94a]. historische [Kim82]. History [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92a, Ano92d, Ano92e, Ano92j, Ano92k, Ano93a, Ano93c, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano95i, Ano96, Ano97c, Ano97b, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano03a, Ano04b, Ano05b, Ano06a, Ano07, Ano09, Ano11a, Ano11b, Ats97, Bro83, Fur96, Got94, Has92a, Has98a, Hir95a, Hor06, HJ82, Ito83, Iya85, Jam94, KY96, Kim08c, Kim98, Kon83a, Kra82, Low89b, MHM98, Mau85, Mor04b, Mur84, Mur87a, Nak98, Oit82, Pye00, Ras97, Sai96, Sas81a, Sas81b, Sat06a, Suz12, Tod88a, Tod88b, Uem96, UTN00, Viz05, Wat90b, YO10, Yam03, YW05, Yat83, Yos01, Bay18b, Liu94, Mar85, Mor04a, Oma04, Rie12, Sza85, Tak93c, Tak00, Tan92]. history [YY18, pZ04, Ano93b, Ano01b, Ano05d, Oit82]. Hitherto [GP91]. Hito [Has07a]. Hitoshi [Suz82]. Hiyal [Miu87]. Hiyama [Hig15]. Hogendijk [Ano94b]. Hokkaido [Ano05b, Fur98, Nak18]. Hokusen [Has98a]. Hokusen-sha [Has98a]. Hole [Miz07]. Honour [Ano94b]. Honzōgaku [Yat83]. hope [NKY15]. Horiuchi [iS95a]. hospital [Shi04]. Hospitals [Uns84]. Hours [Hog91]. Hronszky [Hij90]. Hua [YR06]. Hua-Loo-Keng [YR06]. Huainanzi [Ano94b]. Hughes [Ano94b]. Hui [Che94, Hor95, Vol94]. Human [Has07a, Sat17, Ima12, Ima15]. Humanistic [Sak93]. Humboldt [Bie84]. Humean [Mor16]. Huygens [Har82, Duc07, Has89]. Hyakunen [Mat89]. Hydrogen [Kat00]. Hygiene [Sak98, Lee08]. Hygienic [Lee08]. Hyoden [Oka12]. Hyoron [Oka98, Sas81a]. Hyoron-sha [Oka98, Sas81a]. Hyoronsha [Got94]. hypergeometric [RV03].

I. [Hij90]. Ian [Has02a]. ibn [YV90, Suz82]. Ichi [Hor06, Ano83b].

Ichinokawa [Mor04b]. Idai [Shi81]. Idea [Hin89, Hon06, Nag83, Yam89]. Ideas [Che07, Kan88, Hir84, Kaw93a, Kaw93b]. idées [Hir84, Kaw93a, Kaw93b, Yos82]. ideological [Ich13b]. ideology [Kan13]. Igaku [Kur00]. Igarashi [Suz82]. ihr [Sza81]. ihre [Bie85, Wäc92]. II [Ano83a, Ano84a, Ano85a, Ano86a, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano18, Tom81, Yos98, Ama01, Ara89b, Fuj83b, Fum02, Har84b, Har92b, Has17a, Kim95, KMT89, MS03, Pan92, Sai85, Sat92a, Tak93b, Yam89, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano92j, Ano92k, Ano94e, Ano94f, Ano94g, Ano94h]. iii [Kon83a, Sat92b]. Iijima [Sat06a]. ik [JN12]. Ikari [Kus81a]. Illinois [Ano94a]. Illnesses [Smi18]. illustrated [Tuc16]. illustrations [Chu11a, Low11]. IM [Yus06, Sza81, Uns84]. image [Kim09a, Nak09]. Imaginary [Kim09a]. immutability [Tak13]. Impact [Low89a, Low11]. Imperial [Kaj08, Lee06, Sat06a, Yam82, Lee08, Mat97, Sas01, Tan06]. Imperialism [Kim03, Lee06, Pye84, Shi04, Pye05]. Imperialist [De 03]. Imperio [Pye00]. Imperium [Pye00]. Implications [Ich95]. implied [Hol03]. Importance [Kim82]. Important [Bro83]. Improvement [Has89, Eul87]. Imwon [Chu11a]. Inclined [Ito11]. included [Sug86]. Incongruence [Ich13a]. Independence [Sat95]. index [Hig01a, Yos01]. India [Sat88, Bab05, Gup91, Jad02, Mak15, Oha95, Par03, Yad00]. Indian [HKY89, Hay96, Hay99b]. indices [Sat88]. Indies [Pye84]. Indo [Kus81a]. Indonesia [Maa07]. induced [Smi18]. indulgence [Smi18]. Industrial [Oit82, Sat94]. Industrialisation [MS95]. Industriearchaologie [Oit82]. Industry [MS03, Yos83a]. Infini [Cos85, Bla99]. infiniment [Gui86]. infinitely [Gui86]. Infinitésimales [Sas03]. Infinitorum [Mur80]. Infinity [Bla99, Cos85, Hor95]. Influence [YR06, Hay98b, Hay99a, IE18, Shi03]. Influences [Kim97]. informal [Ger13]. Information [Ano83c, Ano83d, Ano84b, Ano84c, Ano85b, Ano85c, Ano86c, Ano86b, Ano91j]. Infrared [Fuj83a, Fuj83b]. ingènieurs [Hor82]. Ingenious [Miu87]. Inheritance [Hay98a, Høy03, Mur88, Mur01]. Initial [Oza09]. Inkster [Has02a]. Inoue [Yam95]. inquiry [Ano94a, Ano94a]. inscribed [Has12b]. Institut [Has91]. Institute [Ano04b, Dim09, Sai13]. Institutes [Mat89]. Institutions [Kaw93a, Kaw93b, Tuc14, Kaw04]. instrument [Sza81, Sza81]. Instrumentation [Kim11, Shi03]. Instruments [Has10c, Oka11]. integers [Luo03]. Integral [Sas81a]. Integrated [Sat06a]. Integration [Sat07]. intellectual [THI17]. interdependencies [Maa91]. Interest [Mur90]. interface [Sat12b]. Interference [Kon81]. internal [Sat86, Sat87]. International [Ano92a, KMT89, Sat93, Zha07, Ano94a]. Interplay [Boa06]. Interpretation [Oga89, Mur05b, Xia99, Yos84b, Yos84b]. Interpretations [Mur90]. Intersection [Hir95a]. intervalles [Ama82]. intervals [Ama82]. Intervention [Sai13]. Interwar [Has02b]. introduced [Tak87a]. Introducing [Hay98b]. Introduction [Bay18b, Bur93b, Fur07, Has10c, Has11, Has12a, Has14a, HM17, HY15, Ich06, Ich13a, Kik00, Low08, MS03, Nak05a, NY09, Sat06b, Suz12, Tak15, Yan08, Yos81b, Has16, Ich16a, Liu94, Mét02, Nak97, Obe04, She06a, Yos01]. Introductions [Kus81a]. Intuition [Has92b]. Inventarisation [Bro83]. Inventing [Has99]. inverse [Eul87, Tak85]. Investigations [Kim08a]. Invisible [Nak15]. Invited [Har82]. Isaac [Kan88]. ISBN [Has98a, Hig01a, Hij90, Hor06, Ike10, Izu05, Kan06, Miu03, Mor04b, Nak00a, Oka98, Sat05, Sat06a, She06a, She06b, Yaj07, Yos98]. Ishiwara [Oka12, Sig00]. Ishiyama [Tod88b]. Islam [Ano94a, Sas07]. Islamic [Shi03, Yab97]. isolation [iT86]. Issue [Has10c, Ich13a, Kaw05, YW05, Bay18b, Has16, HM17, Ich16a, Oki13b, Suz12, YY18]. Issues [Has98b, GG01]. Italian [MV18]. Italian-style [MV18]. itiro [Tak84a, Dar88]. Ito [Kaw88, Sas80, Suz82]. Itokawa [Sat12b]. iv [Kon83a, Sat96]. Iwanami [Kaw82, Kur00, Mur84, Oka12]. Iwanami-shoten [Kaw82]. ix [Hij90, Mur84, Oka98]. Iyanaga [Ano95h, Sas80].

J [Ano94a, Ano94b, Har82, Nak00a, iS95a, Yag84, Yam83b, Yos82, Yos84a]. J. [Kon97, Pye84, Tak93c]. Jûyoku [Kat04]. Jacob [She06a]. Jacobi [Nak00a]. Jacques [Ano94a, Jam90]. Jahrhundert [Sza81, Uns84, Wäc92]. Jahrhunderts [Lor86]. Jami [Has91, Has10a]. Jamison [Sat88]. Janarizumu [Oka12]. Janez [Juž07]. January [Yos01]. Japan [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano92j, Ano92k, Ano93c, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano95i, Ano96, Ano97c, Ano97b, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano03a, Ano04b, Ano05b, Has98a, Hor06, Kan06, Low89b, Mur84, Sat88, YW05, Yat83, Ano93c, Ano94b, Ano97b, Ano01b, Ano05d, Ano11a, Ano11b, ANS^{+}82, Ara95, Ats97, Bar94, Bay12, Bay18a, Bie84, Cla02, Dai18, Fur96, Has02b, Has12a, HM17, Has17b, Hay98a, Hig15, Hir18, Ich95, Ich06, Iha95, Ito04, Iya95, JN12, Joh18, Kaj03, Kam95, Kan84, Kar08, Kat00, Kaw83, Kik00, Kim08c, Kim08a, Kim11, Kim95, Kim98, Kim03, Kob02, Lap08, Lee08, Lia06, Loh17, LY89]. Japan [Low89b, Low08, Low11, Mak89, MV18, Mat98, Mat04a, MHM98, Mat02, Mea17, Mét02, MT82, Mih15, Mor04a, MS95, NNSA09, NRGY05, Nak94a, Nak97, NM02, Nak15, Nak09, Nis03, Nis92, Obe04, Ogu06, Oh14, Oma04, Oza05, Oza09, Oza10, PO05, Sai96, SW15, Sas96, Sas01, iS95b, Sat01, Sat12a, Sat13, SR87, Sat89, Sat90, Sat92a, Sat92b, Sat93, Sat94, Sat96, Shi00, Shi98, Suz81, Suz12, Tak87a, THI17, Tan06, Tsu94, Uem96, UTN00, Yaj98, Yam82, Yam95, Yam09, Yok18, YS97, You18, Has02a]. Japanese [Bar15, Fuj88, Iya85, Kus81a, Mat89, Mat97, Nak07a, Sas81b, Sat09, Sat91, Set09, Suz82, Yaj07, Yos81b, Ano93a, Fuj16, Fum02, HR15, Has12b, Hay04, Hon06, Juž09, Kar08, KY96, Kaw81, Kaw17, Kim97, KMT89, Low89a, Low89b, Low10, MS03, Miu90, Miz98, Miz07, Mot01, Mur87a, Nak86, Oga96, Sak98, Sas99a, Sat98b, Sat07, Sat12b, Tak97, Tan90, Toc18, Tom81, iT86, Wat90b, Yaj06]. Japon [Ano95h, Iya95]. japonaises [iS95a]. Jealousy [Ich13b]. Jean [Jam90, Ano94a, HJ97, Jam90]. Jean-Claude [Jam90]. Jean-Francois [HJ97]. Jesuit [Juž09, Juž15]. Jesuits [Has10a]. Jidai [Koj10]. Jiegenfang [Tia99]. Jingoki [Shi81]. Jiryoku [Kat04]. jitsuzo [Hor06]. Jo [Fur97]. Johann [Kim82, Kim82, Lam09, Tom81]. Johannes [Sat85]. John [Ano94b, Sug92]. Joint [Ano09]. Joji [Kik00]. Joseon [Chu11b]. Joseph [Hir95a, Has95, Nak94b]. Joske [Fum89]. Journal [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano92j, Ano92k, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, JN12, Hay04]. Journalism [Oka12]. Journals [Ano09, Zha18]. Juan [Jim98]. Juhasseiki [Nag80]. Jun [Oka12, Sig00]. June [Ano95i, Ano97c, Ano03a, Ano05b].

Kōron [Kaw84]. Kōronsha [Iya85]. Kōryū [Yat83]. Kōwa [Yat83]. Kagaku [Fuj88, Jid91, Oka12, Mor04b]. Kagaku-Gijutu-Seisaku [Jid91]. Kagakugijutsushi [Yos01]. Kagakusha [Fuj88, Nak85]. Kagakushi [Kon83a, Ito83, Tod88a, Tod88b, Ano83a, Ano86a, Ano92j, Ano92k, Ano94e, Ano94f, Ano94g, Ano94h]. Kagakusi [Ano84a, Ano85a, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano05a]. Kage [Kaw86]. Kai [Tod88b]. Kaichi [Miu90]. Kaihoshiki [Sat98b]. Kaiser [Mai06]. Kaiso [Tsu00]. Kaji [Fur98, Ich17]. Kaku [Koj10]. Kalender [Sza88]. Kamaishi [Wit08]. Kamatani [Mat89]. Kampo [Sug04]. Kanazawa [Ano02]. Kangnian [Hor00]. Kangxi [Han15]. Kankaku [Kan06]. Kankei [Yos01]. Kansei [Has92b]. Kanto [Ogu06]. Kapitel [Bie85]. kara [Has92b, Koj10, Oka98]. Karl [She06b, Yos83b]. Karriere [IE18]. Kashiwagi [Tod88a]. Katahiro [Hor87, Xu99]. Katen [Oka98]. Kawahara [Yos81b]. Kazuo [Høy03, iS94]. Kazutoshi [Kan06]. Kegan [Sat93]. Kei [Mat89]. Keiji [Kur00]. Keijo [Lee06]. Keisei [Ito94, Nak07a, Oka98, Sas81a]. Keiso [Ito94]. Keisoshobo [Mor04b]. Keizo [Yos81b]. Ken [Hor06]. Ken-Ichi [Hor06]. Keng [YR06]. Ken’ichi [Bar94, Bur93b]. Kenkyu [Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano92j, Ano92k, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Mat89, Tod88b, Ano83a, Ano84a, Ano85a, Ano86a, Ano05a]. Kenzo [Kaw82]. Kepler [HJ97]. Khaifeng [Nee80]. Khaifeng-fu [Nee80]. Khwārizmī [Hog91]. Kigen [Koj10, Kur00]. Kiji [Mea17]. Kikai [Kaw82]. Kikan [Mat89]. Kikuchi [Bar15]. Killing [Kot06]. Kimiyo [Has92b]. Kimoto [Wäc92]. Kimura [Yat83, Ste96]. kind [Kob02]. Kindai [Nak07a, Kan06]. Kindaika [Mat89]. kinetic [Nak97]. Kingendai [Mor04b]. Kingendai-Shi [Mor04b]. Kinsei [Hor06, Sas81a]. Kitab [Miu87]. Kitasato [IE18]. Kitasatos [IE18]. Kitashima [IE18, IE18]. Kitazawa [Kan06]. kitchen [Rie12]. Kiyoshi [Tak92a, Tak00]. Kiyosi [Has01, Yos81b]. Kizuita [Has07a]. kleiner [Oit82]. Kluwer [Hij90]. Knight [Tod88a]. Knowledge [Cla02, Has08, Hij90, Jam91, Nag11, Nak86, Pye00, Wit08, Bur12, JN12, Mat04b, Sat12a, You18]. Kobe [Ano03a]. Kobori [Nag80]. Kobunkan [Yos01]. Koizumi [iT84]. Kokoromi [Iya85]. Kokuritsu [Mat89]. Kondō [Got94]. Kontroversen [Bie85]. Korea [Chu11b, Chu11a, Has12a, Kim09a, Kim09b, Lap08, LH12, LXF99, Lim08, Miy08a, Moo08, Moo12, Oh14, Par03, Shi04]. Korean [Dim09, Hon06, Hor02, Miy08b, Sas09, Shi04, Shi08, Yin13]. Koseikaku [Sas81b]. Koseisha [Sas81b]. Koseisha-Koseikaku [Sas81b]. Koshiro [Har87a, Sas81a]. kouiki [Sat06a]. Koyanagi [Has92b]. Kozo [Fuj88]. Kramers [Kon83b]. Kratzer [Fuj83a, Fuj83b]. Krishnaji [Hay02a]. Kronecker [Bie85]. Kugo [Hor02]. Kuhn [Oga80, Nak07b]. Kumar [Bur93a]. Kume [Yos01, Yos01]. Kunio [Yam94]. Kunitake [Yos01]. Kuomintang [Wen16]. Kuriyama [Kan06]. Kushyar [YV90]. Kusukawa [Hig01a]. Kuwaki [Tan92]. Kyōritsu [Kaw86]. Kyoiku [Kaw84]. Kyoritsu [Kaw86, Kaw88, Nag80, Sas80]. Kyoritsu-shuppan [Nag80, Sas80]. Kyoto [Ano92i, Ano94i, Nag11, Smi02]. Kyusyu [Bur93b].

L [Ano94b, Suz82]. ’l-tibb [Suz82]. Līlāvatī [Hay02a]. Lab [Bar15]. Labban [YV90]. Labor [Sak98]. Laboratories [Bro04]. Laboratory [Miz07, Rie12]. labyrinth [Tak93b, Tak93a]. L’accueil [Yos82]. Lagrange [Oka98, Hir95b, Pan91, Pan92]. Lal [Hay02a]. Lambert [Lam09]. Landing [Nis03]. Lane [Yos81c]. largely [Hay02a]. LaSalle [Ano94a]. Last [Kat99, Mur84, pZ04]. Late [Moo08, Hor00, Shi04, THI17, Han15, Jam91, Low11, NM02]. Latin [Bur93b]. latitude [Sza85]. Latus [Kus14]. Launch [Sat01]. Lavoisier [Kaw11, Kaw13, Yos84b]. Law [Fur98, Tak93b, Tak93a, Hee06, Iya83, She08, Yam83a]. Lawrence [Yag10]. Laws [Kim10, Mat98, HJ97, Jad02]. League [Yok18]. learn [Siu95]. Learned [Tak97]. Learning [Yat83, Ras96]. least [She99]. l’École [Gui86]. led [Ama00, Ama01]. legendaren [Bie84]. legendary [Bie84]. légende [Yos82]. legitimacy [Nak13]. Leibniz [Ano94a, Hay98b, Hay02b, Xia99]. Leibnizian [Yam03]. Leonard [Has07b]. Leonardo [Miu81]. leprosy [Mor04a]. Lerner [Yos98]. lesson [Siu95]. Letter [Ano91k, Ano92f, Mur87b, Nis92, She06a, iT92]. lettres [Yos98]. L’Évolution [Ano94b]. Leydesdorff [Fum89]. l’histoire [Yos82]. l’Humanité [Ano94b]. Li [Kaw91]. Liba [Ano94a]. Liber [Miu81]. Librairie [iS95a, Yos82, Yos84a]. Lichtmann [Has08]. Lifa [HJ97]. Life [Kaw83, Mor04b, She06b, Yam83a, Yam94, Har82, Sat98a]. Light [Boa06, Fuj82, Kon81, Mal09, Sai85, Ano94a, Ano94b, Kon00]. Lights [Kaw86]. limin [Har87b]. Lincei [Hak90]. Line [Kon98]. link [HJ97]. Liouville [Hir84]. Lisse [Har82]. list [Ano94b]. Literary [Ito16]. Literati [Han15]. Liu [Che94, Hor95, Vol94]. Livre [Mur92a]. livres [Ras94b]. Lloyd [Izu05]. local [Bur12]. Locating [Lim08]. Loet [Fum89]. Logarithms [Miu89, Jad02, Mur97]. Logic [Hee06, Ito94, Maa91, Sug13]. London [Ano94a, Bar15, Fur97, Has10a, Hij90, Sat93, Sat94, Yos81c, Sat88]. Longer [Kim08b]. Loo [YR06]. Look [Hon06, KIF96]. Lost [Sat81]. lower [Chu11b]. lower-class [Chu11b]. Ltd [Sat88]. Lu [Jim98]. Lucky [HR15]. Luís [Has10a]. Lumbricus [Kus14]. Lumières [Kaw93a, Kaw93b]. lxvii [Nak00a]. Lysenko [Fuj16, Sai13, Sai16]. Lysenkoists [Fuj16].

M [Ano94b, Har82, Hij90, Mor04b, She06b, Tod88a]. M. [Ano94b]. Mādhava [GP91]. Mūsā [Ras96, Sat85]. Mach [Bla88, Ita82, Tak87b, Yan90]. Machinery [Kaw82]. Machines [Has10b]. MacMillan [Bar15, Sat88]. Macromolecular [Fur82]. Madame [Kaw90, Kaw93a, Kaw93b, Kaw04, Kaw05, Kaw13]. made [Has92b]. madeno [Has92b]. Magic [Ses91]. Magnetic [Kat00]. Magnetism [Kat04]. Magnitude [Nha88]. Magnitudes [Sat81]. Major [Ano94b, Liu94]. Make [Yam89]. Makes [Wit08]. Making [Cla02, Kar10, Lee08]. Makyo [Wat90a]. Malacology [Set09]. Malaria [Sat06a]. mammoth [Ano94a]. Mammouth [Ano94a]. Man [Sat05]. Management [Chu11a, Bay12, Hig15]. Managing [Sat12b]. Manifestations [GG01]. Manifesto [HY15]. Manila [Kar10]. Mansourah [Ano97a]. Manuscripts [Tsu00]. many [Sat86, Sat87]. Marāthī [Hay02a]. Mararia [Sat06a]. Marcantonio [Hig01b]. March [Ano95h]. Marie [Kaw11]. Marie-Anne [Kaw11]. Marin [Har87b]. Marine [Mat97, Miz07]. Markov [She06c]. Marseille [Sza85, Sza85]. Martzloff [Jam90]. Marx [Yos83a, Yos83b]. Marxism [Fre97]. Mary [Fur97]. Masakazu [Nak83, Nak85, Wat93]. Masanori [Fur98, Ich17]. Masao [Kaw86, Has98a, Kaw86, Mur05a]. Masayuki [Ito94]. Mass [Yan90, Loh17]. Massachusetts [Yam83b]. Masson [Jam90]. Material [Ish06, Nag11, Yat83, Nak97, Oki13b]. materialism [Fre01]. Materials [Ōmo84, Yos01]. mathematic [Oki13a]. Mathematica [Ano94b]. Mathematical [Dho85, Har84a, Har84b, Has91, Jam91, Kus81a, LXF99, Miu03, Mur03b, Tak84b, Yos81b, Ada95, Ama82, Ano94b, Che94, DP88, Hay95, Hay02b, Hor02, Høy03, Iya95, Mur91, Mur92b, Mur94, Mur98, Mur01, Mur05b, Nak94a, Sza85, Yin13, Bla99]. Mathematician [MT82, Qu96, Sas99b]. Mathematicians [NM02, Hay99b, Juž15]. mathematicien [Iya99, Sas99b]. Mathematics [Ano94b, Cos85, Got94, Hor06, Iya85, Kaw88, Mur87a, Mur88, Mur90, Nag80, Sas80, Sas81a, Sas81b, Sas85, Sas96, Siu95, Chi94, Far95, Ger13, GG01, Hay96, Hay02a, Hig01b, Hol03, Hor00, Jam94, Jul99, Kob02, Liu94, Maa91, Mar85, Mur84, Mur92c, Mur96, Mur99, Mur02, Mur03a, Mur03b, Nak13, Oga96, Oki13b, Sat98b, She03b, SV99, Hir95a, Ike10]. Mathematik [Nak00a]. Mathematiker [Nak00a]. Mathematiker-Vereinigung [Nak00a]. Mathématique [Cos85, Ama82, DP88, Far95]. Mathématiques [Jam90, Sas03, Ano97a, Ano95h, Bla99, Har84a, Har84b, Hig01b, Iya95, Jul99, Mar85, iS95a, Has91]. mathematischen [Sza85]. Mathematizing [Has83]. Mathesis [Hay02b]. Matter [Sug92, Tsu94]. Matters [Ōmo84]. Maurice [Sat05, Sat05]. Max [Fre01]. Maxime [Chr96]. May [Ano94i, Ano96, Ano97c, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano04b]. Meaning [Tak85, Chr96]. Means [Oit82]. Measured [She07]. Measurement [Kat00]. measurements [Sza81]. Mechanical [Yag81, Nak02, Nav91]. Mechanics [Ito94, Nak83, Oka93, Oka98, Oza09, Lor86, Mor16, Nak91, Pan91, Pan92]. Mechanik [Lor86, Nak00a]. Mechanism [Yos83b]. mechanisms [Ger13].

Mechanization [Nak84]. Media [Nak86, Duc07, Loh17]. Medica [Yat83]. medical [JN12, Mat04b, Mor04a, Mor03, Oh14]. medical-treatment [Mor04a]. Medicine [Ano09, Kur00, MHM98, Nak15, Sak98, Sat06a, Suz12, Bay18b, Dai18, Hir18, Ima11, Mih15, Obe04, Oh14, Shi04, Shi08, Smi18, Sug04, You18, pZ04]. medicines [Mat04b]. Medieval [Ano94b, Bur93b, Bur93a]. Méditerranée [Ano97a]. Meeting [Ano93b, Ano94i, Ano95i, Ano96, Ano97c, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano03a, Ano04b, Ano05b, Ano06a, Ano07, Ano11a]. megutte [Sas81a]. Mei [Jam94, Mar81]. meii [JN12]. Meiji [Fuj88, Kaw84, Mat89, PO05, Ano95h, Bur12, Cla02, De 03, Fuj88, Has17b, Hay98a, Iya95, Kaw84, Kik00, Kim11, Kim03, Low08, Mea17, Nak18, Sug04]. Memoir [Yag81]. Mémoires [Has91]. memorial [Nak11a]. Memoriam [Ano83b, Low14, Sza88]. memories [Loc00]. Memory [Kik96, Pye97, Yag10, Ano94a, Mur90]. Mendeleev [Fur98, Fur98]. Mendelism [Mat04a]. mental [Sat12a]. meridian [Mat04b]. Mersenne [Har87b]. Mesopotamia [Yus08]. Messungen [Sza81]. mesure [Ano94b]. Métailié [Has08]. metals [Yam14]. Meteorological [Miy08a]. Meteorology [Miy08a]. Method [Sat86, Sat87, Eul87, Hay94, Maa91, Nak94a]. méthode [Eul87]. Méthodes [Bla99, Has91]. Methodological [Fum02, Mak89, Uem96]. Methodology [HO89, Kon02]. Methods [Bla99, Fum02, YO10]. Metrology [Tak92b]. Michel [Ano94b, Yos84a, Yos98]. Michel-Pierre [Yos98]. Michi [Ito94]. Michio [Kus81a]. Micronesia [Sak98]. Microscope [Yam10]. Mid [Han15, Kim10]. Mid-19th [Kim10]. Mid-Qianlong [Han15]. Middle [Kaw88, Sat85, Ras96, Tak08]. Middleman [Has07b]. Mie [Tod88a]. Mikami [Sas81b]. Military [Nak15, iT86, Lee08]. Mills [ANS^{+}82]. Ming [Jam91, Yab97]. minime [Har87b]. mining [Wac82]. Ministry [iT84]. minute [Tak09]. Mirrors [Lap08]. Misconduct [Ish06]. Mises [She03b]. missing [HJ97, Loc00]. Mission [Uns84]. missionaries [MV18]. Missionshospitäler [Uns84]. Misuzu [Kon83a, Tak84a]. Misuzu-shobo [Kon83a]. Mitsubishi [Sat94]. Mitsukuni [Ano92i]. Mitsuo [Hir95a, Ito94]. Mitsutomo [Nak05b]. Mixed [Nak13, Oki13b, Oki13a]. Miyamoto [Jid91, Jid91]. Miyoshi [Kaw84]. Mobilization [Kol06, Tan06, Yam95]. Modal [Tan83]. Model [Fuj82, Ito94, Kon98, Kon00, Yam14]. Models [De 03, Has10b, Mak89]. Modern [Ano94b, Ano04a, Fuj91, Fur97, Has98a, Has02a, Hor06, Izu05, Kaj03, Kan06, Kar10, Kim08c, Mor04b, MS95, Nak07a, Sas81a, Sas96, Suz12, Tsu81, Chi94, Fre01, Hay02a, Mét02, Oga96, Ras99, Sat12a, Tuc14, Zha18]. modernes [Ras99]. Modernity [Bab05, Nak07a, Low10]. Modernization [Mat89, Ogu06, Shi08]. Mokei [Ito94]. Molecular [Fuj91, Has99, Oba92]. Molières [Nak94b]. Molland [Ano94b]. monde [Yos98]. Monjo [Yos01]. montanwissenschaftliches [Wac82]. Monuments [Shi98]. Moonlight [Ike10]. Morse [Yaj98]. mort [Ano94a]. Moscow [Kis06]. Motilal [Hay02a]. Motion [Ito11]. motomete [Hor06]. Motoo [Ste96]. Movement [Oha95, Hay98b, LH12, Sai16, Yok18]. movements [Yam09]. MR1845726 [Høy03]. Multiple [Iya83]. multiplication [Mur02]. Murakami [Ito83]. Murata [Iya85]. Muroi [Høy03]. Musa [Miu87]. Muscle [Yam83a]. Museum [Ano05b, MV18]. museums [Lig16]. musical [Ama82]. musicaux [Ama82]. Müssen [Oit82]. Myth [Oba92].

N [Hay02a]. Nīlakantha [Hay99b]. Nagaoka [CK90]. Nagasaki [Ito94, Sat94]. Nagoya [Has92b, Mak89]. Naimpally [Hay02a]. naissance [Mur86]. Nakamura [Sas81a, Har87a]. Nakaoka [Nak07a]. Nakayama [Sat93, Tod88b, Low14]. Nam [Hor02]. Naples [Miz98]. NASA [Has07a, Has07a, Sat17]. Nation [Bay12, Yaj07]. National [Har87b, Mat89, Sat07, Ano93c, Ano97b, Ano01b, Ano05d, Yok18]. Nationalism [Mat89]. Nations [Bur93a, Sat91, Sat95, Yos83a]. Natural [Ano94a, Moo08, Ogu06, Pye00, Shi98, Yat83, Hon96, MV18, Tak93b, Tak93a]. Nature [YMN97, Kaw15]. Naturwissenschaften [Oza10]. Navigation [Has17a, Ano18]. Navy [Mat97]. Nazi [She03b]. Necrologie [Har87a]. Needham [Has95]. Negotiating [Lim08]. nen [Mur84]. neonatal [Mih15]. nerve [Mat04b]. Netherlands [Sat13]. networks [Low10, THI17]. Neugebauer [Mur90]. Neutralist [Ste08]. Neutrality [Hon06]. New-Quantum [Kat00]. Newcomb [She02]. News [Ano83e, Ano84d, Ano85d, Ano86d, Ano91l, Ano92g, Ano92h, Ano94i, Ano95i, Ano96, Ano97c, Ano97b, Ano98, Ano00, Ano01a, Ano02, Ano03a, Ano04b, Ano04a, Ano05b, Ano05a, Ano06a, Ano07, Ano11a, Tuc16, Ano99]. newspaper [Sat12a]. Newton [Kaw86, Nak83, Oka98, Duc07, Has92a, Kan88, Kaw86, Lor86, Tak13, Yan90]. Newtonian [Nak98]. Newtonianism [Nak94b]. newtonienne [Yos84a]. Newtons [Lor86]. Nexus [Kar08]. Nihon [Got94, Hor06, Kan06, Oka98, Yat83, Iya85, Nak07a, Yat83]. nine [Che94]. Nineteenth [Low11, Shi04, Joh18, Sho90, Yos81a, You18]. nineteenth-century [Shi04, Sho90, You18]. Nippon [Bie84, Mat89, Sas81a, Sas81b, Bie84, Mur84]. Nipponium [Kaj03, Yos00, YKK05]. Nishi [Kim10]. Nishina [Kim08b]. Nishiwaki [NKY15]. Nisio [Ano92f, Kon83a, Oka12]. No [Ano83f, Ano83g, Ano84e, Ano85e, Ano85f, Ano86f, Ano86e, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91i, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92j, Ano92k, Ano93a, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95f, Ano95g, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Fuj88, Fur98, Hor06, Høy03, Ito83, Ito94, Iya85, Kan06, Kat04, Kaw84, Kaw86, Kaw88, Kaw82, Koj10, Kur00, Mat89, Miu03, Mor04b, Mur84, Mur92b, Mur94, Mur01, Nag80, Nak83, Nak85, Nak07a, Oka98, Oka12, Sas80, Sas81a, Sat06a, Tod88a, Yat83]. Nobuhiro [Kaw84]. Nobuo [Kaw17]. Noddacks [YKK05]. Non [Bro83, Has12b, Nak98]. Non-Newtonian [Nak98]. Non-standard [Has12b]. Non-Western [Bro83]. Norbert [YR06]. normalizing [Sai16]. Northeast [Yaj06]. Norwood [Miu89]. Nos [Ano84f]. Note [Hay91, Hir84, Yus09, Høy03, Shi03, Yus04, Hir84]. Notes [Kon92, Kus81a, Ras94b, Yos81b, She06a, Ras94b]. notice [Sat86, Sat87]. Notices [Ano03b, Ano04c, Ano05c, Ano06b]. notion [Hir95b, Tak85]. nouvelle [Yos84b]. Nouvelles [Har84a, Har84b]. Nozoe [Kaj08]. Nozomu [Sak93]. Nuclear [Bad05, Bad10, Kim09b, LY89, NRGY05, Nis92, UTN00, Wal05, Yam82, YW05, Yam09, HY15, Hig15, Ich16b, Kim09a, NKY15]. Number [The85, Kon00, Nak94a]. Numbers [Miu88, Sla95, She08]. numerical [Sug86]. nutrition [Rie12]. Nyūton [Nak83]. Nye [Fur97]. Nyumon [Tod88a, Tod88b].

o [Has07a, Hor06, Koj10, Sas81a]. O. [Mur90]. Obituary [Ano92i, Eza01, Has95, Has01, Kik96, Kim98, Loc00, Mot01, Mur05a, Nak05b, iS94, Shi92, Suz99, Wat93, Har87a, Loc00]. Object [Wat90a]. Objections [Yan80]. Objectivity [Hon06]. objects [Tuc16]. Obninsk [Ich16b]. Observation [Miy08a, Yus06]. Observations [Har84a, Har84b, Kim11, Yam10, Kus14, She06c]. Observatory [Ano94b, Pye84]. Observing [Has14b]. Obstacles [Has80]. obtain [Tuc16]. obtuse [Hay94]. obtuse-angled [Hay94]. occasion [Ano94b]. occidental [Has91]. Occupied [Ich95]. occurred [Ano94a]. Ocean [Ota17]. Ochanomizu [YMN97]. Official [SV99]. Ogawa [Kaj03, Yos00, YKK05]. Ohta [Ste08]. Ohya [Shi92]. Oka [Yam94]. Old [Kat00, Kon98, MT82, Cro80, Mur97, Mur98, Yus04, Yus06, Yus09]. Old- [Kat00]. once [Ano94b]. One [Tak09, Ano94b, Sat86]. only [Ano94a]. Ontology [Tan83]. Open [Ano94a]. Opening [Iya99]. Operation [Nak84, Oka93]. optical [Kon97]. oral [Rie12]. Order [Sat06a]. organ [Ima12]. Organization [Sho90]. Oriental [CK90, Mat04b, Pye05]. Origin [Fuj83a, Fuj83b, Hin89, Koj10, Yam89, Yat83, Oga93, Tak85]. original [Tod88a]. Originator [iT84]. Origine [Tak85]. Origins [Kur00, Oba92, HY15]. Osaka [Yam82]. Oscillator [Kon98, Kon00]. Oudemans [Pye84]. Outcome [Has91]. Outlook [Nug88]. ouvrages [Har92b, Har92a]. Ovales [Har85]. ovals [Har85]. Overseas [Kim08b, RS07]. Overview [Oit82, Mar85]. Oxford [Izu05, Sat05, Sas81c]. Oyodo [Jid91].

P [Ano94b, Har82, Har87b, IL17, Nak85, Ano95h, Jam90, Mur84, Yos82, Yos84a]. P. [Fuj82, Oka93]. Pacific [Ota17]. Padroado [Has10a]. Padua [Oga89]. pages [Miu87]. painter [Chu11b]. Palgrave [Bar15]. Palgrave-MacMillan [Bar15]. panopticon [Lig16]. paper [Fre01]. Papers [Ano93a, Kon83a, Ano94b, Har82]. Paradox [Sas01, She03a]. Parallel [Ger13]. Paris [Ano94a, Ano94b, Jam90, iS95a, Yos82, Yos84a, Yos98]. Parity [Gav86]. Part [Has92a, Hin89, Yam89, Kaw93a, Kaw93b, Sat86, Sat87]. participates [Hir18]. Participation [Kaw90]. Particle [KMT89, Low89a, Low89b, Mak89]. Particulate [Sug92]. partie [Kaw93a, Kaw93b]. partition [Yus04]. Pascal [Has92b, Har84a, Har84b]. pass [SV99]. Past [Ōmo84, Has12b]. Pasukaru [Has92b]. Path [Ito94]. patients [FH12, Mor04a]. patterns [Maa91]. Patwardhan [Hay02a]. Paul [Har87b, Sat93, Set09]. Pauling [Abe81]. ’Peaceful [LY89]. Pearson [She06b]. peculiar [Tak08]. pédagogie [Dho87]. pedagogy [Dho87]. Pennsylvania [Has99]. Penser [Hig01b]. people [LH12, Mar85]. perceive [Hor95]. Percy [Oka04]. Perfect [Miu88, Wit08]. perfectionnement [Eul87]. peril [NKY15]. Period [Sas81a, Han15, Nak18, Nak94a, Yus06, NM02, Yus09]. Periodic [Fur98]. Periphery [Cha05]. Persia [Toy82]. Personal [Nak07b, Yag84, Mea17]. Perspective [Ich06, Nak98, Sat88, Uem96]. Perspectives [Bar94, Ham11, Izu05]. perspectivism [Ito83]. petits [Gui86]. Phadke [Hay02a]. Phase [Low89b]. phénomènes [Yos84a]. Phenomenology [Kaw82, Oga80]. Pheru [Ike10]. Philadelphia [Has99]. Philosophical [Ano94a, Izu05]. philosophie [Hig01b]. Philosophique [iS95a, Yos82, Yos84a]. Philosophy [Hig01b, iI82, Ada95, Fre01, Hon96, Mor16, Hig01a]. Photochemical [Boa06]. Physical [Kik00, Kim10, Nag11, Sho90, Sug92, Yam89, YS97]. Physicist [Mot01]. Physicists [KMT89, Tak97, Ich13b, Wen16]. physico [Nak13]. physico-mathematics [Nak13]. Physicochemical [Fum02]. Physics [Ats97, CK90, Fur97, Has92b, HO89, Ish06, Ito04, Kaw93a, Kaw93b, Kim97, Low89a, Low89b, Nak98, Tak87a, Uem96, Yam82, Has14b, Hom05, Kan13, Oka04, Tak98, Fre01]. physiology [Ima15]. physique [Kaw93a, Kaw93b, Kaw04]. pi [Hol03]. Pictures [Ham11, Ham11]. pièces [Har87b]. Pierre [Iya99, Ras99, Sas99b, Yos98]. Pietro [Yos82]. Pilot [Nis03]. Pioneer [MHM98, Oka12, Oza10]. pioneering [Tak00]. Pionier [Oza10]. Pirogov [She01b]. Pisano [Miu81]. Pitiscus [Miu86]. Plan [Hin89, Hay02b]. Planck [Bla88, Kon02]. Plane [Sat81]. Planes [Ito11]. Planetary [Has92a]. plant [Ich16b]. Planude [Chr96]. plasmatikón [Chr96]. Plastic [MS03]. plates [Has12b]. plusieurs [Tak85]. Plutonium [Hin89]. Pneumatic [Boa06]. Pogson [Nha88]. Policies [Ich95]. Policy [Hak90, Jid91, Mat98, SR87, Mor04a, Yok18]. Political [Ich13a, Low89a, Sat17, Ano94b, Shi08]. politicians [Has12a]. Politics [Wan06, Yaj07, Fre01, LH12, Sat12b]. Polity [Koj13]. pollution [FH12]. pollution-related [FH12]. polygons [Sat98b]. Polymer [Has99, Zha07]. Polytechnical [Gui86]. Polytechnique [Gui86]. Populaire [Mar85]. Popular [Kaw07, Mea17, Joh18, Tuc16]. Popularization [Sat12a, Has16, HM17, Ito16, Lig16]. popularizer [Sug13]. Popularizing [Loh17]. Population [Sak98]. Porter [She06b]. Portrait [Hor06]. Position [Lim08]. possession [Bur12]. Possible [Yan80]. Post [Ara95, SR87]. Post-War [Ara95]. Postwar [Has98b, Mat98, Sat93, Ger13, Has12a, Yam09]. pour [Ano94a, Has91]. Power [Dim09, GP91, Iha95, LY89, Sat01, Ich16b]. powers [Luo03]. Pp [Hij90, Izu05, Bur93b, Bur93a, Fuj88, Has98a, Hig01a, Hor06, Ike10, Ito83, Ito94, Kan06, Kaw84, Kaw86, Kaw88, Kon83a, Kur00, Miu03, Mor04b, Nag80, Nak83, Nak00a, Oka98, Sas80, Sas81b, Sat88, Sat05, Sat06a, She06a, She06b, Suz82, Tod88a, Tod88b, Tom81, Yaj07, Yat83, Yos81b, Yos98, Yos01, Yos81c]. pp. [Har82, Has10a, Kaw82, Kus81a, Mat89, Oga80, Sas07, Yam83b]. PRA [Sat17]. Practical [Nak84]. Practice [Boa06, Kim08a, Wit08]. practiced [Mar85]. Practices [Kik08]. pratiquée [Mar85]. Pre [Sat94, Hir18, Oga96]. pre- [Hir18]. pre-modern [Oga96]. Pre-war [Sat94]. précession [Nak00b]. précis [Has91]. Predated [Yaj98]. Preface [The85, YW05, Yan91]. Prefaces [Jam90]. Preis [Bie85]. préliminaire [Hir95b]. Preliminary [Miu86]. Prelude [Mat98]. Premature [Kon98]. Première [Hir84]. premiers [Ras94b]. Prentice [Fur97]. Prescribed [Suz81, Hay99b]. prescription [Mor03]. Present [Ōmo84]. presentation [DP88, DP88]. Preservationists [Shi98]. President [Ano11b]. Press [Ano94a, Ano94b, Bur93b, Bur93a, Fur98, Has92b, Has99, Hor06, Izu05, Jid91, Miu03, Oga80, Sat88, Sat91, Sat05, Yaj07, Yam83b]. prevention [Bay18a]. prewar [Yok18, Bay12]. Price [Yos01, Yag84]. Principle [Ito11, Tan87, She99, Kon92]. principles [Lor86, Nak02]. print [Loh17]. prints [Low11]. Prior [Nha88]. Privat [Nak94b]. Prize [Bie85]. Probabilistic [Sat17]. probability [She01c, She03a]. Problem [Kaw07, Kim08c, Bur12, Mur94, Nak91, Ras94b, Tak85, Yus09]. problème [Ras94b, Tak85]. Problems [Mur88, Høy03, Mur92c, Mur96, Mur98, Mur01, Mur05b, Nak02, Sat86, Sat87]. procedures [Che94]. Proceedings [Low89b]. Process [Has92b, HJ82, Nak97, Oga96, Ota17, Sat98b]. Processes [Yam83a]. Producer [Koj13]. Product [Suz81]. Production [Bab05, Has08, You18]. Prof [Sza88]. Prof. [Ano04a, Mot01, Sza88]. professionals [Nak18]. Professor [Ano92i, Has01, Kik96, Toc18]. Program [Ano06a, Ano07, Ano11a, iT84]. Programs [Kon10, Sat12b]. Progress [Bot89]. Project [Has02b, Kim09b, Kis06, Viz05]. Projects [YW05]. Promoter [iT84]. Proof [Mau85, Hay94, Nak94a]. proofs [Hay94]. propagation [Kaw15]. properties [Sla95]. Proportion [Sas85]. Proportions [Iya83]. Propose [HY83]. Proposition [Sat86, Sat87, Hay94]. Protestant [Low08]. proved [Kat99]. Province [Juž09]. Provincial [Juž09]. pseudoconvexité [Tak85]. pseudoconvexity [Tak85]. Psychiatric [Nak15, Bur12, Sat12a]. psychiatrists [Bur12]. psychiatry [Oma04]. Ptolemy [Ano94a, Ano94a]. Public [Fum89, Sak98]. publications [Ano94b]. publiée [Har87b]. published [Yos01, Zha18]. Publisher [Yos01]. Publishers [Hay02a, Hij90]. Publishing [Fum89, Kaw15]. Pulte [Nak00a]. Pursuit [Fur97, Tak13, Has14b]. Pyong [Hor02]. Pyong-gil [Hor02]. Pytheas [Sza85, Sza85].

Qanun [Suz82]. Qi [Tsu94]. Qianlong [Han15]. Qin [Luo15]. Qing [Ham11, Jam91, Moo08, Tia99]. Qing-Dynasty [Ham11]. Quadrature [Sat85, Sat86, Sat87]. Quadrilaterals [Kus81b]. Qualities [Has07a]. quanta [Got00a]. Quantum [Abe81, Ara87, Ara89a, Ara89b, Fre97, Ito94, Kat00, Kon98, Nug88, Oka93, Oza09, Abi00, Ama00, Ama01, Got00a, Got00b, Koj00, Kon00, Mor16, Nav91, Sig00, Tak00]. quarrel [Gui86]. qu’elle [Mar85]. Quelques [Har92b, Har92a, Hir95b, Loc00, Mur92a]. Querelle [Kaw90, Gui86]. quest [Obe04]. question [Kaw13]. qui [Hir95b].

R [Har82, Izu05, Miu87, Nak91]. rêve [Kaw93a, Kaw93b]. race [HY15]. Radha [Hay96]. Radiation [Ich15, Kon10, Kon92, Loh17, Abi00, Ama00, Ama01, Kon02, NKY15, Tak09, Tak93c]. radioactive [Hig15]. radiological [HY15]. raft [Yam14]. Railway [Miu90]. Railways [Sat07]. Ramanujan [RV03, Sat98a]. Rangaku [Yat83]. rapides [Has91]. rapport [Has91]. Rashed [Sas03, Sas07]. Ratio [Sai86, Sai93]. Rational [Hay91]. rationale [Hay02a]. rays [Loh17]. Re [Yan90, Yos00]. Re-evaluation [Yos00]. Re-examination [Yan90]. Reaction [Sas85]. Reactions [Ogu06]. Really [HY83]. Reappraisal [Kim09b]. Rebuttal [Yan80]. Received [Ano94a, Ano94b]. Reception [Oza09, Tan90, Lor86, Mat04a]. Recherche [Har87b]. Recherches [Nak00b]. Recognition [Uch90]. Recollections [Nak07b]. recombinant [NNSA09]. reconciler [Nak94b]. Reconsidered [Tan87]. Reconsidering [Sai13]. Reconstructed [Tan83, Yam10]. Reconstruction [Sat86, Sat87, Hay94, Sat98b]. Recording [Kim11]. Records [Kam95, Oh14]. Redefining [Kik00]. Redondi [Yos82]. Reevaluation [iS95b]. Reexamination [Mur92b, Mur94, Mur05b]. Reexamining [Nis03]. Reference [De 03, Toc18]. Reflections [Izu05]. Réflexions [KIF96]. Reforestation [Moo12]. Refraction [Hee06]. regarding [Nak94a]. Regime [Bot89, Moo12]. Regional [Sat06a]. Regions [NM02]. regular [Sat98b]. regulation [NNSA09]. rehabilitation [Ano94b]. Reidel [Fum89]. Reject [Ita82]. Rekishi [Sas81a]. Rekishi-Bisekibungaku [Sas81a]. related [FH12]. relatif [Dho87]. relating [Dho87]. Relation [Ōmo84, Yam83a, Hom05]. Relationship [KMT89, Bie85, Hay02b, Wen16]. Relativity [Ita82, Nug88, Tan87]. relevant [Yos01]. Reliability [Has10b]. Reliance [Sat91]. religieux [Har87b]. Relocating [Bur12]. remarkable [Eul87]. Remarks [Kon81, Oga89, Mur92a, She00]. remarquable [Eul87]. remarques [Hir95b, Mur92a]. Reminiscences [Yag84]. Renaissance [Has92a, Ito83, Cro80]. René [Has92a]. Renormalization [Ara87, Ara89a, Ara89b]. repercussion [Mur86]. Report [Ano93c, Ano97b, Ano01b, Ano05d]. repr [Ito94]. Represent [Has10b]. Representation [IL17]. representations [Chu11b]. Reproduction [Suz81]. Republic [Mar85]. République [Mar85]. Research [Boa06, Dim09, Has02b, Has07b, HJ82, Ish06, Kaj03, Kaj08, Kim95, Mat89, NRGY05, Nak11b, Nis92, Oit82, Sas81c, Tod88b, iT84, iT86, UTN00, Wan06, Yam89, YKK05, pZ04, Ano94a, Hir18, Nak13, Tak93c]. Researches [Has98b, Kon10, Sug92, YS97]. residual [Tak09]. Resistance [Nak11b, Nak13]. resolution [Nak00b]. resolving [Nak02]. respect [Toc18].

Response [Ano09]. Result [Kim08b]. Resumption [KMT89]. retentissement [Mur86]. Revaluation [Ito83]. Reverse [Ito83]. Reverse-perspectivism [Ito83]. Review [Bar15, Bur93b, Bur93a, Fuj88, Fum89, Fur97, Fur98, Got94, Har82, Har87b, Has92b, Has92a, Has98a, Has99, Has02a, Has07a, Has08, Has10a, Hay02a, Hig01a, Hij90, Hir95a, Hor06, Ike10, Ito83, Ito94, Iya85, Izu05, Jam90, Jid91, Kan06, Kat04, Kaw84, Kaw86, Kaw88, Kaw82, Koj10, Kon83a, Kur00, Kus81a, Mat89, Miu87, Miu03, Mor04b, Mur84, Nag80, Nak83, Nak00a, Nak85, Nak07a, Oga80, Oka98, Oka12, Pye97, Sas80, Sas81a, Sas81b, Sas81c, Sas03, Sas07, Sat98a, iS95a, Sat09, Sat88, Sat91, Sat93, Sat94, Sat95, Sat05, Sat06a, Set09, She06a, She06b, Suz82, Tak84a, Tak92b, Tod88a, Tod88b, Wäc92, Yaj07, Yam83b, Yos81b, Yos82, Yos84a, Yos98, Yos01, Yos81c]. Reviews [Yat83]. Revised [Sas81b]. revisit [Hor95]. revisited [RV03]. Revival [Yam03]. Revolution [Ano94a, Has92a, Nug88, Yam03, Moo12]. Revolutionary [Abe81, Bro89]. Revolutions [Cha05, Moo12]. Rezeption [Lor86]. Richard [Fum89, Miu89]. Riemann [Kot01]. Righi [CK90]. rigor [DP88]. rigueur [DP88]. Rikigaku [Ito94, Nak83, Oka98]. Rise [Has92a, Low89b, Pye84, Obe04]. Risk [Sat17]. Ritsumeikan [Ano94i]. Ritter [Oza10, Oza10]. Road [Bro07, Mat89]. Robert [Yos83b]. Roberval [Har92b, Har92a, Hay98b]. Rodopi [Ano94b]. Roger [Ano94a]. Rokakuho [Tod88a]. Role [Abe81, Fuj91, Kis06, Nak11b, Set09, Izu97, Low10, Mor03, Nak91]. Romberg [Xu99]. Ronbun [Has92b]. Ronbunshū [Kon83a]. Ronri [Ito94]. root [Hay99b]. root-extraction [Hay99b]. Roots [Kar08, Mur99]. Rosenfeld [Ano94b]. Roshdi [Sas03, Sas07]. Rotation [Yan80]. Rough [Ats97]. Roundworm [Bay18a]. Routes [Has83]. Routledge [Sat94]. royal [Lig16]. Rs [Ike10]. Rudolf [KIF96]. Rudwick [Har82]. Rule [Sak98, Tak84b]. ruler [Has12b]. rules [Hay99b]. Russell [HY15]. Russia [YW05, Bro04, Kan13, Sai16]. Ryoji [Sza88]. Ryoshi [Ito94].

S [Ano94b, Har82, Kon83a, Oga80, Bur12]. S. [HY83]. Sā’id [Bur93a]. Sūgaku [Iya85]. Sūgakushi [Iya85, Got94]. Śaṅkara [Hay99b]. Sachiko [Hig01a]. Sadi [Yos82, KIF96]. Saemaul [LH12]. Safe [Has17a, Ano18]. Safety [Nis03]. Saguru [Koj10]. Saiensu [Nak83]. Saigusa [iI82]. Saikurotoron [Koj10]. Sakamoto [Kaw82]. Sakata [HO89, Mak89]. Sake} [Kam95]. Sakhya [Ike10]. Sakurai [Kik00]. Salem [Bur93a]. Saneh [Sat95]. Sankei [Mur80, Sat98b]. Sanskrit [GP91]. Santong [Kaw91]. Santong-Li [Kaw91]. Sapporo [Ano05b, Fur98]. Saraiva [Has10a]. Sasaki [Got94, Ano04a, Hir95a, Miu03]. Satio [Ano92f, Nis92]. Sato [Has07a, Sas80, Suz82, Hor06]. Satofuka [Has02a, Nak11a]. Saving [Che07]. Savior [Kim09a]. SBPC [Bot89]. Scale [Nha88, Has12b]. Scanning [Yam10]. SCAP [Ich95]. scholar [Chu11b]. Scholars [Jam91]. Scholarship [Mal09]. School [Oka11, Gui86]. schools [Sas99a]. Sci [Høy03]. Science [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92a, Ano92d, Ano92e, Ano92j, Ano92k, Ano93b, Ano93c, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Ano95i, Ano96, Ano97c, Ano97b, Ano98, Ano99, Ano00, Ano01a, Ano01b, Ano02, Ano03a, Ano04b, Ano05b, Ano05d, Ano06a, Ano07, Ano09, Ano11a, Ano11b, Bab05, Bar94, Bot89, Bro83, Che07, Cla02, Cro80, Fre01, Fur97, Has98b, Has99, Has10a, Has12a, Has80, Ich06, Ich13a, Ito83, Izu05, Jid91, Kar10, KY96, Kaw81, Kaw05, Kim08c, Koj13, Kol06, Kon83a, Lee06, Lia06, Low89a, Low08, MS95, Mot98, Nag98, Nak98, Nak05a, Oka11, Oka12, Pye05, Sas01, SR87, Sat89, Sat06b, Sho90, Tan06, Tod88a, Tod88b, Tsu81]. Science [Tuc14, Uem96, Wat90a, Wat90b, Yam95, Yos01, Zha07, Ano93a, Han15, Has16, HM17, Hor87, Ich16a, Ito16, Izu97, Kan13, Lig16, Low11, Mak15, Mat83, Mea17, Oza10, Rie12, Tak92a, Tan92, Toc18, Tuc16, Wac82, Bur93a, Has98a, Sat93, Tuc14, Ano04a]. Science-Making [Cla02]. Sciences [Fum89, Fur07, Maa07, Mor04b, Ras97, YY18, Fum89]. Scientiarum [Ano18, Ano83f, Ano83g, Ano84f, Ano84e, Ano85e, Ano85f, Ano86f, Ano86e, Ano91b, Ano91c, Ano91d, Ano91h, Ano92b, Ano92c, Ano93d, Ano93e, Ano94c, Ano94d, Ano95f, Ano95g]. Scientific [Cha05, Dar88, Has07b, Has10a, Has10c, Ich95, Juž07, Kar08, Nug88, Sas81c, Sat90, Sat92a, Sat92b, Sat96, She06b, Ano94a, IE18, Kaw93a, Kaw93b, Kaw11, Obe04, Tak92a, Tak08, Has92a, Hij90, Ich95, Ito16]. Scientification [LH12]. Scientifique [Har87b, Kaw93a, Kaw93b]. scientifiques [Kaw93a, Kaw93b]. Scientist [Low89b]. Scientists [Hon06, Kaw17, Nak85, Shi98, Wan06, HR15, Ich15, Ich16b, Toc18]. Scientization [Che07]. Scriba [Ano94b]. scrofulous [Dai18]. scrutiny [Tak98]. search [Cro80, Kon00]. Searching [Hor06, Koj10]. seasonal [Has12b]. Second [Low89b, Sas81b, Ich15, Nak00b, Ich06, Lia06, Mai06, NRGY05, Yam95]. Secrecy [iT86]. Section [Ich95]. Sections [Sai85]. see [Miy99]. Seen [NM02]. segment [Hay94]. Seimei [Mor04b]. Seimei-Kagaku [Mor04b]. Seimi [Tsu00]. Seiritsu [Ito94, Yat83]. Seisaku [Jid91]. Seismicity [Yaj07]. Seitaro [Toc18]. Seiyō [Iya85]. Seki [Hor06, Sat98b]. Selectionist [Ste08]. Self [NNSA09, Sat91]. Self-regulation [NNSA09]. Self-Reliance [Sat91]. selon [Mar81]. Sema’am [Bur93a]. Senbyoshitsu [Dai18]. Senkusha [Oka12]. sens [Chr96]. Sensation [Kan06]. sentence [Mur97]. Seo [Chu11a]. Separation [YKK05]. Sept [Ano92f, Nis92]. septembre [Ano97a]. Sericulture [NM02]. Series [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano92j, Ano92k, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, GP91, Hor06, RV03]. service [LH12]. ses [Har85, Kaw93a, Kaw93b]. Sesiano [Ano94b]. set [Hay99b, Mur86]. setback [Nak94a]. Sets [She06a]. setting [Hom05]. Seuil [Ano94a]. Seventeenth [Hig01a, Sas85, Ano94a, Mat83, Sas99a, Sat13, Lim08, Sas81c, Yam03]. seventeenth-century [Mat83, Sat13]. Seventh [Ano92a]. several [Tak85]. Sex [Sak98]. Shūkiritsu [Fur98]. Sha [Ito83, Kaw84, Kaw86, Kaw88, Nak83, Yat83, Has98a, Oka98, Sas81a, Tom81]. Shadows [Kaw86]. Shakir [Miu87]. Shankara [Hay02a]. shaped [Shi08]. Sharaf [Far95]. sheds [Ano94a, Ano94b]. Shi [Mor04b, Mur84, Hor00, Jim98]. Shibasaburô [IE18]. Shibasaburo [IE18]. Shiboruto [Yat83]. Shift [Tsu94]. Shigehisa [Kan06]. Shigeru [Low14, Sat93, Tod88b]. Shiki [Suz81]. Shima [Sat12b]. Shimodaira [iS94]. Shinichi [Shi92]. Shinkansen [Sat07]. Shinkaron [Kaw86]. Shinpei [Low10]. Shintai [Kan06]. Shipyard [Sat94]. Shiryo [Tod88a]. Shiso [Miu03]. Shizenshi [Yat83]. Shizue [Koj10]. Shizuma [Sza88]. sho [Kaw81, Kaw81]. Sho-sho-gaku [Kaw81]. Shobo [Ito94, Tak84a, Kon83a]. Shoichi [HO89, Jid91]. Shokichi [Ano95h, Sas80]. Shokubutsu [Yat83]. shokuminnchiigaku [Sat06a]. short [Oma04]. Shortly [MS03]. Shoten [Kur00, Mur84, Oka12, Kaw82]. Shu [Kus81a, Yos81b]. Shuntaro [Sas80, Suz82, Kaw88]. Shuppan [Kaw86, Kaw88, Tod88b, Yat83, Yos81b, Kus81a, Nag80, Sas80, Suz82]. Shuppankai [Hor06]. Shuppansha [Ito94]. Shushu [Tak08]. Shyam [Hay02a]. Si [Yam10]. Siberia [Lap08]. Siebold [Yat83, Ōmo84]. siècle [Ano97a, DP88, Gui86, Hig01b, Yos81a, Kaw93a, Kaw93b, Ano94b]. siècles [Pen04]. Sigeko [Ano92f, Oka12]. Significance [Kaw15, Sug92]. Signification [Tak85]. similarities [Mur80]. similarity [Mat04b]. Simon [She02]. Sin [Tak84a, Dar88]. Sin-itiro [Tak84a, Dar88]. Sina [Suz82]. Since [Kim95, Iya95, Liu94]. Sine [GP91, Hay91, Hog91]. Singh [Hay02a]. Sino [Hor02]. Sinology [Has91]. situs [Hay98b]. six [Ano94a]. Sixteenth [Hig01a, Hig01b, Hay99b, Sas85]. Sixteenth-Century [Hay99b]. Sized [Has10b]. Sketch [Ats97]. Skilled [Nak84]. Slater [Kon83b, Kon97, Kon98]. slips [Luo15]. Slovenian [Juž09]. Small [Mur92c, She08, Gui86]. smallpox [Shi04]. Snelders [Har82]. snow [Has14b]. Social [Fum89, Iha95, Kaj03, Low89a, Nag83, She01c, Suz12]. Socialized [Hij90]. société [Kaw93a, Kaw93b]. Societies [Ito16, RS07, Tak97]. Society [Ano93c, Ano94i, Ano95i, Ano96, Ano97c, Ano97b, Ano98, Ano99, Ano00, Ano01a, Ano01b, Ano02, Ano03a, Ano04b, Ano05b, Ano05d, Ano06a, Ano07, Ano11a, Ano11b, Bot89, Nak15, Sat93, Kaw93a, Kaw93b, Ano93b, Mai06]. Sociology [Fum89]. Sohn [Nak00a]. Solid [Ats97]. Solids [Sat81]. Solla [Yag84]. Somashekhara [Hay02a]. Somastuvan [Hay99b]. Some [Bro83, Has80, Kon81, Loc00, Oga89, Ōmo84, Sat89, Sat90, Sat92a, Sat92b, Sat96, GG01, She00, Har92b, Har92a, Mur92a]. Sommerfeld [Oza05]. Sommerfelds [Oza05]. son [Mur86]. Soné [Kat00]. sono [Ito94]. Sony [Ota17]. Sophia [Ano94b]. Sources [Tod88a, Ima15]. South [Dim09, Kim09b, Lap08, Moo12]. Southeast [Nag83]. souvenirs [Loc00]. Soviet [Bro07, Ger13, Ich13a, Ich13b, Ich15, Ich16b, Ich16a, Kan13, Kis06, Koj13, Kol06, Sai13, Viz05, Zha07]. Space [Sat01, Sat12b]. Spaceflight [Sat17]. Spanish [Jim98]. sparks [Has14b]. Special [Bay18b, The85, Has10c, Has16, HM17, Ich13a, Ich16a, Nug88, Oki13b, Suz12, YY18, YW05, Mor03, Toc18]. Specially [YO10]. specific [Obe04].

Specification [Tsu94]. Spectra [Fuj83a, Fuj83b]. Spectroscopy [Fuj91]. speculis [Bur93b]. Sphere [Gup91, Sat85]. spheres [Ait81, Yos98]. Spherical [Lim08]. spirit [Mat83, DP88]. spirits [Mat83]. Spontaneous [Kon92]. spread [SW15]. sprout [Nak94a]. square [Mur99, Mur03b]. Squares [Ses91, She99]. squaring [Mur91, Mur03a]. Śrīdhara [Hay95]. Srinivasa [Sat98a]. St [Ano00]. Stability [Has07b]. Stage [Nak86, Nug88]. Stages [Hin89]. Stance [Sat17]. Standard [Has02b, Has12b]. Standards [Has17a, Ano18]. State [Ano94b, Ats97, CK90, Kis06, Has12a, Ito16]. Statement [Ano11b, Xia99]. States [Bad05, Kra82, Nis03]. Station [Miz98]. Statistical [She06b, Abi00, Nav91]. statistical-mechanical [Nav91]. statistician [She01b, She02]. statistics [She01c]. Status [Hin89]. Staudinger [Has99, Fur82]. Staudt [Sla95]. Stay [Kim08b, Oza05]. Steam [Mat97]. Steiner [Bie85, Bie85]. Steiner-Preis [Bie85]. Stellar [Nha88]. Stellung [Wäc92]. Stenian [Yam03]. Stent [Oba92]. stereometry [Izu97]. Sterilization [Mat98]. Stone [Toy82]. Stone-Working [Toy82]. Story [Nag98, Ano94a]. Straban [Sza85]. Strabon [Sza85]. strange [Hig15]. Strategy [Yam10]. Strong [Has91]. structures [Ger13, Miy99]. Struggle [Kim09b, Tak93b, Tak93a]. Struggles [PO05]. Struggling [Bot89]. Stuart [Fum89]. student [Has14b]. Studies [Ano94b, Has10c, Moo08, Mur87a, Sai96, Sas81c, Sat89, Yam83a, Ama00, Ama01, Hir18, Ich15, Iya95, Tak00, Har82]. Study [Ich06, Kaw05, Kim08b, Kim10, Lee06, Ōmo84, Suz81, THI17, Yaj06, Yam95, Hor02, Kaw11, Kaw13, Mét02, Mor04a, Nak13, Sas99a, Tak08, YT91, Zha18, Ano91i]. style [Dho87, MV18]. Success [Kon10]. Successful [YKK05]. such [Ano94a]. Sugaku [Kus81a, Mur84, Sas81a, Yos81b, Kaw88, Miu03, Nag80, Sas80]. Sugaku-Shiso [Miu03]. sugakushi [Hor06]. Sugiura [Hir95a]. Suguakushi [Sas81b]. Suketoshi [Kik96]. Sulyo [Hor02]. Sumerian [Mur98]. sums [Luo03]. supported [Nak18]. Surface [Lim08, Sat85, Yam10]. Survey [ANS^{+}82, Kno85, Yin13]. Surveying [Sat13]. Survive [Bot89]. Susa [Høy03, Mur92b, Mur94, Mur01]. Susantha [Sat95]. Susceptibility [Kat00]. Sutherland [Hom05]. Sutras [Kus81b]. SVAYAMBHU [Sat98a]. Swets [Har82]. Sylla [She06a]. Syllogistic [Tan83]. symbolic [Sug13]. Symposium [Ano04a, Har82]. System [De 03, Has07a, Has17a, Miy08a, Oha95, Tan06, Ano18, FH12]. Systems [Sat07, Pan91, Pan92, Shi08].

T [Yos98, Har85, Har87b]. T. [Fuj91, Yam83b]. Tōkyō [Kaw86, Kaw86]. Tōkyō/Masao [Kaw86]. Tōzai [Yat83]. Tūsī [Far95]. Table [Ano83f, Ano83g, Ano84f, Ano84e, Ano85e, Ano85f, Ano86f, Ano86e, Ano92j, Ano92k, Ano93d, Ano93e, Ano05a, Ano13, Hog91, Jam90, Mur97]. tablet [Yus04, Yus06]. Tada [Set09]. Tadaaki [Wäc92]. Tadashiro [Yam95]. Taichi [IE18, IE18]. Taihoku [Kaj08]. Taikoku [Mat89]. Taisei [JN12, Sat98b]. Taishitsu [Dai18]. Taiwan [Kaj08]. takīltum} [Mur03b]. Takabayashi [Eza01, Loc00]. Takahashi [Bur93b]. Takakazu [Hor06]. Takao [Kus81a].

Také [Kat00]. Takebe [Hor87, Mur80, Sat98b, Xu99]. Takehiko [Loc00, Eza01, Sat09]. Takenosuke [Jid91]. Takeshi [Sat91]. Taketani [Ito94, Mot01]. Tamagawa [Nak85]. Tamamushi [Ano83b]. Tamotsu [Iya85]. Tanakadate [YS97]. tangentes [Eul87]. tangents [Eul87]. Tanja [Nak83]. Tannery [Har87b]. Taton [Has92a]. Tatsuo [Suz82]. Taub [Ano94a]. taught [Ano94b]. Taysir [YV90]. Teaching [Kik08, Wat90b]. Technical [Ara95, Has08, Ich95, Kaw84, Oit82, Sat17, YO10, LH12]. Technikgeschichte [Oit82]. technische [Oit82]. Techno [Mat89]. Techno-Nationalism [Mat89]. Technological [Ich95, Juž07, Kim95, Kim03, Lap08, MS03, Sat88, Sat95]. Technologie [Kim82, Nag98, Wäc92, Yos82]. Technology [Ano09, Bar94, De 03, Has98b, Has02a, Has80, Ich06, Iha95, iI82, Jid91, Kar10, Kim09b, Kim82, Kim98, Kra82, Lia06, MS03, Nag98, Nak07a, Nis03, Oit82, Ota17, Sat01, Sat13, SR87, Sat91, Sat93, Tan06, Yam95, Yos01, Has12a, HM17, Low11, NNSA09, Sat12b, Ano04b, Sat94, Yam95, Sat09]. Technoscience [LH12]. teikoku [Sat06a]. Telegraphs [Nak84]. Telescope [Has89]. telle [Mar85]. Temperance [Yok18]. temps [Bla99]. Tengenjutsu [iS95b]. Tenmongaku [Kus81a, Yos81b]. Tennis [She06a]. Tension [Oga80, Oga80]. tentative [Tak08]. Tenth [IL17, Kno85]. Tenth-Century [IL17]. Terada [Has14b]. Term [YO10, Chr96, Mur03b]. terme [Chr96]. terminology [Kob02]. Terms [He07, Hay02a]. Terrains [Kim08a]. Tetsujutsu [Mur80]. Tetsuo [Kaj08]. Tetsuro [Nak07a]. Tetu [Kon83a, Kon83a]. Texas [Bur93a]. Text [Has08, IL17, Yus09, Dho87, Høy03, Mur91, Mur92b, Mur94, Mur01, Mur05b]. Textbooks [Dho85, Tak87a]. texte [Dho87]. texts [Hor02, Kaw11, Yin13, Yos01]. Textual [Tak87b]. Thakkura [Ike10]. Their [Fur96, YKK05, Mur80, Toc18]. Theodore [She06b]. Theorem [Kat99]. Theoretical [Che91, Hom05, Ima15, Oka04, Tak98]. théorie [Ama82, Hir84, Hir95b, Mur86, Tak85, Yos81a, Pan91, Pan92]. theories [Abi00]. Theory [Ano91i, Ano04a, Ara87, Ara89a, Ara89b, Boa06, Fuj82, Has83, Hay98a, Ita82, Kat00, Kaw83, Kon83b, Kon98, Low89b, Mak89, Nug88, Oha95, Sai85, Sas85, She07, Tan90, Tsu94, Yag81, Yaj98, Yan80, Yos83b, Abi00, Ama82, Ama00, Ama01, Ano94b, Got00a, Got00b, Hir84, Koj00, Kon02, Kon97, Kon00, Mur86, Nak97, Nak94b, Nav91, Oga93, Sat98b, She01c, Sig00, Tak85, Tak00, Yos81a, YT91]. Thermal [Ama00, Ama01, Kon10, Abi00, Tak93c]. thermodynamics [Abi00]. these [Miy99]. Thing [Nag11]. things [MV18]. thinkers [Ano94a]. Thinking [MT82, Nak85, Hig01b]. Third [Che91, Sat05]. thirties [Yad00]. Thomas [Nak07b, Oga80]. Thomson [Hay99a]. Thought [Ito83, Miu03, Sat86, Sat87, Ano94b, Kot01, Mat04b]. Threat [Ano09]. Three [Ano94b, Nak86, Tsu00, Nak91, Ras94b]. three-body [Nak91]. Tianshui [Luo15]. tianyuan [Tia99]. Tianyuanshu [iS95b]. tibb [Suz82]. Tilling [Ham11]. Timaean [Izu97]. Time [Bla99, Has12b, Yam83b]. Tinemue [Sat85]. TMS [Yus04]. together [She06a]. Tohae [Hor02]. Tohru [Sas80]. toilets [Bay12]. Tokai [Jid91]. Tokugawa [Kaw81]. Tokyo [Ano04b, Fuj88, Got94, Has98a, Hor06, Ito83, Ito94, Jid91, Kaw84, Kaw82, Kon83a, Kur00, Kus81a, MV18, Mat89, Miu03, Mur84, Nag80, Nak83, Oka98, Oka12, Sas80, Sas81a, Sas81b, Sat91, Sat95, Suz82, Tak84a, Tod88b, Yat83, Yos81b, Yos01, Ano01a, Fuj88, Kaw88, Kik08, Tod88a, Tod88b, YMN97]. Tome [Har87b]. Tomoko [Ste08]. Tomonaga [Dar88, Tak84a]. tones [Luo15]. topic [Pan91, Pan92]. Topography [Pye84]. Torahiko [Has14b]. Toshiko [Ito04, Kaw17, YMN97]. Toshio [Kim98]. Toshisada [Sas81b]. Tradition [Che91, Has91, Nak07a, Sat90, Sat92a, Sat92b, Sat96, Hay02a, Ima11]. Traditional [MS95, Sug04, Chi94, Sat98b, Shi08, SV99, Wac82]. Traditions [Bur93b, Nak80]. traditionsreiches [Wac82]. training [Hor82]. trajectory [Duc07]. Transfer [De 03, Kim09b, Mat97, Sat91, Ota17, Sat13]. Transformation [Has92a, Ich95, Tak13]. transformations [Kot06]. transistor [Ota17]. Transition [Han15, YY18]. Translated [Miu87, She06a, Bur93a]. Translating [Has17b, JN12]. Translation [Ano94b, Bur93b, Kus81a, MS95, Suz82, Tom81, Yos81b, Hay02a, SW15]. translations [Kaw11]. Translator [Miu87]. Transmission [Lia06, Hor02, Mak15]. transmitted [Kob02]. Transnational [HY15]. Transpacific [Has02b]. Trapezoidal [Yus04]. trapezoids [Yus06]. Trauma [Nak15]. travaux [Hir95b]. Treatise [Ses91, Hay95, Hay02a]. Treatises [Chu11a, Has92b]. treatment [Mor04a, She06c]. treatments [Pan91, Pan92]. Tri [Kus81b]. Tri- [Kus81b]. triangulation [She01a]. Tribute [Ste96]. Trigger [Nis92]. trigonométrie [Has91]. Trigonometry [Miu86, Miu89]. Trilogy [Ano04a]. trois [Ras94b]. True [Hor06, She07]. truth [Cro80]. Tsuboi [Toc18]. Tsunami [Kim11]. tuberculosis [Dai18]. Tunneling [Yam10]. Tunnels [Has10b]. Turbine [Mat97]. turn [Sat12a]. turns [Tan92]. twelve [Luo15]. Twentieth [Dai18, Kim97, Mat02, Loh17, Sat12a]. Twentieth-Century [Dai18, Loh17]. Two [Has83, Kaw07, Kaw17, Kim97, Mur96, Nak80, Mat04b, Sat87]. Tycho [Has92a]. Tyndall [Sug92]. Tyson [Kus14].

U [Nak85]. U.S. [Kar08, Zha07]. U.S.A [Hig01a]. Überblick [Oit82]. Uchida [Tod88a, Kaw81]. Udagawa [Tsu00]. Ueber [Kot06]. ukiyo [Low11]. ukiyo-e [Low11]. Umwälzung [Wäc92]. Unauthenticated [Sug04]. Uncertainty [Oka93, Oka03]. underlies [Mur80]. Underlying [Hog91]. Understanding [Fuj83a, Fuj83b, Tsu94, Duc07]. Understood [Kim10, Ano94a]. Union [Bro07, Kol06, Zha07]. unit [iT86]. United [Sat91, Sat95, Bad05, Kra82, Nis03]. Univ [Oga80]. Univ. [YMN97]. universalis [Hay02b]. Universe [Ano94a]. Universities [Wat90b]. University [Ano94a, Ano94b, Ano94i, Ano01a, Bur93b, Bur93a, Fur98, Has92b, Has99, Hor06, Izu05, Jid91, Miu03, Sat91, Sat95, Sat05, Yaj07, Yam83b, Ano92i, Ano00, Ano02, Ano03a, Ano05b, Kaj08, Kik08, Kis06, Lee06, Nag11, Yam82].

Unknown [GP91]. unpublished [Fre01]. Upland [NM02]. upon [Yam95]. USA [Low89b]. Utrecht [Pye84].

V [Ano94b, Har82, Sza81]. Vāriyar [Hay99b]. vaccination [Shi04]. Vacuum [Juž10]. Vaisesika [Oha95]. Vajkard [Juž07]. Value [HKY89, Hon06, She07, Hol03, Mur92b]. Value-Neutrality [Hon06]. Values [HKY89]. Valvasor [Juž07]. variables [Tak85]. variety [Ano94a]. Various [YO10]. VAT [Yus09, Mur90, Mur91]. Vatican [Ano94b]. Vedic [Hay02a]. Vehicle [Sat01]. Vera [Has08]. Vereinigung [Nak00a]. Verlagsgesellschaft [Nak00a]. version [Ras94b]. Vestigia [Ano94b]. veterum} [Har92b, Har92a]. VI [Sai93]. via [Duc07]. Victims [Sas09]. Victorian [Has16, Ito16, Lig16, Tuc16]. View [Kan84, Kaw91, Kaw81, Yam83a, SR87]. Vieweg [Nak00a]. vii [Yos01]. viii [Bur93b, Kaw82, She06b]. village [Mor04a]. Violation [Gav86]. Virtual [Kon98]. virus [iT86]. Visible [Fuj91, Yos81c]. vision [Nak18]. Visitor [Juž09]. Visser [Har82]. visual [Tuc14]. Visualizing [Yam14]. Vitalism [Mih15]. Vives [Kaw90]. Vol [Ano83a, Ano84a, Ano85a, Ano86a, Ano91e, Ano91f, Ano91g, Ano92b, Ano92c, Ano92d, Ano92e, Ano92j, Ano92k, Ano93a, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95f, Ano95g, Ano95a, Ano95c, Ano95d, Ano95e, Ano95b, Has92a, Ito94, Mur84, Tak84a, Tom81, Yos01, Has91, Kon83a]. Vols [Sas03, Got94, Kon83a, Tom81, Kat04]. Volume [Gup91, Ano94b, Hay94]. vom [Bie84, Wäc92]. Vorlesungen [Nak00a]. Vrin [Ano94a, iS95a, Yos82, Yos84a]. vs [Has83].

W [Har82, Hir95a, Oka93]. W. [Hay99a, Kot06, Nak91]. Waard [Har87b]. während [Wäc92]. Wallis [Mur80]. Wang [Hor00]. War [Ano18, Ara95, Fum02, Has17a, Kim95, KMT89, Nak15, SR87, Sat06b, Yam89, Ich13b, Sat94, Fum02, Ich06, Les06, Lia06, Mai06, NRGY05, Wan06, Yam95]. warfare [iT86]. Warmelehre [Tak87b]. Warp [Has08]. Wartime [Yaj06, Hir18]. was [Ano94b, Kat99, Kob02, Sat07, Zha18]. Wasan [NM02]. Waseda [Ano01a]. waste [Bay12]. Watanabe [Has98a, Kaw86, Miu90, Mur05a]. Wataru [Sat06a]. Water [ANS^{+}82]. Way [Nak85, Ogu06, Tak93b, Tak93a]. Ways [MT82]. Weapons [Bad05, NRGY05, Wal05, YW05]. Weaving [Ham11]. Weft [Has08]. Wegener [Tan90]. Weierstraß [Bie85, Bie85]. Weil [Ada99, Yad00]. Weismann [Oga93]. Welfare [iT84]. wenda [HJ97]. Wende [Wäc92]. Wending [Jam94, Mar81]. Were [Kim10, NM02]. Werken [Sug86]. Werskey [Yos81c]. West [Yat83, Has98a, Mat97, Yab80]. Western [Ano94a, Iya85, Bro83, Cla02, De 03, Han15, Low11, MS03, Mat04b, Nak80, Oza10, Pye97, Pye05, Shi04, Tsu94]. Westernization [Tsu00]. westlichen [Oza10]. where [NM02]. Which [Kat00, Ano94a]. Whirling [Nak11b]. Whitehead [Tan87]. Whitley [Fum89]. Wiener [YR06]. Wiesbaden [Nak00a]. Wilhelm [Mai06]. Wilkins [Sat05, Sat05]. William [Hom05]. Williams [Oka04]. Willis [Yos83b]. Wilson [Has92a]. Wind [Has10b, Pye97]. Winter [Bad10]. Wissenschaft [Wäc92]. wissenschaftliche [IE18]. Wolff [Lor86]. Wolffs [Lor86]. Woman [PO05, Ste08, Kaw93a, Kaw93b]. Women [Kaw11, PO05, Chu11b]. Woods [Miz07]. Words [Iya99]. Work [GP91, Oga89, Wal05, Yam94, Har82, Jam94, KIF96, Sat98a, She00, She06c]. Working [Toy82]. Works [Bro83, Got94, Kus81a, Ōmo84, Sat81, Tak84a, Yos81b, Har92b, Har92a, Sat86, Sat87, Shi92, Sug86]. Workshops [Low89b]. World [Ano18, Bro83, Bur93a, Has10a, Ich06, Kaw91, Lia06, Mai06, NRGY05, Yam95, Ich16b, Fum02, Has17a, Kim95, KMT89, Yam89]. World-View [Kaw91]. Worlds [Izu05, Ger13]. writing [Dho87]. Writings [Har84a, Har84b]. written [Mur98]. wrong [Xia99]. Wu [Hor00]. WW [MS03].

X [Nak00a, Miu87, Loh17, Mur92a]. X-rays [Loh17]. xi [Has98a, Izu05]. xii [Ito94]. xiii [Has10a]. xiv [Ito94]. XIX [Gui86, Yos81a]. xl [Yos81b]. XV [Har87b]. XVI [Har87b, Miu03]. XVIe [Hig01b]. XVII [Pen04]. XVIII [Ano94b, Jam90, Mor04b, DP88, Pen04]. XVIIIe [Ano97a]. xx [Jam90, Kur00, Mor04b, Sas07, She06a]. xxi [Bur93a, Hig01a, Mur84]. xxiii [Oga80, Yus04]. XXXII [Has91].

Y. [Mor04b]. Yāzdī [Ras94a]. Yōichirō [Ito83]. Yōitsu [Got94]. Yōjirō [Yat83, Yat83]. Yū [MHM98]. Yabuuti [Yos81b, Has01]. Yajima [Kik96]. Yamada [Kaw17, Kur00]. Yamamoto [Kat04, Oka98]. Yamaoka [Sak93]. Yamashiro [THI17]. Yamazaki [Kim98]. Yano [Kus81a]. Yasu [Has99]. Yasuke [Kus81a]. Yasushi [Has07a, NKY15]. year [Ano94b]. Yearbook [Fum89]. years [Liu94, Mur84, pZ04]. yen [Hor06, Kan06, Sat06a]. Yoan [Tsu00]. Yoichiro [Set09]. Yokkaichi [FH12]. York [Ano94b]. Yoshida [Ano92i]. Yoshikawa [Yos01]. Yoshinaka [Nak83, Nak85, Wat93]. Yoshio [Sas81b, Hig15, Kim08b]. Yoshitaka [Kat04, Oka98]. Yoshiyuki [Bar15]. Yuan [Yab97]. Yuasa [Ito04, Kaw17, Nak05b, YMN97]. Yugu [Chu11a]. Yukawa [Ano92f, Bro89, Nis92]. Yukiko [Sat94]. Yun [Chu11b]. Yunosawa [Mor04a]. Yuxiu [Che07].

Zahlenrechnungen [Sug86]. zeitgenössischen [Bie84]. zeitgeschichtlicher [Oit82]. Zeitlinger [Har82]. Zentrum [Wac82]. zero [Mur91]. Zimara [Hig01b]. Zipangu [Bie84, Bie84]. Zone [Bar15]. Zoological [Hay04, Miz98]. Zoshu [Sas81b]. zum [Bie84, Wäc92]. Zur [Sza85, Lor86, Nak00a, Tom81]. zwischen [Bie85].

References

Abe:1981:PRR

[Abe81] Yuko Abe. Pauling’s revolutionary role in the development of quantum chemistry. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):107–124, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837258.

Abiko:2000:ETF

[Abi00] Seiya Abiko. Einstein’s theories of the fluctuation and the thermal radiation: the first quantum theory through statistical thermodynamics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):130–147, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Adachi:1995:DMP

[Ada95] Norio Adachi. The dawn of mathematical philosophy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1):1–23, June 1995. CODEN HISCDU. ISSN 0285-4821.

Adachi:1999:ECF

[Ada99] Norio Adachi. Elliptic curves: from Fermat to Weil. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):27–35, July 1999. CODEN HIS-CDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Aiton:1981:CSC

[Ait81] E. J. Aiton. Celestial spheres and circles. Historia Scientiarum = International journal of the History of Science Society of Japan, 19 (44):75–114, 1981. CODEN HISCDU. ISSN 0285-4821.

Amano:1982:DDC

[Ama82] Chikako Amano. La ‘division du canon’ et la théorie mathématique des intervalles musicaux. (French) [The ’division of the canon’ and the mathematical theory of musical intervals]. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):97–115, 1982. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837278.

Amano:2000:TRSb

[Ama00] Kiyoshi Amano. Thermal radiation studies that led to the genesis of quantum theory. I. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10 (2):185–210, November 2000. CODEN HISCDU. ISSN 0285-4821. Translated by Seiji Takata and Shin-ichi Hyodo, Discovery of energy quanta and development of early quantum theory.

Amano:2001:TRSa

[Ama01] Kiyoshi Amano. Thermal radiation studies that led to the genesis of quantum theory. II. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3): 256–280, March 2001. CODEN HISCDU. ISSN 0285-4821. Translated by Seiji Takata and Shin-ichi Hyodo.

Anonymous:1983:CKK

[Ano83a] Anonymous. Contents of the Kagakushi Kenkyu (Journal of the History of Science, Japan), Series II, Vol. 21 (1982). Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):122–123, March 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1983:MBI

[Ano83b] Anonymous. In memoriam: Bun-ichi Tamamushi (1898–1982). Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):103, September 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837447.

Anonymous:1983:ICa

[Ano83c] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 24 (??):124–125, March 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1983:ICb

[Ano83d] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 25 (??):109–110, September 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1983:N

[Ano83e] Anonymous. News. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):104–107, September 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1983:TCHa

[Ano83f] Anonymous. Table of contents of the Historia Scientiarum (no. 23). Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):121, March 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1983:TCHb

[Ano83g] Anonymous. Table of contents of the Historia Scientiarum (no. 24). Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):108, September 1983. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:CKK

[Ano84a] Anonymous. Contents of the Kagakusi Kenkyu (Journal of the History of Science, Japan), Series II, Vol. 22 (1983). Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):128–129, March 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:ICa

[Ano84b] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 26 (??):130–131, March 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:ICb

[Ano84c] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 27 (??):128, September 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:N

[Ano84d] Anonymous. News. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):123–126, September 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:TCHb

[Ano84e] Anonymous. Table of contents of the Historia Scientiarum (no. 26). Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):127, September 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1984:TCHa

[Ano84f] Anonymous. Table of contents of the Historia Scientiarum nos. 18–25 (1981–1983). Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):121–127, March 1984. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1985:CKKa

[Ano85a] Anonymous. Contents of the Kagakusi Kenkyu (Journal of the History of Science, 154 Japan), Series II, Vol. 23. Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):154–155, March 1985. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1985:ICa

[Ano85b] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 28 (??):156, March 1985. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1985:ICb

[Ano85c] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):165, September 1985. CODEN HISCDU. ISSN 0285-4821. Special Issue, “The Historiography of the History of Mathematics”.

Anonymous:1985:N

[Ano85d] Anonymous. News. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):161–163, September 1985. CODEN HISCDU. ISSN 0285-4821. Special Issue, “The Historiography of the History of Mathematics”.

Anonymous:1985:TCHa

[Ano85e] Anonymous. Table of contents of the Historia Scientiarum no. 27 (1984). Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):153, March 1985. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1985:TCHb

[Ano85f] Anonymous. Table of contents of the Historia Scientiarum (no. 28). Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):164, September 1985. CODEN HISCDU. ISSN 0285-4821. Special Issue, “The Historiography of the History of Mathematics”.

Anonymous:1986:CKKb

[Ano86a] Anonymous. Contents of the Kagakushi Kenkyu (Journal of the History of Science, Japan), Series II, Vol. 24 (1985). Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):102–103, March 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1986:ICc

[Ano86b] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 30 (??):104, March 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1986:IC

[Ano86c] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 31 (??):130, September 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1986:N

[Ano86d] Anonymous. News. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):125–127, September 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1986:TCHc

[Ano86e] Anonymous. Table of contents of the Historia Scientiarum no. 29 (1985). Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):101, March 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1986:TCH

[Ano86f] Anonymous. Table of contents of the Historia Scientiarum no. 30 (1986). Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):129, September 1986. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:A

[Ano91a] Anonymous. Announcement. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1 (1):84, June 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:CHSa

[Ano91b] Anonymous. Contents of Historia Scientiarum no. 41 (1991). Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):115, ???? 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:CHSb

[Ano91c] Anonymous. Contents of Historia Scientiarum no. 42 (1991). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):85, June 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:CHSc

[Ano91d] Anonymous. Contents of Historia Scientiarum, no. 43 (1991). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):171, December 1991. CODEN HIS-CDU. ISSN 0285-4821.

Anonymous:1991:CKKa

[Ano91e] Anonymous. Contents of the Kagakusi Kenkyu (Journal of History of Science, Japan), Series II, vol. 29 (no. 176), 1990. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):116, ???? 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:CKKb

[Ano91f] Anonymous. Contents of the Kagakusi Kenkyu (Journal of History of Science, Japan), Series II, vol. 30 (no. 178), 1991. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):86, June 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:CKKc

[Ano91g] Anonymous. Contents of the Kagakusi Kenkyu (Journal of History of Science, Japan), Series II, vol. 30 (no. 179), 1991. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):172, December 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:EHS

[Ano91h] Anonymous. Errata of Historia Scientiarum no. 42. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):167–169, December 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:EGF

[Ano91i] Anonymous. Errata to “The Growth of Fourier’s Theory of Heat Conduction and His Experimental Study (in no. 43). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):170, December 1991. CODEN HISCDU. ISSN 0285-4821. See [YT91].

Anonymous:1991:IC

[Ano91j] Anonymous. Information for contributors. Historia Scientiarum = International journal of the History of Science Society of Japan, 42 (??):117–??, ???? 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:LE

[Ano91k] Anonymous. Letter to the Editor. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):83, June 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1991:N

[Ano91l] Anonymous. News. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):77–82, June 1991. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:ASI

[Ano92a] Anonymous. Announcement: The Seventh International Conference on the History of Science in East Asia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):161–166, December 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:CHSa

[Ano92b] Anonymous. Contents of Historia Scientiarum, vol. 1, no. 1–3. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):234–235, March 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:CHSb

[Ano92c] Anonymous. Contents of Historia Scientiarum vol. 1, no. 3 (no. 45). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):80, June 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:CKKa

[Ano92d] Anonymous. Contents of the Kagakusi Kenkyu (Journal of History of Science, Japan), Series II, vol. 30 (no. 180) 1991. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):236–237, March 1992. CODEN HIS-CDU. ISSN 0285-4821.

Anonymous:1992:CKKb

[Ano92e] Anonymous. Contents of the Kagakusi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 181), 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):81, June 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:EFS

[Ano92f] Anonymous. Errata to “Forum: Satio Hayakawa’s Letter to Hideki Yukawa (19 Sept. 1955)” by Sigeko Nisio (in No. 44). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):79, June 1992. CODEN HISCDU. ISSN 0285-4821. See [Nis92].

Anonymous:1992:Na

[Ano92g] Anonymous. News. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):229, March 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:Nb

[Ano92h] Anonymous. News. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):76–79, June 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:ODM

[Ano92i] Anonymous. Obituary: [Dr. Mitsukuni Yoshida, Professor Emeritus of Kyoto University]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3): 225–228, March 1992. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1992:TCKa

[Ano92j] Anonymous. Table of contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 182), 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):167, December 1992. CODEN HIS-CDU. ISSN 0285-4821.

Anonymous:1992:TCKb

[Ano92k] Anonymous. Table of contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 183), 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):168, December 1992. CODEN HIS-CDU. ISSN 0285-4821.

Anonymous:1993:CPV

[Ano93a] Anonymous. Call for papers (vol. 3, no 3: History of Japanese science). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):202, March 1993. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1993:NAM

[Ano93b] Anonymous. New: Annual meeting of History of Science Society. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):103–106, July 1993. CO-DEN HISCDU. ISSN 0285-4821.

Anonymous:1993:NRA

[Ano93c] Anonymous. New: Report of the activities of the National Committee of Japan and the History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):107–114, July 1993. CODEN HIS-CDU. ISSN 0285-4821.

Anonymous:1993:TCHa

[Ano93d] Anonymous. Table of contents of Historia Scientiarum, vol. 2, no. 1 (no. 46). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):239, March 1993. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1993:TCHb

[Ano93e] Anonymous. Table of contents of Historia Scientiarum, vol. 2, no. 2 (no. 47). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):240, March 1993. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:BRJ

[Ano94a] Anonymous. Books received: Jean Gayon, ed., Buffon 88: Actes du Colloque international pour le bicentenaire de la mort de Buffon, Paris, J. Vrin, 1992. A Collection of essays compiled in memory of Jacques Roger. Forty six essays discuss a variety of aspects of Buffon’s scientific activities. Liba Chaia Taub, Ptolemy’s Universe: The Natural Philosophical and Ethical Foundations of Ptolemy’s Astronomy, Chicago and LaSalle, Illinois, Open Court, 1993. An inquiry which sheds light on not only the natural philosophical but also ethical aspects of Ptolemy’s Almagest. Claudine Cohen, Le destin du Mammouth, Paris, Éditions du Seuil, 1994. A story as to how Western thinkers such as Leibniz and Buffon understood mammoth. H. Floris Cohen, The Scientific Revolution: A Historiographical Inquiry, Chicago and London, The University of Chicago Press, 1994. A broad research on how to define the Scientific Revolution and why it occurred in seventeenth century Europe in comparison with Islam countries and China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):179, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Anonymous:1994:BRM

[Ano94b] Anonymous. Books received: M. Folkerts and J. P. Hogendijk, eds., Vestigia Mathematica: Studies in Medieval and Early Modern Mathematics in Honour of H. L. L. Busard, Rodopi, Amsterdam & Atlanta, 1993. The book was edited on the occasion of H. L. L. Busard’s 70th year of birth, containing papers by P. Bockstaele, H. J. M. Bos, M. Folkerts, B. Hughes, G. Molland, B. A. Rosenfeld, C. J. Scriba, J. Sesiano and 11 others and the list of publications by Busard. Eric Brian, La mesure de l’État: Administrateurs et geomètres au XVIII^{ieme} siècle, Albin Michel, Paris, 1994, one volume of L’Évolution de l’Humanité, sheds light on a political aspect of the mathematical thought of 18th century France. John S. Major, Heaven and Earth in Han Thought: Chapters Three, and Four, and Five of the Huainanzi, State University of New York Press, Albany, 1993, discusses a cosmological theory in Han China, containing Christopher Gullen’s essay A Chinese Eratosthenes of the Flat Earth as Appendix A. Annibale Fantoli, Galileo, For Copernicantsm and for the Church, Translation by George V. Coyne, Vatican Observatory Publications, Vatican City State, 1994, is a volume for Galileo’s rehabilitation in the Catholic Church. The author has once taught at Sophia University, Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4 (1):67, August 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CHSa

[Ano94c] Anonymous. Contents of the Historia Scientiarum, vol. 3, no. 1 (no. 49). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):249, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CHSb

[Ano94d] Anonymous. Contents of the Historia Scientiarum, vol. 3, no. 2 (no. 50). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):250, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CKKa

[Ano94e] Anonymous. Contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 184) 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):245, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CKKb

[Ano94f] Anonymous. Contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 185) 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):246, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CKKc

[Ano94g] Anonymous. Contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 186) 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):247, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:CKKd

[Ano94h] Anonymous. Contents of the Kagakushi Kenkyu (Journal of History of Science, Japan), Series II, vol. 31 (no. 187) 1992. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):248, March 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1994:NAM

[Ano94i] Anonymous. News: The 42nd Annual Meeting of the History of Science Society of Japan on May 28–29, 1994 at Ritsumeikan University, Kyoto. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):61–65, August 1994. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1995:CKKa

[Ano95a] Anonymous. Contents of Kagakusi Kenkyu (Journal of History of Science, Japan) Series II, Vol. 32 (No. 188), 1993. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):257, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CKKe

[Ano95b] Anonymous. Contents of Kagakusi Kenkyu (Journal of History of Science, Japan) Series II, Vol. 32 (No. 192), 1994. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):261, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CKKb

[Ano95c] Anonymous. Contents of Kagakusi Kenkyu (Journal of History of Science, Japan) Series II, Vol. 33 (No. 189), 1994. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):258, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CKKc

[Ano95d] Anonymous. Contents of Kagakusi Kenkyu (Journal of History of Science, Japan) Series II, Vol. 33 (No. 190), 1994. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):259, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CKKd

[Ano95e] Anonymous. Contents of Kagakusi Kenkyu (Journal of History of Science, Japan) Series II, Vol. 33 (No. 191), 1994. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):260, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CHSa

[Ano95f] Anonymous. Contents of the Historia Scientiarum vol. 4, no. 1 (no. 52). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):263, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:CHSb

[Ano95g] Anonymous. Contents of the Historia Scientiarum vol. 4, no. 2 (no. 53). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):264, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Anonymous:1995:ESI

[Ano95h] Anonymous. Errata: Shokichi Iyanaga, Évolution des études mathématiques au Japon depuis l’ère Meiji, 4(3) p. 205 March (1995). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):223, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan. See [Iya95].

Anonymous:1995:NAM

[Ano95i] Anonymous. News: The 42nd Annual Meeting of the History of Science Society of Japan, June 3–4, 1995. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):221–223, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Anonymous:1996:NAM

[Ano96] Anonymous. News: The 43rd Annual Meeting of the History of Science Society of Japan, May 25–26, 1996. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):63–65, July 1996. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1997:AEC

[Ano97a] Anonymous. Announcement: École CIMPA, Histoire dés mathematiques en Méditerranée de l’Antiquité au XVIIIe siècle, 12–26 septembre 1997 à La Mansourah (Égypte). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3):243, March 1997. CODEN HISCDU. ISSN 0285-4821. March.

Anonymous:1997:NRN

[Ano97b] Anonymous. News: Report on the National Committee of Japan and on the History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):71–79, June 1997. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1997:NAM

[Ano97c] Anonymous. News: The 44th Annual Meeting of the History of Science Society of Japan, May 31–June 1, 1997. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):67–69, June 1997. CODEN HISCDU. ISSN 0285-4821.

Anonymous:1998:NAM

[Ano98] Anonymous. News: The 45th Annual Meeting of the History of Science Society of Japan, May 30–31, 1998. Historia Scientiarum.

Second Series. International Journal of the History of Science Society of Japan, 8(2):207–210, December 1998. CODEN HISCDU. ISSN 0285-4821. See errata [Ano99].

Anonymous:1999:ENA

[Ano99] Anonymous. Errata: News: The 45th Annual Meeting of the History of Science Society of Japan, May 30–31, 1998. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(3):273, March 1999. CODEN HISCDU. ISSN 0285-4821. See [Ano98].

Anonymous:2000:NAM

[Ano00] Anonymous. News: The 47th Annual Meeting of History of Science Society of Japan on May 20 and 21, 2000, at St. Andrew’s University. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):100–103, July 2000. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2001:NAM

[Ano01a] Anonymous. News: The Annual Meeting of the History of Science Society of Japan, May 26–27, 2001, at Waseda University, Tokyo. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):??, July 2001. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2001:RAN

[Ano01b] Anonymous. Report on the activities of the National Committee of Japan and the History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):??, July 2001. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2002:NAM

[Ano02] Anonymous. News: The 49th Annual Meeting of History of Science Society of Japan, May 25 and 26, 2002 at Kanazawa University. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):??, July 2002. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2003:NAM

[Ano03a] Anonymous. News: The 50th Annual Meeting of the History of Science Society of Japan, June 7 and 8, 2003, at Kobe University. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):??, July 2003. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2003:NB

[Ano03b] Anonymous. Notices of books. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13 (2):??, November 2003. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2004:NSC

[Ano04a] Anonymous. News: Symposium to commemorate the completion of a trilogy on “The Historical Theory of Modern Science” by Prof. Chikara Sasaki. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):??, March 2004. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2004:NAM

[Ano04b] Anonymous. News: The 51st Annual Meeting of the History of Science Society of Japan, May 29 and 30, 2004, at Tokyo Institute of Technology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(1):??, July 2004. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2005:NBa

[Ano04c] Anonymous. Notices of books. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14 (2):??, November 2004. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2005:NTC

[Ano05a] Anonymous. News: Table of contents of Kagakusi Kenkyu (2004). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):??, July 2005. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2005:NAM

[Ano05b] Anonymous. News: The 52nd Annual Meeting of the History of Science Society of Japan, June 4 and 5, 2005, at the Historical Museum of Hokkaido and Sapporo Gakuin University. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):??, July 2005. CODEN HISCDU. ISSN 0285-

Anonymous:2005:NBb

[Ano05c] Anonymous. Notices of books. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15 (1):??, July 2005. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2005:RAN

[Ano05d] Anonymous. Report on the activities of the National Committee of Japan and the History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):??, July 2005. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2006:NAM

[Ano06a] Anonymous. News: The 53rd Annual Meeting of the History of Science Society Program. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16 (2):??, November 2006. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2006:NB

[Ano06b] Anonymous. Notices of books. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15 (3):??, March 2006. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2007:NAM

[Ano07] Anonymous. News: The 54th Annual Meeting of the History of Science Society Program. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17 (1):??, July 2007. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2009:AJU

[Ano09] Anonymous. Announcement: Journals under threat: A joint response from history of science, technology and medicine. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1):??, July 2009. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2011:NPA

[Ano11a] Anonymous. News: The Program of the 57th Annual Meeting of the History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(3):??, March 2011. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2011:SPH

[Ano11b] Anonymous. Statement by The President of The History of Science Society of Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(2):??, December 2011. CODEN HISCDU. ISSN 0285-4821.

Anonymous:2013:TC

[Ano13] Anonymous. Table of contents. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):237–238, March 2013. CODEN HISCDU. ISSN

0285-4821. URL http://hssj.info/wp-content/uploads/2017/
07/Vol22-3_Contents_Vol22.pdf.

Anonymous:2018:ECS

[Ano18] Anonymous. Erratum: The construction of the system of aeronautical standards for safe air navigation before World War II (Historia Scientiarum). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):278, 2018. CODEN HISCDU. ISSN 0285-4821.

Aoki:1982:SWM

[ANS^{+}82] Kunio Aoki, Tohru Nakagawa, Seibun Sakakibara, Akira Sasabe, and Toshiyoshi Kikuchi. A survey of water mills in Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):81–97, 1982. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837289.

Aramaki:1987:FRT

[Ara87] Seiya Aramaki. Formation of the renormalization theory in quantum electrodynamics. Historia Scientiarum = International journal of the History of Science Society of Japan, 32(??):1–42, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837358.

Aramaki:1989:DRTa

[Ara89a] Seiya Aramaki. Development of the renormalization theory in quantum electrodynamics (I). Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??): 97–116, 1989. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435094;   https://ci.nii.ac.jp/
naid/110009837391.

Aramaki:1989:DRTb

[Ara89b] Seiya Aramaki. Development of the renormalization theory in quantum electrodynamics (II). Historia Scientiarum = International journal of the History of Science Society of Japan, 37(??): 91–113, 1989. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435099;   https://ci.nii.ac.jp/
naid/110009837399.

Arakawa:1995:BFT

[Ara95] Kiyoshi Arakawa. Basic features of the technical development in Japan during the post-war half century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):199–214, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Atsushi:1997:RSH

[Ats97] Katsuki Atsushi. A rough sketch of history of solid state physics in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):107–124, December 1997. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Baber:2005:SCP

[Bab05] Zaheer Baber. Science, colonialism and the production of science and modernity in India. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2): 139–158, November 2005. CODEN HISCDU. ISSN 0285-4821.

Badash:2005:USD

[Bad05] Lawrence Badash. The United States and the development of nuclear weapons. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(3):160–163, March 2005. CODEN HISCDU. ISSN 0285-4821.

Badash:2010:NWE

[Bad10] Lawrence Badash. Nuclear winter and the end of the age of agriculture. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(2):150–163, December 2010. CODEN HISCDU. ISSN 0285-4821.

Bartholomew:1994:PST

[Bar94] James R. Bartholomew. Perspectives on science and technology in Japan: The career of Fukui Ken’ichi. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):47–54, August 1994. CODEN HISCDU. ISSN 0285-4821.

Bartholomew:2015:BRY

[Bar15] James R. Bartholomew. Book review: Yoshiyuki Kikuchi, Anglo-American Connections in Japanese Chemistry: The Lab as Contact Zone (London: Palgrave-MacMillan, 2013). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(2):99–101, February 2015. CODEN HISCDU. ISSN 0285-4821.

Bay:2012:NBD

[Bay12] Alexander R. Bay. Nation from the bottom up: Disease, toilets and waste management in Prewar Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(2):142–158, December 2012. CODEN HISCDU. ISSN 0285-4821.

Bay:2018:DEC

[Bay18a] A. R. Bay. Disease, environment, and causation: Roundworm prevention and eradication in 20th-Century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):150–173, 2018. CODEN HISCDU. ISSN 0285-4821.

Bay:2018:SIN

[Bay18b] A. R. Bay. Special issue new directions in the history of medicine in East Asia: Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27 (2):125–126, 2018. CODEN HISCDU. ISSN 0285-4821.

Biermann:1984:JLZ

[Bie84] Kurt-R. Biermann. Japan: vom legendaren Zipangu bis zum zeitgenössischen Nippon bei Alexander von Humboldt. (German) [Japan: from the legendary Zipangu to the contemporary Nippon according to Alexander von Humboldt]. Historia Scientiarum = International journal of the History of Science Society of Japan, 27 (??):1–10, September 1984. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837316.

Biermann:1985:KSP

[Bie85] Kurt-R. Biermann. Kontroversen um den Steiner-Preis und ihre Folgen: Ein Kapitel aus den Beziehungen zwischen Weierstraß und Kronecker. (German) [Controversy over the Steiner Prize and its consequences: a chapter in the relationship between Weierstraß and Kronecker]. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):117–124, September 1985. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac. jp/naid/110009837343. Special Issue, “The Historiography of the History of Mathematics”.

Blackmore:1988:MCP

[Bla88] John Blackmore. Mach competes with Planck for Einstein’s favor. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):45–89, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435088;
https://ci.nii.ac.jp/naid/110009837382.

Blay:1999:MMC

[Bla99] Michel Blay. Méthodes mathématiques et calcul de l’infini au temps de Fermat. (French) [Mathematical methods and calculation of infinity in Fermat’s time]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9 (1):57–71, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Boantza:2006:LPC

[Boa06] Victor D. Boantza. Light in the pneumatic context: Aspects of interplay between theory and practice in early photochemical research. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):105–128, November 2006. CODEN HISCDU. ISSN 0285-4821.

Botelho:1989:SSB

[Bot89] Antonio Jose J. Botelho. Struggling to survive: The Brazilian Society for the Progress of Science (SBPC) and the authoritarian regime (1964–1980). Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):45–63, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837407.

Brookman:1983:HSN

[Bro83] F. Henry Brookman. The history of science in the non-Western world: An inventarisation of some important works during the 20th Century. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):93–100, September 1983.

CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837445.

Brown:1989:YGR

[Bro89] Laurie M. Brown. Yukawa in the 1930s: A gentle revolutionary. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):1–21, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435089;
https://ci.nii.ac.jp/naid/110009837386.

Brooks:2004:CLR

[Bro04] Nathan M. Brooks. Chemistry laboratories in Russia: 1700–1917. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(2):89–100, November 2004. CODEN HISCDU. ISSN 0285-4821.

Brooks:2007:RBC

[Bro07] Nathan M. Brooks. The road to “big chemistry” in the Soviet Union during the 1920s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(3):264–274, March 2007. CODEN HISCDU. ISSN 0285-4821.

Burnett:1993:BRS

[Bur93a] Charles Burnett. Book review: S.ā’id al-Andalusī, Science in the Medieval World “Book of the Categories of Nations”, translated and edited by Sema’am I. Salem and Alok Kumar, (Austin: University of Texas Press, 1991, xxi + 118 pp., $29.00). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(2):157–158, December 1993. CODEN HISCDU. ISSN 0285-4821.

Burnett:1993:BRK

[Bur93b] Charles Burnett. Book review: Ken’ichi Takahashi, The Medieval Latin Traditions of Euclid’s Catoptrica: A Critical Edition of De speculis with an Introduction, English Translation and Commentary, (Fukuoka: Kyusyu University Press, 1992, viii + 373 pp., Yen 10,300). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):95–98, July 1993. CODEN HISCDU. ISSN 0285-4821.

Burns:2012:RPK

[Bur12] Susan L. Burns. Relocating psychiatric knowledge: Meiji psychiatrists, local culture(s), and the problem of fox possession. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(2):88–109, December 2012. CODEN HISCDU. ISSN 0285-4821.

Chayut:2005:PSR

[Cha05] Michael Chayut. From the periphery to scientific revolutions. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):1–22, July 2005. CODEN HIS-CDU. ISSN 0285-4821.

Chemla:1991:TAC

[Che91] Karine Chemla. Theoretical aspects of the Chinese algorithmic tradition (First to Third Century). Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):71–94, ???? 1991. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435118;
https://ci.nii.ac.jp/naid/110009837435.

Chemla:1994:DCE

[Che94] Karine Chemla. Different concepts of equations in The nine chapters on mathematical procedures [four Chinese characters] and in the commentary on it by Liu Hui (3rd century). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):113–137, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Chen:2007:GYI

[Che07] Shou Chen. Gu Yuxiu’s ideas of science and scientization in context of saving China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2):89–102, December 2007. CODEN HISCDU. ISSN 0285-4821.

Chikara:1994:AMT

[Chi94] Sasaki Chikara. Asian mathematics from traditional to modern. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):69–77, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Christianidis:1996:MPS

[Chr96] Jean Christianidis. Maxime Planude sur le sens du terme diophantien plasmatikón. (French) [Maxime Planude on the meaning of the term diophantine plasmatikón]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):37–41, July 1996. CODEN HISCDU. ISSN 0285-4821.

Chung:2011:AIC

[Chu11a] Hyung-Min Chung. Agricultural illustrations of 19th century Korea: ‘Imwon gyeongjeji’ (treatises on management of forest and garden) by Seo Yugu. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(1):42–65, July 2011. CODEN HISCDU. ISSN 0285-4821.

Chung:2011:ACG

[Chu11b] Saehyang Chung. Art, class and gender in Joseon dynasty Korea: representations of lower-class women by the scholar–painter Yun Duseo. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(1):20–42, July 2011. CODEN HISCDU. ISSN 0285-4821.

Carazza:1990:OEN

[CK90] Bruno Carazza and Helge Kragh. An Oriental in Europe: Nagaoka, Righi, and the state of physics 1910. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??): 37–44, 1990. CODEN HISCDU. ISSN 0285-4821. URL https://

ci.nii.ac.jp/naid/110009837421.

Clancey:2002:FKC

[Cla02] Gregory Clancey. Foreign knowledge: Cultures of Western sciencemaking in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(3): 245–260, March 2002. CODEN HISCDU. ISSN 0285-4821.

Costabel:1985:DMI

[Cos85] Pierre Costabel. Descartes et la mathématique de l’infini. (French) [Descartes and the mathematics of infinity]. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):37–49, September 1985. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837340. Special Issue,

“The Historiography of the History of Mathematics”.

Crombie:1980:SAR

[Cro80] A. C. Crombie. Science and the arts in the Renaissance: the search for truth and certainty, old and new. Historia Scientiarum = International journal of the History of Science Society of Japan, 18(42): 233–246, 1980. CODEN HISCDU. ISSN 0285-4821.

Daidoji:2018:FCT

[Dai18] K. Daidoji. The formation of constitutional (Taishitsu) medicine in Early Twentieth-Century Japan: The scrofulous constitution (Senbyoshitsu) and tuberculosis. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27 (2):199–217, 2018. CODEN HISCDU. ISSN 0285-4821.

Darrigol:1988:ESB

[Dar88] Olivier Darrigol. Elements of a scientific biography of Tomonaga Sin-itiro. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):1–29, 1988. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435085;   https://ci.nii.ac.jp/naid/110009837379.

DeMaio:2003:DES

[De 03] Silvana De Maio. The development of an educational system at the beginning of the Meiji Era: Reference models from Western countries transfer of agricultural technology from imperialist. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3):183–199, March 2003. CODEN HIS-CDU. ISSN 0285-4821.

Dhombres:1985:FMT

[Dho85] J. Dhombres. French mathematical textbooks from Bézout to Cauchy. Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):91–137, March 1985. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837329.

Dhombres:1987:TER

[Dho87] Jean Dhombres. Sur un texte d’Euler relatif à une équation fonctionnelle: Archaismes, pédagogie, et style d’écriture. (French) [On a text of Euler relating to a functional equation: Archaisms, pedagogy, and writing style]. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):77–99, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435079;   https://ci.nii.ac.jp/naid/110009837369.

Dimoia:2009:APA

[Dim09] John Dimoia. Atoms for power?: The Atomic Energy Research Institute (AERI) and South Korean electrification, (1948–1965). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):170–183, December 2009. CODEN HISCDU. ISSN 0285-4821.

Dhombres:1988:ERP

[DP88] J. Dhombres and M. Pensivy. Esprit de rigueur et présentation mathématique au XVIIIème siècle: le cas d’une démonstration d’Aepinus. (French) [Spirit of rigor and mathematical presentation in the Eighteenth Century: the case of a demonstration of Aepinus]. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):21–42, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435082;
https://ci.nii.ac.jp/naid/110009837374.

Ducheyne:2007:HUT

[Duc07] Steffen Ducheyne. Huygens’s understanding of trajectory: via media between Galileo and Newton. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17 (1):1–19, July 2007. CODEN HISCDU. ISSN 0285-4821.

Euler:1987:RPM

[Eul87] L. Euler. Un remarquable perfectionnement de la méthode inverse des tangentes. (French) [A remarkable improvement of the inverse method of tangents]. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):100–123, 1987. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837370.

Ezawa:2001:OTT

[Eza01] Hiroshi Ezawa. Obituary: Takehiko Takabayashi (1919–1999). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):286–304, March 2001. CO-DEN HISCDU. ISSN 0285-4821.

Fares:1995:AAM

[Far95] Nicolas Farès. Aspects analytiques dans la mathématique de Sharaf al-Dīn al-T .ūsī. (French) [Analytical aspects in the mathematics of Sharaf al-Dīn al-T .ūsī]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1): 39–55, June 1995. CODEN HISCDU. ISSN 0285-4821.

Feng:2012:YAB

[FH12] Danyang Feng and Takehiko Hashimoto. Yokkaichi asthma and building the system of certification of pollution-related patients in the 1960s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(3):211–227, March 2012. CODEN HISCDU. ISSN 0285-4821.

Freire:1997:QCM

[Fre97] Olival Freire, Jr. Quantum controversy and Marxism. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):137–152, December 1997. CODEN HIS-CDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Freire:2001:SPP

[Fre01] Olival Freire, Jr. Science, philosophy and politics in the fifties: on the Max Born’s unpublished paper entitled “Dialectical materialism and modern physics”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):248–254, March 2001. CODEN HISCDU. ISSN 0285-4821.

Fujisaki:1982:PDT

[Fuj82] Chiyoko Fujisaki. P. Drude’s theory of dispersion of light and atomic model (1900–1913). Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):19–67, 1982. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837275.

Fujisaki:1983:DKD

[Fuj83a] Chiyoko Fujisaki. From Deslandres to Kratzer (I): Development of the understanding of the origin of infrared band spectra (1880– 1913). Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):53–75, March 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837298.

Fujisaki:1983:DKI

[Fuj83b] Chiyoko Fujisaki. From Deslandres to Kratzer (II): Development of the understanding of the origin of infrared band spectra (1913– 1920). Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):57–86, September 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837443.

Fujii:1988:BRK

[Fuj88] Kiyohisa Fujii. Book review: Kozo Hirota, Meiji no Kagakusha (Japanese Chemists in the Meiji Era), 1988, 230 pp., Yen 1,400, Tokyo Kagaku Dojin, Tokyo. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):94–95, 1988. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837385.

Fujisaki:1991:RHD

[Fuj91] Chiyoko Fujisaki. The role of T. Heurlinger in the development of modern visible molecular spectroscopy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):133–162, December 1991. CODEN HISCDU. ISSN 0285-4821.

Fujioka:2016:JLA

[Fuj16] Tsuyoshi Fujioka. Japanese Lysenkoists after Lysenko’s downfall. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):15–24, March 2016. CO-DEN HISCDU. ISSN 0285-4821.

Fumihiko:1989:BRS

[Fum89] Satofuka Fumihiko. Book review: Sociology of the Sciences Yearbook 1987 ‘The Social Direction of the Public Sciences’, Edited by Stuart Blume, Joske Bunders, Loet Leydesdorff and Richard Whitley, D. Reidel Publishing Company. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):73–74, 1989. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837409.

Fumihiko:2002:MAH

[Fum02] Tochinai Fumihiko. The methodological aspect of the “historicity of controversies” in the Japanese geological community after World War II: The change of attitude to physicochemical methods before and after the War. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1): 59–73, July 2002. CODEN HISCDU. ISSN 0285-4821.

Furukawa:1982:HSE

[Fur82] Yasu Furukawa. Hermann Staudinger and the emergence of the macromolecular concept. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):1–18, 1982. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837274.

Furukawa:1996:CHH

[Fur96] Yasu Furukawa. From chemistry to history: Historians of chemistry and their community in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6 (2):87–108, December 1996. CODEN HISCDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Furukawa:1997:BRM

[Fur97] Yasu Furukawa. Book review: Mary Jo Nye, Before Big Science: The Pursuit of Modern Chemistry and Physics, 1800–1940 (London: Prentice Hall, 1996). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7 (1):65–66, June 1997. CODEN HISCDU. ISSN 0285-4821.

Furukawa:1998:BRM

[Fur98] Yasu Furukawa. Book review: Masanori Kaji, Mendeleev no Shūkiritsu Hakken (Mendeleev’s Discovery of the Periodic Law) (Sapporo: Hokkaido University Press, 1997). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):203–205, December 1998. CODEN HISCDU. ISSN 0285-4821.

Furukawa:2007:IGD

[Fur07] Yasu Furukawa. Introduction: Globalization and diversity in chemical sciences. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(3):221–223, March 2007. CODEN HISCDU. ISSN 0285-4821.

Gavroglu:1986:REF

[Gav86] K. Gavroglu. Recent evidence and further comments on parity violation. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):115–123, September 1986. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435076;   https://ci.nii.ac.jp/naid/110009837357.

Gerovitch:2013:PWF

[Ger13] Slava Gerovitch. Parallel worlds: Formal structures and informal mechanisms of postwar Soviet mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):181–200, March 2013. CODEN HISCDU. ISSN 0285-4821.

Grattan-Guinness:2001:MMA

[GG01] Ivor Grattan-Guinness. Manifestations of mathematics in and around the Christianities: some examples and issues. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):48–85, July 2001. CODEN HISCDU. ISSN 0285-4821.

Goto:1994:BRY

[Got94] Kunio Gotō. Book review: Yōitsu Kondō (edited by Chikara Sasaki), Sūgakushi Chosakushū (Collected Works on History of Mathematics), 5 vols. (Tokyo: Nihon Hyoronsha, 1994). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):175–178, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Goto:2000:DEQ

[Got00a] Kunio Goto. Discovery of energy quanta and development of early quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):105–111, November 2000. CODEN HISCDU. ISSN 0285-4821. Historia Sci.

(2) 10 (2000), no. 2.

Goto:2000:EQC

[Got00b] Kunio Goto. Einstein’s quantum conditions in early quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):148–155, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Gold:1991:HUS

[GP91] David Gold and David Pingree. A hitherto unknown Sanskrit work concerning Mādhava’s derivation of the power series for sine and cosine. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):49–64, ???? 1991. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435116;   https://ci.nii.ac.jp/naid/110009837433.

Guitard:1986:QIP

[Gui86] Thierry Guitard. La querelle des infiniment petits a l’École Polytechnique au XIX^{e} siècle. (French) [The quarrel over the infinitely small at the Polytechnical School in the 19th Century]. Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):1–61, March 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435068;
https://ci.nii.ac.jp/naid/110009837347.

Gupta:1991:VSA

[Gup91] Radha Charan Gupta. On the volume of a sphere in Ancient India. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):33–44, ???? 1991. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435114;   https://ci.nii.ac.jp/naid/110009837431.

Hakata:1990:DPA

[Hak90] Toshio Hakata. The diplomatic policy of Accademia dei Lincei and its cultural background. Historia Scientiarum = International journal of the History of Science Society of Japan, 39(??): 1–14, 1990. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435103;   https://ci.nii.ac.jp/
naid/110009837411.

Hammers:2011:PEQ

[Ham11] Roslyn Lee Hammers. Perspectives in Early Qing-Dynasty Pictures of tilling and weaving and pictures of cotton. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(1):3–19, July 2011. CODEN HISCDU. ISSN 0285-4821.

Han:2015:CLA

[Han15] Qi Han. Chinese Literati’s attitudes toward Western science: Transition from the Late Kangxi period to the Mid-Qianlong period. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(2):76–87, February 2015. CODEN HISCDU. ISSN 0285-4821.

Hara:1982:BRS

[Har82] Kokiti Hara. Book review: Studies on Christiaan Huygens, Invited papers from the symposium on the life and work of Christiaan Huygens, Amsterdam, 22–25 August 1979, Edited by H. J. M. Bos, M. J. S. Rudwick, H. A. M. Snelders, R. P. W. Visser. Swets & Zeitlinger B.V. Lisse, 1980, v + 321 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):123–130, 1982. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837280.

Hara:1984:NOEa

[Har84a] Kokiti Hara. Nouvelles observations sur les Écrits mathématiques de Pascal. (French) [New observations on the mathematical writings of Pascal]. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):1–17, March 1984. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837306.

Hara:1984:NOEb

[Har84b] Kokiti Hara. Nouvelles observations sur les Écrits mathématiques de Pascal (II). (French) [New observations on the mathematical writings of Pascal (II)]. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):11–25, September 1984. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837317.

Hara:1985:CDD

[Har85] Kokiti Hara. Comment Descartes a-t-il découvert ses ovales?. (French) [how did Descartes discover his ovals?]. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):51–82, September 1985. CODEN HIS-

CDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/

110009837341. Special Issue, “The Historiography of the History of Mathematics”.

Hara:1987:NKN

[Har87a] K. Hara. Necrologie: Koshiro Nakamura (1901–1986). (French) [Obituary: Koshiro Nakamura (1901–1986)]. Historia Scientiarum = International journal of the History of Science Society of Japan, 32(??):147–150, 1987. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837363.

Hara:1987:BRC

[Har87b] Kokiti Hara. Book review: Correspondance du P. Marin Mersenne, religieux minime, commencée par Paul Tannery, publiée et annotée par Cornelis de Waard et Armand Beaulieu, Éditions du Centre National de la Recherche Scientifique, Tome XV, 1983, pièces limin. et 670p., et t. XVI, 1986, pièces limin. et 592 p., 23.5cm × 18.5cm. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):125–128, 1987. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837371.

Hara:1992:QOGb

[Har92a] Kokiti Hara. Quelques ouvrages de géométrie more veterum de Roberval. I. (French) [Some more veterum geometry works from Roberval. I]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):13–44, June 1992. CODEN HISCDU. ISSN 0285-4821. With unpublished paper, “Quaedam ad librum quintum elementorum Euclidis”, by Gilles Personne de Roberval.

Hara:1992:QOGa

[Har92b] Kokiti Hara. Quelques ouvrages de géométrie more veterum de Roberval. II. (French) [Some more veterum geometry works from Roberval. II]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):109–117, December 1992. CODEN HISCDU. ISSN 0285-4821.

Hassan:1980:SOH

[Has80] Ahmad Y. Hassan. Some obstacles hindering the advance of science and technology in the Arab countries. Historia Scientiarum = International journal of the History of Science Society of Japan, 19 (??):31–46, 1980. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837245.

Hashimoto:1983:AVB

[Has83] Takehiko Hashimoto. Ampère vs. Biot: Two mathematizing routes to electromagnetic theory. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):29–51, March 1983. CODEN HISCDU. ISSN 0285-4821. URL https://

ci.nii.ac.jp/naid/110009837297.

Hashimoto:1989:HDI

[Has89] Takehiko Hashimoto. Huygens, dioptrics, and the improvement of the telescope. Historia Scientiarum = International journal of the History of Science Society of Japan, 37(??):51–90, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435098;   https://ci.nii.ac.jp/naid/110009837398.

Hashimoto:1991:BRM

[Has91] Keizo Hashimoto. A mathematical outcome of the strong tradition of French Sinology; Catherine Jami, Les méthodes rapides pour la trigonométrie et le rapport précis du cercle (1774): tradition chinois et apport occidental en mathématiques, Mémoires de l’Institut des Hautes Études Chinoises, vol. XXXII, Collège de France, 1990. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):111–114, ???? 1991. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837437.

Hashimoto:1992:BRS

[Has92a] Keizo Hashimoto. Book review: Scientific Revolution in the Context of Astronomical Transformation: René Taton & Curtis Wilson (eds.), Planetary Astronomy from the Renaissance to the Rise of Astrophysics: Tycho Brahe to Newton (The General History of Astronomy, Vol. 2, Part A). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1 (3):163–172, March 1992. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:1992:BRK

[Has92b] Takehiko Hashimoto. Book review: Kimiyo Koyanagi, Pasukaru: Chokkan kara Dantei made: Butsuri Ronbun Kansei madeno Dotei (Pascal: From Intuition to Affirmation — The Process of Completing Physics Treatises) (Nagoya: Nagoya University Press, 1992). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):159–160, December 1992. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:1994:GCE

[Has94] Takehiko Hashimoto. Graphical calculation and early aeronautical engineers. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):159–183, March 1994. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:1995:OJN

[Has95] Keizo Hashimoto. Obituary: Joseph Needham (1900–1995). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):215–219, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Hashimoto:1998:BRM

[Has98a] Takehiko Hashimoto. Book review: Masao Watanabe, Science and Cultural Exchange in Modern History: Japan and the West (Tokyo: Hokusen-sha, 1997). xi + 406 pp. Yen14,000. ISBN 4-938424-73-8. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):247–250, March 1998. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:1998:SAR

[Has98b] Takehiko Hashimoto. Science after 1940: Recent historical researches and issues on postwar American science and technology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):87–96, August 1998. CO-DEN HISCDU. ISSN 0285-4821.

Hashimoto:1999:BRY

[Has99] Takehiko Hashimoto. Book review: Yasu Furukawa, Inventing Polymer Science: Staudinger, Carothers, and the Emergence of Molecular Chemistry (Philadelphia: University of Pennsylvania Press, 1998). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):121–??, July 1999. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2001:PDK

[Has01] Keizo Hashimoto. Obituary: Professor Dr. Kiyosi Yabuuti (12/02/1906–02/06/2000). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10 (3):305–309, March 2001. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2002:BRI

[Has02a] Takehiko Hashimoto. Book review: Ian Inkster and Fumihiko Satofuka, eds., Culture and Technology in Modern Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(3):268–270, March 2002. CODEN HIS-CDU. ISSN 0285-4821.

Hashimoto:2002:CSP

[Has02b] Takehiko Hashimoto. The contest over the standard: The project of the transpacific flight and aeronautical research in Interwar Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(3):226–244, March 2002. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2007:BRY

[Has07a] Takehiko Hashimoto. Book review: Yasushi Sato, NASA o Kizuita Hito to Gijutsu (NASA and Its Engineering Cultures: Human Qualities in Grand System Building), 2007. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2):155–157, December 2007. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2007:LBS

[Has07b] Takehiko Hashimoto. Leonard Bairstow as a scientific middleman: Early aerodynamic research on airplane stability in Britain, 1909– 1920. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2):103–120, December 2007. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2008:BRF

[Has08] Takehiko Hashimoto. Book review: Francesca Bray, Vera Dorofeeva-Lichtmann, and George Métailié, eds., Graphics and Text in the Production of Technical Knowledge in China: The Warp and the Weft, 2007. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):253–255, March 2008. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2010:BRL

[Has10a] Takehiko Hashimoto. Book review: Luís Saraiva, and Catherine Jami, eds., The Jesuits, the Padroado and East Asian Science (1552–1773), (London: World Scientific, 2008), xiii + 229 pp. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(1):57–60, July 2010. CODEN HIS-CDU. ISSN 0285-4821.

Hashimoto:2010:HFD

[Has10b] Takehiko Hashimoto. “How far do experiments on models represent experiments on full sized machines?”: The examination and dispute on the reliability of the wind tunnels in Britain, 1909–1917. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(2):96–122, December 2010. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2010:ISI

[Has10c] Takehiko Hashimoto. Introduction: Special issue: Historical studies on scientific experiments and instruments. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(2):61–62, December 2010. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2011:I

[Has11] Takehiko Hashimoto. Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(3):165–166, March 2011. CODEN HISCDU. ISSN 0285-

Hashimoto:2012:IST

[Has12a] Takehiko Hashimoto. Introduction: Science, technology, and the state in East Asia: Experts and politicians in postwar Korea and Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(3):159–160, March 2012. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2012:JCT

[Has12b] Takehiko Hashimoto. The Japanese clocks and time in the past: Non-standard seasonal time inscribed on scale plates of foot-ruler clocks. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(1):22–39, July 2012. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2014:I

[Has14a] Takehiko Hashimoto. Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(3):165–166, March 2014. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2014:OCS

[Has14b] Takehiko Hashimoto. Observing cracks, sparks, and snow crystals: Torahiko Terada and his students’ pursuit for the “physics of form”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(3):214–240, March 2014. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2016:SIP

[Has16] Takehiko Hashimoto. Special issue the popularization of science in Victorian Britain: introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(3):161–163, March 2016. CODEN HISCDU. ISSN 0285-4821.

Hashimoto:2017:CSA

[Has17a] Takehiko Hashimoto. The construction of the system of aeronautical standards for safe air navigation before World War II. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(1):4–24, August 2017. CODEN HIS-CDU. ISSN 0285-4821.

Hashimoto:2017:TAE

[Has17b] Takehiko Hashimoto. Translating and annotating Edison’s biography in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(3): 192–209, March 2017. CODEN HISCDU. ISSN 0285-4821.

Hayashi:1991:NBR

[Hay91] Takao Hayashi. A note on Bhāskara I’s rational approximation to sine. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):45–48, ???? 1991. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435115;   https://ci.nii.ac.jp/naid/110009837432.

Hayashi:1994:RPP

[Hay94] Eiji Hayashi. A reconstruction of the proof of Proposition 11 in Archimedes’s method: proofs about the volume and the center of the gravity of any segment of an obtuse-angled conoid. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):215–230, March 1994. CODEN HIS-CDU. ISSN 0285-4821.

Hayashi:1995:SDA

[Hay95] Takao Hayashi. Śrīdhara’s authorship of the mathematical treatise Gan.itapañcavim .śī. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):233– 250, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Hayashi:1996:BRC

[Hay96] Takao Hayashi. A bibliography (1958–1995) of Radha Charan Gupta, historian of Indian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):43–53, July 1996. CODEN HISCDU. ISSN 0285-4821.

Hayashi:1998:CTD

[Hay98a] Makoto Hayashi. Cell theory in its development and inheritance in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):115–132, December 1998. CODEN HISCDU. ISSN 0285-4821.

Hayashi:1998:IMG

[Hay98b] Tomohiro Hayashi. Introducing movement into geometry: Roberval’s influence on Leibniz’s Analysis situs. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):53–69, August 1998. CODEN HISCDU. ISSN 0285-4821.

Hayashi:1999:IFW

[Hay99a] Haruo Hayashi. The influence of Fourier on W. Thomson. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(2):165–180, November 1999. CODEN HISCDU. ISSN 0285-4821.

Hayashi:1999:SRR

[Hay99b] Takao Hayashi. A set of rules for the root-extraction prescribed by the Sixteenth-Century Indian mathematicians, Nīlakan.t.ha Somastuvan and Śaṅkara Vāriyar. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9 (2):135–153, November 1999. CODEN HISCDU. ISSN 0285-4821.

Hayashi:2002:RLL

[Hay02a] Takao Hayashi. Review of: Līlāvatī of Bhāskarācārya: a treatise of mathematics of Vedic tradition with rationale in terms of modern mathematics largely based on N. H. Phadke’s Marāthī translation of Līlāvatī [Motilal Banarsidass Publishers, Delhi, 2001] by Krishnaji Shankara. Patwardhan, Somashekhara Amrita. Naimpally and Shyam Lal Singh. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(2): 168–174, November 2002. CODEN HISCDU. ISSN 0285-4821.

Hayashi:2002:LCM

[Hay02b] Tomohiro Hayashi. Leibniz’s construction of Mathesis universalis: a consideration of the relationship between the plan and his mathematical contributions. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(2): 121–141, November 2002. CODEN HISCDU. ISSN 0285-4821.

Hayashi:2004:BJZ

[Hay04] Makoto Hayashi. The beginning of the Japanese Zoological Journal. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):241–249, March 2004. CODEN HISCDU. ISSN 0285-4821.

He:2007:ECT

[He07] Juan He. The establishment of Chinese terms for chemical elements: 1901–1932. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(3):275–293, March 2007. CODEN HISCDU. ISSN 0285-4821.

Heeffer:2006:LDD

[Hee06] Albrecht Heeffer. The logic of disguise: Descartes’ discovery of the Law of Refraction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2): 144–165, November 2006. CODEN HISCDU. ISSN 0285-4821.

Higashi:2001:BRC

[Hig01a] Shinichiro Higashi. Book review: Constance Blackwell, Sachiko Kusukawa, eds., Philosophy in the Sixteenth and Seventeenth Centuries: Conversations with Aristotle. xxi + 415 pp., figs., bibl., index. Aldershot, Great Britain/Brookfield, U.S.A.: Ashgate, 1999. ISBN 0-86078-668-4. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1): 127–129, July 2001. CODEN HISCDU. ISSN 0285-4821.

Higashi:2001:PMA

[Hig01b] Shin’ichiro Higashi. Penser les mathématiques au XVIe siècle. La philosophie des mathématiques chez Marcantonio Zimara (1475/76 — avant 1537). (French) [Thinking mathematics in the sixteenth century. The philosophy of mathematics of Marcantonio Zimara (1475/76 — before 1537)]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11 (2):143–167, November 2001. CODEN HISCDU. ISSN 0285-4821.

Higuchi:2015:SCD

[Hig15] Toshihiro Higuchi. The strange career of Dr. Fish: Yoshio Hiyama, radioactive fallout, and nuclear fear management in Japan, 1954– 1958. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):57–77, August 2015. CODEN HISCDU. ISSN 0285-4821.

Hijioka:1990:BRH

[Hij90] Yoshito Hijioka. Book review: I. Hronszky, M. Feher & B. Dajka (eds)., Scientific Knowledge Socialized, Dordrecht, Boston, London: Kluwer Academic Publishers, 1988. Pp. ix + 440, ISBN 90-288-2284-6, $124.00. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):63–64, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837425.

Hinokawa:1989:OIA

[Hin89] Shizue Hinokawa. Origin of the idea of the atomic bomb: Part 2. Changing status of the plutonium bomb plan in the early stages of American atomic bomb development. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):15–28, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435101;
https://ci.nii.ac.jp/naid/110009837405.

Hirano:1984:NDT

[Hir84] Yoichi Hirano. Note sur les diffusions de la théorie de Galois: Première clarification des idées de Galois par Liouville. (French) [Note on the diffusion of the theory of Galois: First clarification of the ideas of Galois by Liouville]. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??): 27–41, September 1984. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837318.

Hirano:1995:BRS

[Hir95a] Yoichi Hirano. Book review: Sasaki Chikara, Sugiura Mitsuo and Joseph W. Dauben, eds., The Intersection of History and Mathematics (Basel: Birkhäuser, 1994). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4 (3):255–256, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Hirano:1995:QRT

[Hir95b] Yoīchi Hirano. Quelques remarques sur les travaux de Lagrange — qui concernent la théorie des équations algébriques et la notion préliminaire de groupes. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1): 75–84, June 1995. CODEN HISCDU. ISSN 0285-4821.

Hirokawa:2018:WMP

[Hir18] W. Hirokawa. When medicine participates in field studies: Epidemiological research of Hansen’s disease during pre- and wartime Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):218–232, 2018. CO-DEN HISCDU. ISSN 0285-4821.

Hua:1982:RCA

[HJ82] Jueming Hua and Yunfu Jia. A research on the Chinese ancient chime bells: Its history, casting process, design and calculation. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):63–79, 1982. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837288.

Hashimoto:1997:KLC

[HJ97] Keizo Hashimoto and Catherine Jami. Kepler’s laws in China: a missing link? Jean-François Foucquet’s Lifa wenda [4 Chinese characters]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3):171–185, March 1997. CODEN HISCDU. ISSN 0285-4821. March.

Hayashi:1989:IVD

[HKY89] Takao Hayashi, Takanori Kusuba, and Michio Yano. Indian values for π derived from Aryabhata’s value. Historia Scientiarum = International journal of the History of Science Society of Japan, 37(??):1–16, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435095;
https://ci.nii.ac.jp/naid/110009837395.

Hashimoto:2017:SIP

[HM17] Takehiko Hashimoto and Ruselle Meade. Special issue the popularization of science & technology in Japan: Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(2):75–76, February 2017. CODEN HIS-CDU. ISSN 0285-4821.

Hirokawa:1989:SSH

[HO89] Shunkichi Hirokawa and Shuzo Ogawa. Shoichi Sakata: His physics and methodology. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):67–81, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435092;   https://ci.nii.ac.jp/naid/110009837389.

Hogendijk:1991:AKT

[Hog91] Jan P. Hogendijk. Al-Khwārizmī’s table of the “sine of the hours” and underlying sine table. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):1–12, ???? 1991. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435112;   https://ci.nii.ac.jp/
naid/110009837429.

Hollenback:2003:AEI

[Hol03] George M. Hollenback. Another example of an implied pi value of 3[+]1/8 in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(2):120–122, November 2003. CODEN HISCDU. ISSN 0285-4821.

Home:2005:WSS

[Hom05] Roderic W. Home. William Sutherland and the ‘Sutherland– Einstein’ diffusion relation: theoretical physics in a colonial setting. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):125–138, November 2005. CODEN HISCDU. ISSN 0285-4821.

Honma:1996:BNP

[Hon96] Eio Honma. Beeckman’s natural philosophy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(3):225–247, March 1996. CODEN HISCDU. ISSN 0285-4821.

Hong:2006:KSL

[Hon06] Sungook Hong. Korean scientists look at Japanese colonization: The emergence of an idea of the value-neutrality and objectivity. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):268–275, March 2006. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110006459063.

Horiuchi:1982:FIC

[Hor82] Tatsuo Horiuchi. La formation des ingènieurs civils en France 1795– 1848. (French) [The training of civil engineers in France 1795–1848]. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):45–61, 1982. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837287.

Horiuchi:1987:SCT

[Hor87] Annick Horiuchi. La science calendérique de Takebe Katahiro (1664–1739). (French) [The calendrical science of Takebe Katahiro (1664–1739)]. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):3–24, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435077;   https://ci.nii.ac.jp/naid/110009837365.

Horng:1995:HDL

[Hor95] Wann-Sheng Horng. How did Liu Hui perceive the concept of infinity: a revisit. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):207–222, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Horng:2000:DML

[Hor00] Wann-Sheng Horng. Disseminating mathematics in late 19thcentury China: the case with Wang Kangnian and the Shi Wu Bao. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):46–57, July 2000. CO-DEN HISCDU. ISSN 0285-4821.

Horng:2002:SKT

[Hor02] Wann-Sheng Horng. Sino–Korean transmission of mathematical texts in the 19th century: a case study of Nam Pyong-gil’s Kugo Sulyo Tohae. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(2):87–99, November 2002. CODEN HISCDU. ISSN 0285-4821.

Horiuchi:2006:BRS

[Hor06] Annick Horiuchi. Book review: Sato Ken-Ichi, Kinsei Nihon sugakushi, Seki Takakazu no jitsuzo o motomete (History of Mathematics in Early Modern Japan: Searching for the True Portrait of Seki Takakazu), Sugakushi series (History of Mathematics Series). Tokyo: Tokyo Daigaku Shuppankai (University of Tokyo Press), 2005. 423 pp. ISBN: 4-13-061355-3. 6,825 yen. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):279–283, March 2006. CODEN HISCDU. ISSN 0285-4821.

Hoyrup:2003:NIP

[Høy03] Jens Høyrup. A note on: “Inheritance problems in the Susa mathematical text no. 26” [Historia Sci. (2) 10 (2001), no. 3, 226–234; MR1845726] by Kazuo Muroi. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):54–58, July 2003. CODEN HISCDU. ISSN 0285-4821. See [Mur01].

Hamblin:2015:BLD

[HR15] Jacob Darwin Hamblin and Linda M. Richards. Beyond the Lucky Dragon. Japanese scientists and fallout discourse in the 1950s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):36–56, August 2015. CODEN HISCDU. ISSN 0285-4821.

Hirota:1983:DAR

[HY83] Kozo Hirota and Tatsuaki Yamaguchi. Did S. Arrhenius really propose the Arrhenius equation? Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):53–55, September 1983. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837442.

Higuchi:2015:ITO

[HY15] Toshihiro Higuchi and Masakatsu Yamazaki. Introduction: Transnational origins of the Russell–Einstein Manifesto and the radiological dimensions of the nuclear arms race. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):1–7, August 2015. CODEN HISCDU. ISSN 0285-4821.

Ichikawa:1995:TTO

[Ich95] Hiroshi Ichikawa. Technological transformation of Occupied Japan: The implications of the policies and activities of the Scientific and Technical Division of the Economic and Scientific Section of GHQ/ SCAP. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):183–198, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Ichikawa:2006:IPH

[Ich06] Hiroshi Ichikawa. Introduction: A perspective on the historical study of science and technology during the Second World War and the Cold War in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):1–4, July 2006. CODEN HISCDU. ISSN 0285-4821.

Ichikawa:2013:ISI

[Ich13a] Hiroshi Ichikawa. Introduction: Special issue — science and Soviet political authorities: Conflict, cooperation, and incongruence. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):159–160, March 2013. CODEN

HISCDU. ISSN 0285-4821. URL http://hssj.info/wp-content/
uploads/2017/07/Vol22-3_02.pdf.

Ichikawa:2013:SPD

[Ich13b] Hiroshi Ichikawa. Soviet physicists during the war: Jealousy, discord and the ideological dispute. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22 (3):215–226, March 2013. CODEN HISCDU. ISSN 0285-4821.

Ichikawa:2015:RSS

[Ich15] Hiroshi Ichikawa. Radiation studies and the Soviet scientists in the second half of the 1950’s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25 (1):78–93, August 2015. CODEN HISCDU. ISSN 0285-4821.

Ichikawa:2016:SIS

[Ich16a] H. Ichikawa. Special issue: Soviet science beyond the boundaries: introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):1–2, March 2016. CODEN HISCDU. ISSN 0285-4821.

Ichikawa:2016:OWF

[Ich16b] Hiroshi Ichikawa. Obninsk, 1955: The world’s first nuclear power plant and “the atomic diplomacy” by Soviet scientists. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):25–41, March 2016. CODEN HIS-CDU. ISSN 0285-4821.

Ichikawa:2017:MK

[Ich17] Hiroshi Ichikawa. Masanori Kaji (1956–2016). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(2):??, February 2017. CODEN HISCDU. ISSN 0285-4821.

Ishihara:2018:ASC

[IE18] A. Ishihara and U. Enke. About the scientific career of the bacteriologist Taichi Kitashima and the influence of Shibasaburô Kitasato and Emil von Behring [Über die wissenschaftliche Karriere des Bakteriologen Taichi Kitashima und den Einfluss Shibasaburo Kitasatos und Emil von Behrings]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27 (2):254–277, 2018. CODEN HISCDU. ISSN 0285-4821.

Ihara:1995:DEP

[Iha95] Satoshi Ihara. Development of electric power technology and social framework in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2): 127–166, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Iida:1982:HSH

[iI82] Ken ichi Iida. Hiroto Saigusa and his philosophy of technology. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):134–138, 1982. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837294.

Ikeyama:2010:BRS

[Ike10] Setsuro Ikeyama. Book review: Sakhya Ganitasarakaumudi: The Moonlight of the Essence of Mathematics by Thakkura Pheru, 2009, ISBN 978-81-70304-809-8, 278 pp., Rs. 995. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(3):239–241, March 2010. CODEN HISCDU. ISSN 0285-4821.

Isahaya:2017:ERH

[IL17] Yoichi Isahaya and Jyuh Fuh Lin. Entangled representation of heaven: A Chinese divination text from a tenth-century Dunhuang fragment (p. 4071). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(3): 153–171, March 2017. CODEN HISCDU. ISSN 0285-4821.

Imai:2011:HCH

[Ima11] Masahiro Imai. Herophilus of Chalcedon and the Hippocratic tradition in early Alexandrian medicine. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(2):103–122, December 2011. CODEN HISCDU. ISSN 0285-4821.

Imai:2012:ADA

[Ima12] Masahiro Imai. Aristotle on the debate about the central organ of the human body in the 5th and 4th centuries BC. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(1):1–21, July 2012. CODEN HISCDU. ISSN 0285-4821.

Imai:2015:ECT

[Ima15] Masahiro Imai. Erasistratus of Ceos and the theoretical sources for his anatomical physiology of a human being. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(3):103–125, March 2015. CODEN HISCDU. ISSN 0285-4821.

Sato:1994:ODK

[iS94] Ken ichi Sato. Obituary: Dr. Kazuo Shimodaira (1928–1994). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):55–56, August 1994. CODEN HIS-CDU. ISSN 0285-4821.

Sato:1995:BRA

[iS95a] Ken ichi Sato. Book review: Annick Horiuchi, Les mathématiques japonaises à l’époque d’Edo (Paris: Librairie Philosophique, J. Vrin, 1994). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1):85–87, June 1995. CODEN HISCDU. ISSN 0285-4821.

Sato:1995:RTT

[iS95b] Ken ichi Sato. Reevaluation of Tengenjutsu or Tianyuanshu: in the context of comparison between China and Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1):57–67, June 1995. CODEN HISCDU. ISSN 0285-4821.

Ishiguro:2006:RMC

[Ish06] Takehiko Ishiguro. Research misconduct in cutting-edge material physics and electronics at the beginning of the 21st Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):41–61, July 2006. CODEN HISCDU. ISSN 0285-4821.

Tsuneishi:1984:CKP

[iT84] Ken ichi Tsuneishi. C. Koizumi as a promoter of the Ministry of Health and Welfare and an originator of the BCW Research Program. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):95–113, March 1984. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837310.

Tsuneishi:1986:RGM

[iT86] Kei ichi Tsuneishi. The research guarded by military secrecy: The isolation of the E.H.F. virus in Japanese biological warfare unit. Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):79–92, March 1986. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435070;   https://ci.nii.ac.jp/naid/110009837349.

Tsuneishi:1992:LE

[iT92] Kei ichi Tsuneishi. Letter to the editors. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):230–233, March 1992. CODEN HISCDU. ISSN 0285-4821.

Itagaki:1982:WDM

[Ita82] Ryoichi Itagaki. Why did Mach reject Einstein’s theory of relativity? Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):81–95, 1982. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837277.

Ito:1983:BRY

[Ito83] Kazuyuki Ito. Book review: Yōichirō Murakami, Kagakushi no Gyakuenkinhō (The Reverse-perspectivism in the History of Science) — A Revaluation of the Renaissance Thought, Chūōkōron Sha, Tokyo, 1982, 306 pp. Yen 1500. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??): 115–117, March 1983. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837302.

Ito:1994:BRM

[Ito94] Kenji Ito. Book review: Mitsuo Taketani and Masayuki Nagasaki, Ryoshi Rikigaku no Keisei to Ronri (The Formation and the Logic of Quantum Mechanics). Vol. 1: Genshi Mokei no Keisei (The Formation of Atomic Model). Tokyo: Ginza Shuppansha, 1948, 282 pp.; repr. ed. Tokyo: Keiso Shobo, 1972, 300 pp. Yen 1,800. Vol. 2: Ryoshi Rikigaku eno Michi (Path to Quantum Mechanics). Tokyo: Keiso Shobo, 1991, xii + 267 pp. Yen 3,800. Vol. 3: Ryoshi Rikigaku no Seiritsu to sono Ronri (The Completion of Quantum Mechanics and Its Logic), Tokyo: Keiso Shobo, 1993, xiv + 302 pp. Yen 4,400. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):235–245, March 1994. CO-DEN HISCDU. ISSN 0285-4821.

Ito:2005:GPE

[Ito04] Kenji Ito. Gender and physics in Early 20th Century Japan: Yuasa Toshiko’s case. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(2):118–136, November 2004. CODEN HISCDU. ISSN 0285-4821.

Ito:2011:GPD

[Ito11] Kazuyuki Ito. Galileo’s principle of descending motion along inclined planes. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(2):93–102, December 2011. CODEN HISCDU. ISSN 0285-4821.

Ito:2016:SPS

[Ito16] Takashi Ito. The state and the popularization of science in Victorian Britain: Scientific and Literary Societies act of 1843. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(3):216–251, March 2016. CODEN HIS-CDU. ISSN 0285-4821.

Iyama:1983:CFD

[Iya83] Hiroyuki Iyama. A case of fabricated discovery: The Law of Multiple Proportions. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):19–28, March 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837296.

Iyanaga:1985:BRT

[Iya85] Shokichi Iyanaga. Book review: Tamotsu Murata, Nihon no Sūgaku, Seiyō no Sūgaku, Hikaku Sūgakushi no Kokoromi (Japanese Mathematics and Western Mathematics: An Attempt of Comparative History of Mathematics), Chūō Kōronsha, 1981. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):157–158, September 1985. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837345. Special

Issue, “The Historiography of the History of Mathematics”.

Iyanaga:1995:EEM

[Iya95] Shokichi Iyanaga. Évolution des études mathématiques au Japon depuis l’ère Meiji. (French) [Evolution of mathematical studies in Japan since the Meiji era]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):181–206, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2. See errata [Ano95h].

Iyanaga:1999:GWO

[Iya99] S. Iyanaga. Greeting words at the opening of the Colloquium “Pierre de Fermat mathematicien français”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):1–2, July 1999. CODEN HISCDU. ISSN 0285-4821.

Izumi:1997:TSR

[Izu97] Chiye Izumi. Timaean science — the role of stereometry in Greek cosmology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):45–57, June 1997. CODEN HISCDU. ISSN 0285-4821.

Izumi:2005:BRG

[Izu05] Chiye Izumi. Book review: G. E. R. Lloyd, Ancient Worlds, Modern Reflections: Philosophical Perspectives on Greek and Chinese Science and Culture. Oxford: Oxford University Press, 2004. Pp. xi + 222. ISBN: 0-19-927016-3. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15 (1):69–72, July 2005. CODEN HISCDU. ISSN 0285-4821.

Jadhav:2002:LLI

[Jad02] Dipak Jadhav. The laws of logarithms in India. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(3):261–267, March 2002. CODEN HIS-CDU. ISSN 0285-4821.

Jami:1990:BRJ

[Jam90] Catherine Jami. Book review: Jean-Claude Martzloff, Histoire des Mathématiques Chinoises, Paris, Masson, 1988, xx + 375 p., Prefaces by Jacques Gernet and Jean Dhombres (Table of contents in English p. xviii), FF 295.00. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):59–62, 1990. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837424.

Jami:1991:SMK

[Jam91] Catherine Jami. Scholars and mathematical knowledge during the Late Ming and Early Qing. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):95–110, ???? 1991. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435119;   https://ci.nii.ac.jp/
naid/110009837436.

Jami:1994:HMM

[Jam94] Catherine Jami. History of mathematics in Mei Wending’s (1633– 1721) work. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):159–174, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Jido:1991:BRS

[Jid91] Yuji Jido. Book review: Shoichi Oyodo, Miyamoto Takenosuke to Kagaku-Gijutu-Seisaku (Miyamoto Takenosuke and the Policy of Science and Technology), (Tokyo: Tokai University Press, 1989). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):71–74, June 1991. CO-DEN HISCDU. ISSN 0285-4821.

Jimenez:1998:SFF

[Jim98] José Antonio Cervera Jiménez. Spanish friars in the Far East: Fray Juan Cobo and his book Shi Lu. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):181–198, March 1998. CODEN HISCDU. ISSN 0285-4821.

Jannetta:2012:TMK

[JN12] Ann Jannetta and Ellen Nakamura. Translating medical knowledge: Japan’s first medical journal — Taisei meii ik. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(2):68–87, December 2012. CODEN HISCDU. ISSN 0285-4821.

Johnston:2018:CPC

[Joh18] W. Johnston. Cholera and popular culture in Nineteenth Century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):174–198, 2018. CO-DEN HISCDU. ISSN 0285-4821.

Jullien:1999:FMC

[Jul99] Vincent Jullien. Les frontières dans les mathématiques cartésiennes. (French) [The boundaries in Cartesian mathematics]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(3):211–238, March 1999. CODEN HISCDU. ISSN 0285-4821.

Juznic:2007:JVV

[Juž07] Stanislav Južnič. Janez Vajkard Valvasor (1641–1693), F.R.S. and his scientific and technological connections to East Asia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2):134–152, December 2007. CODEN HISCDU. ISSN 0285-4821.

Juznic:2009:CSA

[Juž09] Stanislav Južnič. Chinese–Slovenian astronomer Hallerstein: Provincial and visitor to Japanese Jesuit province. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1):55–71, July 2009. CODEN HIS-CDU. ISSN 0285-4821.

Juznic:2010:FEV

[Juž10] Stanislav Južnič. Far Eastern vacuum and electricity: Augustin Hallerstein and experimental correspondence between Beijing and Europe. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(1):21–46, July 2010. CODEN HISCDU. ISSN 0285-4821.

Juznic:2015:BJM

[Juž15] Stanislav Južnič. Bibliometrics of Jesuit mathematicians in China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(3):126–151, March 2015. CODEN HISCDU. ISSN 0285-4821.

Kaji:2003:ODN

[Kaj03] Masanori Kaji. Ogawa’s discovery of a new chemical element “nipponium”: The emergence of modern chemistry research in Japan and its social background. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12 (3):215–218, March 2003. CODEN HISCDU. ISSN 0285-4821.

Kaji:2008:NTC

[Kaj08] Masanori Kaji. Nozoe Tetsuo’s chemical research at Taihoku Imperial University in Taiwan and its colonial context. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2):132–139, December 2008. CODEN HIS-CDU. ISSN 0285-4821.

Kamatani:1995:SBR

[Kam95] Chikayoshi Kamatani. Sake brewing and its records in Edo Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):117–126, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Kaneko:1984:EVJ

[Kan84] Tsutomu Kaneko. Einstein’s view of Japan’s culture. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):51–76, September 1984. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837320.

Kangnian:1988:INI

[Kan88] Yan Kangnian. On Isaac Newton’s ideas of gravitation and God. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):43–56, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435083;
https://ci.nii.ac.jp/naid/110009837375.

Kanamori:2006:BRS

[Kan06] Osamu Kanamori. Book review: Shigehisa Kuriyama and Kazutoshi Kitazawa eds., Kindai Nihon no Shintai Kankaku (Bodily Sensation in Modern Japan) (2004), 426 pp., ISBN 4-7872-3236-3 C0036, 3,000 yen. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):276–278, March 2006. CODEN HISCDU. ISSN 0285-4821.

Kanayama:2013:BIS

[Kan13] Koji Kanayama. Between ideology and science: Dialectics of dispute on physics in 1920s–1930s Soviet Russia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):201–214, March 2013. CODEN HISCDU. ISSN 0285-4821.

Kargon:2008:UJN

[Kar08] Robert Kargon. The U.S.–Japan nexus: Roots of the Japanese scientific enterprise. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2):62–75, December 2008. CODEN HISCDU. ISSN 0285-4821.

Kargon:2010:MMM

[Kar10] Robert Kargon. Making Manila modern: Science, technology and American colonialism 1898–1915. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(3):209–224, March 2010. CODEN HISCDU. ISSN 0285-4821.

Katsumori:1998:BEC

[Kat98] Makoto Katsumori. Bohr’s early complementarity argument. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):1–19, August 1998. CODEN HIS-CDU. ISSN 0285-4821.

Kato:1999:HFL

[Kat99] Kazuya Kato. How Fermat’s Last Theorem was proved. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):49–56, July 1999. CODEN HIS-CDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Katsuki:2000:ECJ

[Kat00] Atsushi Katsuki. An experimental contribution from Japan to decide which is better: Old- or new-quantum theory — measurement of magnetic susceptibility of hydrogen gas by Soné Také. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):177–184, November 2000. CODEN HISCDU. ISSN 0285-4821.

Katsumori:2005:BRY

[Kat04] Makoto Katsumori. Book review: Yoshitaka Yamamoto, Jiryoku to Jûyoku no Hakken [The Discovery of Magnetism and Gravitation], 3 vols. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(2):154–155, November 2004. CODEN HISCDU. ISSN 0285-4821.

Kawajiri:1981:ACV

[Kaw81] Nobuo Kawajiri. The acceptance of the Christian view of science by a Japanese in the Tokugawa Era: The establishment of Shosho-gaku by Gokan Uchida. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):67–86, 1981. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837267.

Kawamoto:1982:BRK

[Kaw82] Hideo Kawamoto. Book review: Kenzo Sakamoto, Kikai no Genshogaku (The Phenomenology of Machinery), Iwanami-shoten, Tokyo, 1975, viii + 331 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):131–133, 1982. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837293.

Kawamoto:1983:TLE

[Kaw83] Hideo Kawamoto. The theory of life in the Early 20th Century: In Japan and in Europe. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):17–27, September 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837439.

Kawajiri:1984:BRN

[Kaw84] Nobuo Kawajiri. Book review: Nobuhiro Miyoshi, Meiji no Engineer Kyoiku (Technical Education in the Meiji Era), Chūō Kōron Sha, Tokyo, 1983, 219 pp., Yen 460. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):119–120, March 1984. CODEN HISCDU. ISSN 0285-4821. URL https://

ci.nii.ac.jp/naid/110009837312.

Kawajiri:1986:BRM

[Kaw86] Nobuo Kawajiri. Book review: Masao Watanabe, ed., Newton no Hikari to Kage (Newton: Lights and Shadows), 1982, 286 pp., Yen 2200; Kyoritsu Shuppan Sha, Tōkyō/Masao Watanabe, ed., Darwin to Shinkaron (Darwin and Darwinism: the Historical Context), 1984, 288 pp., Yen 2200; Kyōritsu Shuppan Sha, Tōkyō. Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):99–100, March 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837351.

Kawajiri:1988:BRS

[Kaw88] Nobuo Kawajiri. Book review: Shuntaro Ito, ed., Chusei no Sugaku (Mathematics in the Middle Ages), 1987, 671 pp., Yen 12,000, Kyoritsu Shuppan Sha, Tokyo. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):85–86, 1988. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837378.

Kawashima:1990:PMD

[Kaw90] Keiko Kawashima. La participation de Madame du Châtelet à la querelle sur les forces vives. Historia Scientiarum = International journal of the History of Science Society of Japan, 40(??): 9–28, 1990. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435107;   https://ci.nii.ac.jp/
naid/110009837415.

Kawahara:1991:WVS

[Kaw91] Hideki Kawahara. World-view of the Santong-Li. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):65–70, ???? 1991. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435117;
https://ci.nii.ac.jp/naid/110009837434.

Kawashima:1993:ISMa

[Kaw93a] Keiko Kawashima. Les idées scientifiques de Madame du Châtelet dans ses institutions de physique: un rêve de femme de la haute société dans le culture scientifique au Siècle des Lumières. 1ère partie. (French) [The scientific ideas of Madame du Châtelet in her physics institutions: a dream of a woman of high society in the scientific culture of the Century of Enlightenment. Part 1]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):63–82, July 1993. CODEN HISCDU. ISSN 0285-4821.

Kawashima:1993:ISMb

[Kaw93b] Keiko Kawashima. Les idées scientifiques de Madame du Châtelet dans ses institutions de physique: un rêve de femme de la haute société dans le culture scientifique au Siècle des Lumières. 2ème partie. (French) [The scientific ideas of Madame du Châtelet in her physics institutions: a dream of a woman of high society in the scientific culture of the Century of Enlightenment. Part 2]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(2):137–156, December 1993. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2004:BAM

[Kaw04] Keiko Kawashima. Birth of ambition: Madame du Châtelet’s Institutions de physique. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(1): 49–66, July 2004. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2005:IGS

[Kaw05] Keiko Kawashima. The issue of gender and science: A case study of Madame du Châtelet’s Dissertation sur le feu. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):23–43, July 2005. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2007:TPA

[Kaw07] Keiko Kawashima. Two popular accounts of Émilie du Châtelet and the gender problem. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17 (2):121–133, December 2007. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2011:WTS

[Kaw11] Keiko Kawashima. Women’s translations of scientific texts in the 18th century: a case study of Marie-Anne Lavoisier. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(2):123–137, December 2011. CODEN HIS-CDU. ISSN 0285-4821.

Kawashima:2013:EGQ

[Kaw13] Keiko Kawashima. The evolution of the gender question in the study of Madame Lavoisier. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(1): 24–37, July 2013. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2015:EDC

[Kaw15] Keiko Kawashima. Émilie Du Châtelet and the significance of publishing: Dissertation sur la nature et la propagation du feu (1744). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(2):95–118, December 2015. CODEN HISCDU. ISSN 0285-4821.

Kawashima:2017:NYT

[Kaw17] Keiko Kawashima. Nobuo Yamada and Toshiko Yuasa: Two Japanese scientists and the Curie family. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(1):108–124, August 2017. CODEN HISCDU. ISSN 0285-4821.

Kikuchi:1996:RCC

[KIF96] Yoshiyuki Kikuchi, Kenji Ito, and Yasu Furukawa. Rudolf Clausius’ copy of Sadi Carnot’s Réflexions: a new look at Clausius’ access to Carnot’s work. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):31–36, July 1996. CODEN HISCDU. ISSN 0285-4821.

Kikuchi:1996:OMP

[Kik96] Toshiyoshi Kikuchi. Obituary: To the memory of Professor Suketoshi Yajima. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):55–58, July 1996. CODEN HISCDU. ISSN 0285-4821.

Kikuchi:2000:RAC

[Kik00] Yoshiyuki Kikuchi. Redefining academic chemistry: Joji Sakurai and the introduction of physical chemistry into Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(3):215–256, March 2000. CODEN HISCDU. ISSN 0285-4821.

Kikuchi:2008:AFE

[Kik08] Yoshiyuki Kikuchi. Analysis, fieldwork and engineering: Accumulated practices and the formation of applied chemistry teaching at Tokyo University, 1874–1900. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18 (2):100–120, December 2008. CODEN HISCDU. ISSN 0285-4821.

Kimoto:1982:BAT

[Kim82] Tadaaki Kimoto. Die Begriffswandlung der allgemeinen Technologie und die historische Bedeutung der Technologie Johann Beckmanns. (German) [The conversion of general technology and the historical importance of Johann Beckmann’s technology]. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):21–44, 1982. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837286.

Kimoto:1995:TDJ

[Kim95] Tadaaki Kimoto. Technological development in Japan and its historical research since World War II. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):89–102, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Kim:1997:TDI

[Kim97] Dong-Won Kim. Two different influences on the Japanese physics community in the Early Twentieth Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):125–136, December 1997. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Kimoto:1998:OTY

[Kim98] Tadaaki Kimoto. Obituary: Toshio Yamazaki and history of technology in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):227–234, March 1998. CODEN HISCDU. ISSN 0285-4821.

Kimoto:2003:EIT

[Kim03] Tadaaki Kimoto. European imperialism and technological development in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3): 175–182, March 2003. CODEN HISCDU. ISSN 0285-4821.

Kim:2008:TPG

[Kim08a] Boumsoung Kim. Terrains of practice: Geophysical investigations in Japan in the 1880s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2): 121–131, December 2008. CODEN HISCDU. ISSN 0285-4821.

Kim:2008:LSB

[Kim08b] Dong-Won Kim. The longer the stay, the better the result?: Nishina Yoshio’s overseas study in the 1920s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2):151–171, December 2008. CODEN HISCDU. ISSN 0285-4821.

Kim:2008:PEM

[Kim08c] Yung Sik Kim. Problem of Early Modern Japan in the history of science in East Asia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(1): 49–57, July 2008. CODEN HISCDU. ISSN 0285-4821.

Kim:2009:ISI

[Kim09a] Dong-Won Kim. Imaginary savior: the image of the nuclear bomb in Korea, 1945–1960. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2): 105–118, December 2009. CODEN HISCDU. ISSN 0285-4821.

Kim:2009:TTB

[Kim09b] Seong-Jun Kim. Technology transfer behind a diplomatic struggle: Reappraisal of South Korea’s nuclear fuel project in the 1970s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):184–193, December 2009. CODEN HISCDU. ISSN 0285-4821.

Kim:2010:HPL

[Kim10] Sungkhun Kim. How physical laws were understood in Mid-19th Century East Asia: A comparative study of Choe Han-gi and Nishi Amane. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(1):1–20, July 2010. CODEN HISCDU. ISSN 0285-4821.

Kim:2011:DRE

[Kim11] Boumsoung Kim. Detecting, recording and expanding: Instrumentation of earthquake and tsunami observations in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(3):179–195, March 2011. CODEN HISCDU. ISSN 0285-4821.

Kiselev:2006:RMS

[Kis06] G. V. Kiselev. Role of Moscow State University alumni in the Soviet Atomic Project. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):179–205, November 2006. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110006448390.

Konuma:1989:RIR

[KMT89] Michiji Konuma, Chieko Masuzawa, and Yoshio Takada. Resumption of international relationship of Japanese particle physicists after World War II. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):23–41, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435090;   https://ci.nii.ac.jp/naid/110009837387.

Knorr:1985:ETB

[Kno85] Wilbur R. Knorr. Euclid’s tenth book: An analytic survey. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):17–35, September 1985. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/

110009837339. Special Issue, “The Historiography of the History of Mathematics”.

Kobayashi:2002:WKM

[Kob02] Tatsuhiko Kobayashi. What kind of mathematics and terminology was transmitted into 18th-century Japan from China? Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):1–17, July 2002. CODEN HISCDU. ISSN 0285-4821.

Kojima:2000:AQT

[Koj00] Chieko Kojima. Acceptance of quantum theory in France. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):156–162, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Kojima:2010:BRS

[Koj10] Chieko Kojima. Book review: Shizue Hinokawa, Saikurotoron kara Genbaku e: Kaku Jidai no Kigen o Saguru (From Cyclotron to Atomic Bomb: Searching for the Origin of the Atomic Age). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(3):235–238, March 2010. CODEN HISCDU. ISSN 0285-4821.

Kojevnikov:2013:SCP

[Koj13] Alexei Kojevnikov. Science as co-producer of Soviet polity. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):161–180, March 2013. CODEN HISCDU. ISSN 0285-4821.

Kolchinsky:2006:SMS

[Kol06] Eduard I. Kolchinsky. Science mobilization in the Soviet Union. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):15–28, July 2006. CODEN HISCDU. ISSN 0285-4821.

Konno:1981:SRI

[Kon81] Hiroyuki Konno. Some remarks on the interference experiments with faint light. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):95–105, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837257.

Konno:1983:BRT

[Kon83a] Hiroyuki Konno. Book review: Tetu Hirosige, Hirosige Tetu Kagakushi Ronbunshū (Collected Papers of the History of Science of Tetu Hirosige), edited and commented by S. Nisio, 2 vols., Misuzushobo, Tokyo, vol. 1, 1980, iii + 365 pp., Yen 4500, vol. 2, 1981, iv + 378 pp., Yen 5800. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):111–114, March 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837301.

Konno:1983:SEG

[Kon83b] Hiroyuki Konno. Slater’s evidence for the genesis of the Bohr– Kramers–Slater theory. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):39–52, September 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.

nii.ac.jp/naid/110009837441.

Konno:1992:NSE

[Kon92] Hiroyuki Konno. Notes: Spontaneous emission, radiation force, and the Correspondence Principle. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2 (2):151–158, December 1992. CODEN HISCDU. ISSN 0285-4821.

Konno:1997:JCS

[Kon97] Hiroyuki Konno. J. C. Slater’s optical theory in 1925: another dispersion theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):59–64, June 1997. CODEN HISCDU. ISSN 0285-4821.

Konno:1998:STL

[Kon98] Hiroyuki Konno. Slater’s theory of line breadth and the virtual oscillator model: A premature attempt in the old quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):213–226, March 1998. CO-DEN HISCDU. ISSN 0285-4821.

Konno:2000:BSQ

[Kon00] Hiroyuki Konno. Bohr’s search for the quantum theory of dispersion: the number of dispersion electrons, absorption and emission of light and the oscillator model. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10 (2):163–176, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Konagaya:2002:MPR

[Kon02] Daisuke Konagaya. The methodology of Planck’s radiation theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):43–58, July 2002. CODEN HISCDU. ISSN 0285-4821.

Konagaya:2010:SDP

[Kon10] Daisuke Konagaya. Success from different programs: The development of experimental researches on thermal radiation in Germany at the end of the 19th Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20 (2):63–95, December 2010. CODEN HISCDU. ISSN 0285-4821.

Kota:2001:RCH

[Kot01] Shoji Kota. Riemann’s “Commentatio” and his geometrical thought. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):235–247, March 2001. CODEN HISCDU. ISSN 0285-4821.

Kota:2006:CDG

[Kot06] Shoji Kota. The contact of differential geometry and continuous groups of transformations: W. Killing’s “Ueber die Grundlagen der Geometrie’. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):166–178, November 2006. CODEN HISCDU. ISSN 0285-4821.

Kranzberg:1982:HTU

[Kra82] Melvin Kranzberg. The history of technology in the United States. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):1–14, 1982. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837284.

Kuriyama:2000:BRY

[Kur00] Shigehisa Kuriyama. Book review: Yamada Keiji, Chūgoku Igaku no Kigen (Origins of Chinese Medicine) (Tokyo: Iwanami Shoten, 1999, xx + 496 pp. 9500 Yen). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10 (1):92–95, July 2000. CODEN HISCDU. ISSN 0285-4821.

Kusuba:1981:BRI

[Kus81a] Takanori Kusuba. Book review: Indo Tenmongaku Sugaku Shu (A Collection of Hindu Astronomical and Mathematical Works), Japanese Translation with Introductions and Notes by Michio Yano, Takao Hayashi and Yasuke Ikari, Asahi-shuppan, Tokyo, 1980, 517 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):130–132, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837261.

Kusuba:1981:BST

[Kus81b] Takanori Kusuba. Brahmagupta’s sutras on tri- and quadrilaterals. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):43–55, 1981. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837265.

Kusukawa:2014:AOE

[Kus14] Sachiko Kusukawa. Aligning observations in Edward Tyson’s ‘Lumbricus Latus’ (1683). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(3): 167–190, March 2014. CODEN HISCDU. ISSN 0285-4821.

Kawahara:1996:JCH

[KY96] Hideki Kawahara and Michio Yano. Japanese contributions to the history of Chinese science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6 (2):123–158, December 1996. CODEN HISCDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Lamy:2009:ADC

[Lam09] Jérôme Lamy. Antoine Darquier et la cosmologie de Johann Heinrich Lambert: emprunts et distances. (French) [Antoine Darquier and the cosmology of Johann Heinrich Lambert: borrowings and distances]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1):43–54, July 2009. CODEN HISCDU. ISSN 0285-4821.

Lapteff:2008:CTC

[Lap08] Sergey Lapteff. A comparison of technological characteristics of ancient bronze mirrors in China, Korea, Japan, and South Siberia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):242–252, March 2008. CODEN HISCDU. ISSN 0285-4821.

Lee:2006:ICS

[Lee06] Sung Kyu Lee. Imperialism and colonial science: A case study of Keijo Imperial University, 1924–1945. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):247–267, March 2006. CODEN HISCDU. ISSN 0285-4821.

Lee:2008:HGM

[Lee08] Jong-Chan Lee. Hygienic governance and military hygiene in the making of Imperial Japan, 1868–1912. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(1):1–23, July 2008. CODEN HISCDU. ISSN 0285-4821.

Leslie:2006:ECW

[Les06] Stuart W. Leslie. Engineering the Cold War. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):35–40, July 2006. CODEN HISCDU. ISSN 0285-4821.

Lee:2012:TPK

[LH12] Young-Mi Lee and Sungook Hong. Technoscience and politics in Korea in the 1970s: “scientification of all people movement” and the “Saemaul technical service corps”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(3):174–192, March 2012. CODEN HISCDU. ISSN 0285-4821.

Liang:2006:TST

[Lia06] Bo Liang. The transmission of science and technology from Japan to China before the Second World War. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):233–246, March 2006. CODEN HISCDU. ISSN 0285-4821.

Lightman:2016:RPV

[Lig16] Bernard Lightman. The royal panopticon: Victorian museums and the popularization of science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25 (3):164–189, March 2016. CODEN HISCDU. ISSN 0285-4821.

Lim:2008:LCS

[Lim08] Jongtae Lim. Locating a center on the surface of a globe: Negotiating China’s position on the spherical Earth in Seventeenth and Eighteenth-Century China and Korea. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):175–188, March 2008. CODEN HISCDU. ISSN 0285-4821.

Liu:1994:YHM

[Liu94] Dun Liu. 400 years of the history of mathematics in China — an introduction to the major historians of mathematics since 1592. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):103–111, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Lochak:2000:OQS

[Loc00] Georges Lochak. Obituary: Quelques souvenirs d’un ami disparu: Dr. Takehiko Takabayashi (1919–1999). (French) [Obituary: Some memories of a missing friend: Dr. Takehiko Takabayashi (1919– 1999)]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):211–212, November 2000. CODEN HISCDU. ISSN 0285-4821.

Loh:2017:RPP

[Loh17] Shi-Lin Loh. Radiation in print: Popularizing X-rays in the mass media of early twentieth-century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(2):93–111, February 2017. CODEN HISCDU. ISSN 0285-4821.

Lorenz:1986:BCW

[Lor86] Martina Lorenz. Der Beitrag Christian Wolffs zur Rezeption von Grundprinzipien der Mechanik Newtons in Deutschland zu Beginn des 18. Jahrhunderts. (German) [Christian Wolff’s contribution to the reception of fundamental principles of Newton’s mechanics in the early 18th century]. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):87–100, September 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435074;
https://ci.nii.ac.jp/naid/110009837355.

Low:1989:ASI

[Low89a] Morris Fraser Low. Accounting for science: The impact of social and political factors on Japanese elementary particle physics. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):43–65, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435091;
https://ci.nii.ac.jp/naid/110009837388.

Low:1989:HJP

[Low89b] Morris Fraser Low. The history of Japanese physics and the rise of the scientist–historian: Proceedings of the Japan–USA Collaborative Workshops on the History of Particle Theory in Japan, 1935– 1960, Japan–USA Collaboration, Second Phase. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):117–120, 1989. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837392.

Low:2008:ISP

[Low08] Morris Low. Introduction: Science, Protestant Christianity and Darwinism in Meiji Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2): 76–87, December 2008. CODEN HISCDU. ISSN 0285-4821.

Low:2010:CMN

[Low10] Morris Low. Colonial modernity and networks in the Japanese empire: the role of Got Shinpei. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19 (3):195–208, March 2010. CODEN HISCDU. ISSN 0285-4821.

Low:2011:IWS

[Low11] Morris Low. The impact of Western science and technology on ‘ukiyo-e’ prints and book illustrations in Late Eighteenth and Nineteenth Century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(1): 66–87, July 2011. CODEN HISCDU. ISSN 0285-4821.

Low:2014:MSN

[Low14] Morris Low. In memoriam: Shigeru Nakayama (1928–2014). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(1):25–28, August 2014. CODEN HISCDU. ISSN 0285-4821.

Luo:2003:DFS

[Luo03] Jianjin Luo. On the development of the formulae for sums of powers of integers. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(2):101–110, November 2003. CODEN HISCDU. ISSN 0285-4821.

Luo:2015:AAT

[Luo15] Jianjin Luo. An algorithm analysis on the twelve tones in the book bamboo slips of Fangmatan of the Qin Dynasty in City Tianshui. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(2):50–58, February 2015. CODEN HISCDU. ISSN 0285-4821.

Li:1999:MEB

[LXF99] Wenlin Li, Zelin Xu, and Lisheng Feng. Mathematical exchanges between China and Korea. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9 (1):73–83, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français.

Low:1989:BPA

[LY89] Morris F. Low and Hitoshi Yoshioka. Buying the ’peaceful atom’: The development of nuclear power in Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):29–44, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435102;
https://ci.nii.ac.jp/naid/110009837406.

Maas:1991:DCP

[Maa91] Joerg F. Maas. The dissolution of contradictory patterns of explication: method and the interdependencies between logic and mathematics in the 16th and early 17th centuries. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):63–70, June 1991. CODEN HISCDU. ISSN 0285-4821.

Maat:2007:ASC

[Maa07] Harro Maat. Agricultural sciences in Colonial Indonesia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(3):244–263, March 2007. CODEN HIS-CDU. ISSN 0285-4821.

Maier:2006:KWS

[Mai06] Helmut Maier. The Kaiser Wilhelm Society during the Second World War. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):5–14, July 2006. CODEN HISCDU. ISSN 0285-4821.

Maki:1989:DEP

[Mak89] Ziro Maki. The development of elementary particle theory in Japan: Methodological aspects of the formation of the Sakata and Nagoya models. Historia Scientiarum = International journal of the History of Science Society of Japan, 36(??):83–95, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435093;   https://ci.nii.ac.jp/naid/110009837390.

Mak:2015:TBA

[Mak15] Bill M. Mak. The transmission of Buddhist astral science from India to East Asia: The central Asian connection. Historia Scientiarum.

Second Series. International Journal of the History of Science Society of Japan, 24(2):59–75, February 2015. CODEN HISCDU. ISSN 0285-4821.

Malloy:2009:CSN

[Mal09] Sean L. Mallo. Contemporary scholarship and new light on the Abomb decision. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):91–104, December 2009. CODEN HISCDU. ISSN 0285-4821.

Martzloff:1981:GES

[Mar81] Jean Claude Martzloff. La géometrie euclidienne selon Mei Wending. (French) [Euclidean geometry according to Mei Wending]. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):27–42, 1981. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837264.

Martzloff:1985:AHM

[Mar85] Jean Claude Martzloff. Aperçu sur l’histoire des mathématiques chinoises telle qu’elle est pratiquée en République Populaire de Chine. (French) [Overview of the history of Chinese mathematics as practiced in the People’s Republic of China]. Historia Scientiarum = International journal of the History of Science Society of Japan, 28 (??):1–30, March 1985. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837326.

Matsuo:1983:DAS

[Mat83] Yukitoshi Matsuo. “A digression of the animal spirits”: the changes of the concept of ‘spirit’ in the seventeenth-century science (I). Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):1–15, September 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837438.

Matsumoto:1989:BRC

[Mat89] Miwao Matsumoto. Book review: Chikatoshi Kamatani, Gijutsu Taikoku Hyakunen no Kei: Nippon no Kindaika to Kokuritsu Kenkyu Kikan (The Road to Techno-Nationalism: Japanese Modernization and National Research Institutes from the Meiji Era), Heibonsha, Tokyo, 1988, 259 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??): 75–80, 1989. CODEN HISCDU. ISSN 0285-4821. URL https://

ci.nii.ac.jp/naid/110009837410.

Matsumoto:1997:IJN

[Mat97] Miwao Matsumoto. The Imperial Japanese Navy’s connection with a marine steam turbine transfer from the West. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3):209–228, March 1997. CODEN HISCDU. ISSN 0285-4821.

Matsubara:1998:EJS

[Mat98] Yōko Matsubara. The enactment of Japan’s sterilization laws in the 1940s: A prelude to postwar eugenic policy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):187–202, December 1998. CODEN HISCDU. ISSN 0285-4821.

Matsunaga:2002:EET

[Mat02] Toshio Matsunaga. Evolutionism in Early Twentieth Century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(3):218–225, March 2002. CODEN HISCDU. ISSN 0285-4821.

Matsubara:2004:RMJ

[Mat04a] Yoko Matsubara. The reception of Mendelism in Japan, 1900–1920. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):232–240, March 2004. CODEN HISCDU. ISSN 0285-4821.

Matsumura:2004:SBT

[Mat04b] Noriaki Matsumura. The similarity between two medicines: Oriental medical thought of meridian and Western medical knowledge of nerve. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):224–231, March 2004. CODEN HISCDU. ISSN 0285-4821.

Maula:1985:PHH

[Mau85] Erkka Maula. Proof, history and heuristics. Historia Scientiarum = International journal of the History of Science Society of Japan, 29 (??):125–155, September 1985. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837344. Special Issue,

“The Historiography of the History of Mathematics”.

Meade:2017:PSP

[Mea17] R. Meade. Popular science & personal endeavor in early-Meiji Japan: The case of Hatsumei Kiji. Historia Scientiarum. Second Series.

International Journal of the History of Science Society of Japan, 26 (2):77–92, February 2017. CODEN HISCDU. ISSN 0285-4821.

Metailie:2002:CSI

[Mét02] Georges Métailié. Comparative study of the introduction of modern botany in Japan and China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11 (3):205–217, March 2002. CODEN HISCDU. ISSN 0285-4821.

Matsumura:1998:FYP

[MHM98] Noriaki Matsumura, Yoshiyuki Hirono, and Yōko Matsubara. Fujikawa Yū, pioneer of the history of medicine in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):157–172, December 1998. CODEN HISCDU. ISSN 0285-4821.

Mihara:2015:VNM

[Mih15] Sayaka Mihara. Vitalism and neonatal medicine in Japan, 1901– 1945. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(2):119–139, December 2015. CODEN HISCDU. ISSN 0285-4821.

Miura:1981:ALA

[Miu81] Nobuo Miura. The algebra in the Liber Abaci of Leonardo Pisano. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):57–65, 1981. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837266.

Miura:1986:ATP

[Miu86] Nobuo Miura. The applications of trigonometry in Pitiscus: a preliminary essay. Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):63–78, March 1986. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435069;   https://ci.nii.ac.jp/
naid/110009837348.

Miura:1987:BRD

[Miu87] Nobuo Miura. Book review: Donald R. HILL (Translator), The Book of Ingenious Devices (Kitab al-Hiyal) by the Banu Musa bin Shakir, Translated and annotated by D. R. Hill, Dordrecht, 1979, x + 267 pages. Historia Scientiarum = International journal of the History of Science Society of Japan, 32(??):143–145, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837362.

Miura:1988:CBP

[Miu88] Nobuo Miura. Charles de Bovelles and perfect numbers. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):1–10, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435080;
https://ci.nii.ac.jp/naid/110009837372.

Miura:1989:ALT

[Miu89] Nobuo Miura. The applications of logarithms to trigonometry in Richard Norwood. Historia Scientiarum = International journal of the History of Science Society of Japan, 37(??):17–30, 1989. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435096;   https://ci.nii.ac.jp/naid/110009837396.

Miura:1990:FFR

[Miu90] Motohiro Miura. The Firth of Forth railway bridge and a Japanese engineer, Kaichi Watanabe. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):53–57, 1990. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837423.

Miura:2003:BRS

[Miu03] Nobuo Miura. Book review: Sasaki Chikara, Dekaruto no Sugaku-Shiso (Descartes’s Mathematical Thought). University of Tokyo Press, 2003, xvi + 600 pp., ISBN: 4-13-061351-0. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):64–68, July 2003. CODEN HISCDU. ISSN 0285-4821.

Miyake:1999:DFS

[Miy99] Katsuya Miyake. Did Fermat see these structures? Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):37–47, July 1999. CODEN HIS-CDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Miyagawa:2008:MOS

[Miy08a] Takuya Miyagawa. The meteorological observation system and colonial meteorology in Early 20th-Century Korea. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2):140–150, December 2008. CODEN HIS-CDU. ISSN 0285-4821.

Miyajima:2008:NDK

[Miy08b] Kazuhiko Miyajima. A new discovery of Korean astrolabe. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):211–224, March 2008. CODEN HISCDU. ISSN 0285-4821.

Mizoguchi:1998:JBN

[Miz98] Hazime Mizoguchi. Japanese biologists at the Naples Zoological Station, 1887–1956. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):99–114, December 1998. CODEN HISCDU. ISSN 0285-4821.

Mizoguchi:2007:JBM

[Miz07] Hajime Mizoguchi. Japanese biologists at the Marine Biological Laboratory, Woods Hole. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(1): 20–37, July 2007. CODEN HISCDU. ISSN 0285-4821.

Moon:2008:LCE

[Moo08] Joong-Yang Moon. Late circulation of the Early Qing natural studies in 19th Century Korea. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17 (3):189–210, March 2008. CODEN HISCDU. ISSN 0285-4821.

Moon:2012:DGR

[Moo12] Manyong Moon. The dual green revolutions in South Korea: Reforestation and agricultural revolution under the authoritarian regime. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(3):161–173, March 2012. CODEN HISCDU. ISSN 0285-4821.

Morimura:2003:RSG

[Mor03] K. Morimura. The role of special group article in ancient Chinese medical prescription. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1): 1–12, July 2003. CODEN HISCDU. ISSN 0285-4821.

Mori:2004:HYV

[Mor04a] Shuichi Mori. The history of Yunosawa village and the leprosy policy in Japan: a study of a free medical-treatment area for leprosy patients. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(2):137–153, November 2004. CODEN HISCDU. ISSN 0285-4821.

Mori:2004:BRH

[Mor04b] Yukiya Mori. Book review: Y. Hirono, Y. Ichinokawa, and M. Hayashi (eds.), Seimei-Kagaku no Kingendai-Shi (The Modern History of Life Sciences). Keisoshobo, 2002, xx + 376 + xviii pp., ISBN: 4-326-15366-0. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(1):77–82, July 2004. CODEN HISCDU. ISSN 0285-4821.

Morita:2016:ECQ

[Mor16] Kunihisa Morita. Einstein’s criticism of quantum mechanics and Humean philosophy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1): 65–74, March 2016. CODEN HISCDU. ISSN 0285-4821.

Motoyama:1998:DSB

[Mot98] Shozo Motoyama. Diffusion of science in Brazil — the beginning of FAPESP. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):165–172, March 1998. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Motoyama:2001:OJP

[Mot01] Shozo Motoyama. Obituary: A Japanese physicist in Brazil: Prof. Taketani (1911–2000) and the Brazilian culture. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):112–126, July 2001. CODEN HISCDU. ISSN 0285-4821.

Morris-Suzuki:1995:GTT

[MS95] Tessa Morris-Suzuki. The great translation: Traditional and modern science in Japan’s industrialisation. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):103–116, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Matsuhashi:2003:TDJ

[MS03] Toru Matsuhashi and Shigeo Sugiyama. The technological development of the Japanese plastic industry shortly after WW II and the introduction of Western high technology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3):219–232, March 2003. CODEN HISCDU. ISSN 0285-4821.

Michiwaki:1982:CMW

[MT82] Yoshimasa Michiwaki and Osamu Takinami. On the comparison of mathematicians’ ways of thinking in Japan: Old and new. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):117–121, 1982. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837279.

Murata:1980:WAI

[Mur80] Tamotsu Murata. Wallis’ Arithmetica Infinitorum and Takebe’s Tetsujutsu Sankei: What underlies their similarities and dissimilarities? Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):77–100, 1980. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837247.

Murata:1984:BRN

[Mur84] Tamotsu Murata. Book review: Nippon no Sugaku 100 nen Shi (History of mathematics in Japan during the last 100 years), Vol. 1, 1983, xxi + 337 p., Vol. 2, 1984, ix + 394 p., Iwanami Shoten, Tokyo, 1983–84. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):119–122, September 1984. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837325.

Murata:1986:NTE

[Mur86] Tamotsu Murata. La naissance de la théorie des ensembles et son retentissement en France. (French) [The birth of set theory and its repercussion in France]. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):1–24, September 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435071;
https://ci.nii.ac.jp/naid/110009837352.

Murata:1987:CAJ

[Mur87a] Tamotsu Murata. Certain aspects of Japanese studies on the history of mathematics. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):43–59, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837367.

Murata:1987:LE

[Mur87b] Tamotsu Murata. A letter from the Editor. Historia Scientiarum = International journal of the History of Science Society of Japan, 33 (??):1–2, 1987. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110009837364.

Muroi:1988:IPB

[Mur88] Kazuo Muroi. Inheritance problems of Babylonian mathematics. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):11–19, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435081;
https://ci.nii.ac.jp/naid/110009837373.

Muroi:1990:ICB

[Mur90] Kazuo Muroi. Interest calculation in Babylonian mathematics: New interpretations of VAT 8521 and VAT 8528: To the memory of O. Neugebauer. Historia Scientiarum = International journal of the History of Science Society of Japan, 39(??):29–34, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435105;   https://ci.nii.ac.jp/naid/110009837413.

Muroi:1991:EZS

[Mur91] Kazuo Muroi. The expressions of zero and of squaring in the Babylonian mathematical text VAT 7537. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):59–62, June 1991. CODEN HISCDU. ISSN 0285-4821.

Murata:1992:QRL

[Mur92a] Tamotsu Murata. Quelques remarques sur le Livre X des Eléments d’Euclide. (French) [Some remarks on Book X of Euclid’s Elements]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):51–60, June 1992. CODEN HISCDU. ISSN 0285-4821.

Muroi:1992:RSM

[Mur92b] Kazuo Muroi. Reexamination of Susa mathematical text no. 3: alleged value π ≈ 3(1/8). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1): 45–49, June 1992. CODEN HISCDU. ISSN 0285-4821.

Muroi:1992:SCP

[Mur92c] Kazuo Muroi. Small canal problems of Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):173–180, March 1992. CO-DEN HISCDU. ISSN 0285-4821.

Muroi:1994:RFP

[Mur94] Kazuo Muroi. Reexamination of the first problem of the Susa mathematical text no. 9. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3): 231–233, March 1994. CODEN HISCDU. ISSN 0285-4821.

Muroi:1996:THP

[Mur96] Kazuo Muroi. Two harvest problems of Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(3):249–254, March 1996. CO-DEN HISCDU. ISSN 0285-4821.

Muroi:1997:ESO

[Mur97] Kazuo Muroi. An enigmatic sentence in the old Babylonian table of exponents and logarithms. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3):229–230, March 1997. CODEN HISCDU. ISSN 0285-4821. March.

Muroi:1998:EOB

[Mur98] Kazuo Muroi. Early Old Babylonian mathematical problems written in Sumerian. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):199–203, March 1998. CODEN HISCDU. ISSN 0285-4821.

Muroi:1999:ESR

[Mur99] Kazuo Muroi. Extraction of square roots in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(2):127–133, November 1999. CODEN HISCDU. ISSN 0285-4821.

Muroi:2001:IPS

[Mur01] Kazuo Muroi. Inheritance problems in the Susa mathematical text no. 26. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):226–234, March 2001. CODEN HISCDU. ISSN 0285-4821. See note [Høy03].

Muroi:2002:EMB

[Mur02] Kazuo Muroi. Expressions of multiplication in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(2):115–120, November 2002. CODEN HISCDU. ISSN 0285-4821.

Muroi:2003:ESB

[Mur03a] Kazuo Muroi. Expressions of squaring in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(2):111–119, November 2003. CODEN HISCDU. ISSN 0285-4821.

Muroi:2003:MTT

[Mur03b] Kazuo Muroi. Mathematical term takīltum and completing the square in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3):254–263, March 2003. CODEN HISCDU. ISSN 0285-4821.

Murakami:2005:OMW

[Mur05a] Yoichiro Murakami. Obituary: Masao Watanabe (1920–2005). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):205–208, November 2005. CODEN HISCDU. ISSN 0285-4821.

Muroi:2005:RBM

[Mur05b] Kazuo Muroi. A reexamination of the Babylonian mathematical text Erm 15073: a new interpretation of the earthwork problems. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):191–200, November 2005. CODEN HISCDU. ISSN 0285-4821.

Marabini:2018:CMN

[MV18] S. Marabini and G. B. Vai. Christian missionaries and natural things: The Italian-style geological collection of Cimatti’s museum at Chofu, Tokyo, Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27 (3):334–352, 2018. CODEN HISCDU. ISSN 0285-4821.

Nagaoka:1980:BRA

[Nag80] Ryosuke Nagaoka. Book review: Akira Kobori, Juhasseiki no Sugaku (Mathematics in the 18th Century), Kyoritsu-shuppan, Tokyo, 1979, 200 pp., Yen 3,300. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):114–115, 1980. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837251.

Nagazumi:1983:DIS

[Nag83] Akira Nagazumi. The diffusion of the idea of social Darwinism in East and Southeast Asia. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):1–18, March 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110009837295.

Nagamini:1998:LFC

[Nag98] Marilda Nagamini. L’histoire de la FAPESP: Un chapître dans le développement de la science et de la technologie au Brésil. (French) [The story of the FAPESP: A chapter in the development of science and technology in Brazil]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3): 173–180, March 1998. CODEN HISCDU. ISSN 0285-4821.

Nagahira:2011:PAC

[Nag11] Yukio Nagahira. The physical apparatus collection at Kyoto University: Historical material as thing knowledge. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(3):212–219, March 2011. CODEN HISCDU. ISSN 0285-4821.

Nakayama:1980:TAT

[Nak80] Shigeru Nakayama. Two academic traditions in comparison: Western and Chinese. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):47–75, 1980. CO-DEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837246.

Nakajima:1983:BRM

[Nak83] Hideto Nakajima. Book review: Masakazu Yoshinaka, Nyūton Rikigaku no Tanja (The Birth of Newton’s Mechanics), Saiensu Sha, Tokyo, 1982, 243 pp., Yen 1700. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??): 118–119, March 1983. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837303.

Nakagawa:1984:DEP

[Nak84] Yasuo Nakagawa. The development of early practical electromagnetic telegraphs and the mechanization of skilled operation. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):77–89, September 1984. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837321.

Nakayama:1985:BRM

[Nak85] Shigeru Nakayama. Book review: Masakazu Yoshinaka, Kagakusha no Hasso (The Way of Thinking of Scientists), Tamagawa U. P., 1984. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):159–160, September 1985. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/

110009837346. Special Issue, “The Historiography of the History of Mathematics”.

Nakayama:1986:TSD

[Nak86] Shigeru Nakayama. The three stage development of knowledge and the media [in Japanese]. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):101–113, September 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435075;
https://ci.nii.ac.jp/naid/110009837356.

Nakane:1991:RTB

[Nak91] Michiyo Nakane. The role of the three-body problem in W. R. Hamilton’s construction of the characteristic function for mechanics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):27–38, June 1991. CODEN HISCDU. ISSN 0285-4821.

Nakamura:1994:SSC

[Nak94a] Kunimitsu Nakamura. On the sprout and setback of the concept of mathematical “proof” in the Edo period in Japan: regarding the method of calculating number π. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):185–199, March 1994. CODEN HISCDU. ISSN 0285-4821.

Nakata:1994:JPM

[Nak94b] Ryoichi Nakata. Joseph Privat de Molières: reconciler between Cartesianism and Newtonianism in collision theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(3):201–213, March 1994. CODEN HISCDU. ISSN 0285-4821.

Nakamura:1997:PTD

[Nak97] Kunimitsu Nakamura. Process through disaffirmance of “material theory of heat” to introduction and diffusion of “kinetic theory of heat” in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3):187–208, March 1997. CODEN HISCDU. ISSN 0285-4821. March.

Nakata:1998:NNE

[Nak98] Ryoichi Nakata. Non-Newtonian elements in French Newtonian physics: A perspective on a history of science of 18th Century France. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):205–212, March 1998. CODEN HISCDU. ISSN 0285-4821.

Nakane:2000:BRC

[Nak00a] Michiyo Nakane. Book review: Carl Gustav J. Jacobi, Vorlesungen über Analytische Mechanik, Berlin 1847/48. Dokumente zur Geschichte der Mathematik, Band 8. Edited by Helmut Pulte. Deutsche Mathematiker-Vereinigung (Braunschweig/Wiesbaden: Friedr. Vieweg & Sohn Verlagsgesellschaft, 1996, lxvii + 353 pp., ISBN 3-528-06692-X). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1): 96–99, July 2000. CODEN HISCDU. ISSN 0285-4821.

Nakata:2000:DSR

[Nak00b] Ryoichi Nakata. D’Alembert’s second resolution in Recherches sur la précession des equinoxes: comparison with Euler. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):58–76, July 2000. CODEN HISCDU. ISSN 0285-4821.

Nakata:2002:GPR

[Nak02] Ryoichi Nakata. The general principles for resolving mechanical problems in d’Alembert, Clairaut and Euler. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):18–42, July 2002. CODEN HISCDU. ISSN 0285-4821.

Nakayama:2005:CSI

[Nak05a] Shigeru Nakayama. Colonial science: An introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):93–96, November 2005. CODEN HISCDU. ISSN 0285-4821.

Nakayama:2005:OMY

[Nak05b] Shigeru Nakayama. Obituary: Mitsutomo Yuasa (1909–2005). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):201–204, November 2005. CODEN HISCDU. ISSN 0285-4821.

Nakayama:2007:BRT

[Nak07a] Shigeru Nakayama. Book review: Tetsuro Nakaoka, Nihon Kindai Gijutsu no Keisei: ‘Dento’ to ‘Kindai’ no Dainamikusu (The formation of Modern Japanese Technology: A Dynamics of ‘Tradition’ and ‘Modernity’), 2006. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2): 153–154, December 2007. CODEN HISCDU. ISSN 0285-4821.

Nakayama:2007:TKH

[Nak07b] Shigeru Nakayama. Thomas Kuhn — a historian’s personal recollections. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(1):49–53, July 2007. CODEN HISCDU. ISSN 0285-4821.

Nakao:2009:IAB

[Nak09] Maika Nakao. The image of the atomic bomb in Japan before Hiroshima. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):119–131, December 2009. CODEN HISCDU. ISSN 0285-4821.

Nakayama:2011:MFS

[Nak11a] Shigeru Nakayama. In memorial of Fumihiko Satofuka. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(1):88–89, July 2011. CODEN HISCDU. ISSN 0285-4821.

Nakazawa:2011:WAR

[Nak11b] Satoshi Nakazawa. Whirling arm and its role in experimental research on fluid resistance in the Eighteenth Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(3):167–178, March 2011. CODEN HIS-CDU. ISSN 0285-4821.

Nakazawa:2013:MMP

[Nak13] Satoshi Nakazawa. Mixed mathematics, physico-mathematics, and the legitimacy of experience: A case study from eighteenth-century fluid-resistance research. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23 (2):141–163, December 2013. CODEN HISCDU. ISSN 0285-4821.

Nakamura:2015:IWT

[Nak15] Eri Nakamura. “Invisible” war trauma in Japan: Medicine, society, and military psychiatric casualties. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(2):144–??, December 2015. CODEN HISCDU. ISSN 0285-4821.

Nakagawa:2018:FPV

[Nak18] T. Nakagawa. Frontier of professionals: The vision of American advisors that supported Hokkaido development in the early Meiji period. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(3):300–318, 2018. CODEN HISCDU. ISSN 0285-4821.

Navarro:1991:ESM

[Nav91] Luis Navarro. On Einstein’s statistical-mechanical approach to the early quantum theory (1904–1916). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):39–58, June 1991. CODEN HISCDU. ISSN 0285-4821.

Needham:1980:GKF

[Nee80] Joseph Needham. The guns of Khaifeng-fu. Historia Scientiarum = International journal of the History of Science Society of Japan, 19 (??):11–30, 1980. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837244.

Nha:1988:CSM

[Nha88] Il-Seong Nha. Calibration of the stellar magnitude scale prior to Pogson. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):39–44, 1988. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435087;   https://ci.nii.ac.jp/naid/110009837381.

Nisio:1992:SHL

[Nis92] Sigeko Nisio. Satio Hayakawa’s letter to Hideki Yukawa (19 Sept. 1955) — a trigger for controlled nuclear fusion research in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):221–222, March 1992. CO-DEN HISCDU. ISSN 0285-4821. See errata [Ano92f].

Nishiyama:2003:DLA

[Nis03] Takashi Nishiyama. Aeronautical technology for pilot safety: Reexamining deck-landing aircraft in Great Britain, Japan, and the United States. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):13–32, July 2003. CODEN HISCDU. ISSN 0285-4821.

Nakao:2015:YNR

[NKY15] Maika Nakao, Takeshi Kurihara, and Masakatsu Yamazaki. Yasushi Nishiwaki, radiation biophysics, and peril and hope in the nuclear age. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(1):8–35, August 2015. CODEN HISCDU. ISSN 0285-4821.

Nakamura:2002:CFU

[NM02] Kunimitsu Nakamura and Tsutomu Matsumoto. Characteristics of farmers in the upland farming regions (sericulture regions) in Eastern Japan as seen in the regions where Wasan mathematicians were active in the Late Edo Period. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12 (2):100–114, November 2002. CODEN HISCDU. ISSN 0285-4821.

Nafai:2009:SRR

[NNSA09] Hiroyuki Nagai, Yoshio Nukaga, Koji Saeki, and Akira Akabayashi. Self-regulation of recombinant DNA technology in Japan in the 1970s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1):1–18, July 2009. CODEN HISCDU. ISSN 0285-4821.

Nagase-Reimer:2005:NWR

[NRGY05] Keiko Nagase-Reimer, Walter Grunden, and Masakatsu Yamazaki. Nuclear weapons research in Japan during the Second World War. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(3):201–240, March 2005. CODEN HISCDU. ISSN 0285-4821.

Nugaev:1988:SRS

[Nug88] R. N. Nugaev. Special relativity as a stage in the development of quantum theory: a new outlook of scientific revolution. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):57–79, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435084;
https://ci.nii.ac.jp/naid/110009837376.

Nakazawa:2009:I

[NY09] Shiho Nakazawa and Masakatsu Yamazaki. Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):87–90, December 2009. CODEN HISCDU. ISSN 0285-4821.

Obayashi:1992:SMO

[Oba92] Masayuki Obayashi. Stent’s myth in the origins of molecular biology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):139–150, December 1992. CODEN HISCDU. ISSN 0285-4821.

Oberlander:2004:RSM

[Obe04] Christian Oberländer. The rise of scientific medicine in Japan: Beriberi as the driving force in the quest for specific causes and the introduction of bacteriology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13 (3):176–199, March 2004. CODEN HISCDU. ISSN 0285-4821.

Ogawa:1980:HPR

[Oga80] Tsuyoshi Ogawa. Historical phenomenology — a review essay on “The Essential Tension” — Thomas S. Kuhn, The Essential Tension, Univ. of Chicago Press, 1977, xxiii + 366 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):107–111, 1980. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837249.

Ogawa:1989:GWF

[Oga89] Yutaka Ogawa. Galileo’s work on free fall at Padua: Some remarks on Drake’s interpretation. Historia Scientiarum = International journal of the History of Science Society of Japan, 37(??): 31–49, 1989. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435097;   https://ci.nii.ac.jp/
naid/110009837397.

Ogawa:1993:WOD

[Oga93] Mariko Ogawa. Weismann and the origin of death in evolution theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):223–235, March 1993. CODEN HISCDU. ISSN 0285-4821.

Ogawa:1996:PEP

[Oga96] Tsukane Ogawa. A process of establishment of pre-modern Japanese mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(3):255–262, March 1996. CODEN HISCDU. ISSN 0285-4821.

Ogura:2006:GKE

[Ogu06] Megumi Ogura. The Great Kanto Earthquake: A natural disaster and Japan’s reactions on its way of modernization. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):62–79, July 2006. CODEN HISCDU. ISSN 0285-4821.

Oh:2014:CDM

[Oh14] Chaekun Oh. The cold damage medicine in 17th and 18th century Korea: Comparison through the medical records of China, Japan, and Korea. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(1):1–24, August 2014. CODEN HISCDU. ISSN 0285-4821.

Ohami:1995:TMV

[Oha95] Isao Ohami. On the theory of movement in the Vaisesika system in Ancient India. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1):25–34, June 1995. CODEN HISCDU. ISSN 0285-4821.

Oita:1982:TIT

[Oit82] Akira Oita. Technikgeschichte, Industriearchaologie und technische Müssen: Ein kleiner Überblick unter zeitgeschichtlicher Fragestellung. (German) [History of technology, industrial research and technical means: A brief overview of contemporary history]. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):99–109, 1982. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837290.

Okamato:1993:UOP

[Oka93] Takuji Okamato. Uncertainty of operation: P. W. Bridgman and quantum mechanics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):35–62, July 1993. CODEN HISCDU. ISSN 0285-4821.

Okamoto:1998:BRY

[Oka98] Takuji Okamoto. Book review: Yoshitaka Yamamoto, Katen Rikigaku no Keisei: Newton kara Lagrange e (The Formation of Classical Mechanics: From Newton to Lagrange) (Tokyo: Nihon Hyoron-sha, 1997). ix + 373 pp. Yen6,000. ISBN 4-535-78243-1. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):243–246, March 1998. CODEN HISCDU. ISSN 0285-4821.

Okamoto:2003:UCB

[Oka03] Takuji Okamoto. Uncertainty and controllability: Bridgman, Dingler, and Dewey. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3):233–253, March 2003. CODEN HISCDU. ISSN 0285-4821.

Okamoto:2004:PWB

[Oka04] Takuji Okamoto. Percy Williams Bridgman and the establishment of theoretical physics at Harvard. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14 (1):1–48, July 2004. CODEN HISCDU. ISSN 0285-4821.

Okamoto:2011:FHS

[Oka11] Takuji Okamoto. The first higher school’s instruments for science and engineering education. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20 (3):196–211, March 2011. CODEN HISCDU. ISSN 0285-4821.

Okamoto:2012:BRN

[Oka12] Takuji Okamoto. Book review: Nisio Sigeko, Kagaku Janarizumu no Senkusha, Hyoden, A Critical Biography of Ishiwara Jun: A Pioneer in Science Journalism, Tokyo: Iwanami Shoten, 2011. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(1):40–42, July 2012. CODEN HISCDU. ISSN 0285-4821.

Oki:2013:EMM

[Oki13a] Sayaka Oki. The establishment of ‘mixed mathematics’ and its decline 1600–1800. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(2):82–91, December 2013. CODEN HISCDU. ISSN 0285-4821.

Oki:2013:SIM

[Oki13b] Sayaka Oki. Special issue: Mixed mathematics and its material culture. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(2):77–81, December 2013. CODEN HISCDU. ISSN 0285-4821.

Omata:2004:SHP

[Oma04] Waichiro Omata. A short history of psychiatry in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):200–208, March 2004. CODEN HIS-CDU. ISSN 0285-4821.

Omori:1984:SMS

[Ōmo84] Minoru Ōmori. Some matters in the study of von Siebold from the past to the present and new materials found in relation to Siebold and his works. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):92–110, September 1984. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837322.

Otani:2017:TTD

[Ota17] Takushi Otani. Technology transfer as a dialogical process crossing the Pacific Ocean: Sony’s transistor technology transfer. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(2):112–143, February 2017. CODEN HISCDU. ISSN 0285-4821.

Ozawa:2005:ASA

[Oza05] Takeshi Ozawa. Arnold Sommerfelds Aufenthalt in Japan. (German) [Arnold Sommerfeld’s stay in Japan]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):44–65, July 2005. CODEN HISCDU. ISSN 0285-4821.

Ozawa:2009:IRQ

[Oza09] Takeshi Ozawa. On the initial reception of quantum mechanics in Japan, 1925–1928. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1): 29–42, July 2009. CODEN HISCDU. ISSN 0285-4821.

Ozawa:2010:HRP

[Oza10] Takeshi Ozawa. Hermann Ritter (1827–1874), ein Pionier der westlichen Naturwissenschaften in Japan. (German) [Hermann Ritter (1827–1874), a pioneer of Western science in Japan]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(3):225–234, March 2010. CODEN HIS-CDU. ISSN 0285-4821.

Panza:1991:AFM

[Pan91] Marco Panza. The analytical foundation of mechanics of discrete systems in Lagrange’s Théorie des fonctions analytiques, compared with Lagrange’s earlier treatments of this topic. I. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(2):87–132, December 1991. CODEN HISCDU. ISSN 0285-4821.

Panza:1992:AFM

[Pan92] Marco Panza. The analytical foundation of mechanics of discrete systems in Lagrange’s Théorie des fonctions analytiques, compared with Lagrange’s earlier treatments of this topic. II. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):181–212, March 1992. CODEN HISCDU. ISSN 0285-4821.

Park:2003:CCF

[Par03] Sub Park. Countries to colonies: With the focus on the case of Korea and India. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(3):200–214, March 2003. CODEN HISCDU. ISSN 0285-4821.

Penchevre:2004:LAX

[Pen04] Erwan Penchèvre. L’élimination en algèbre aux XVII^{e} et XVIII^{e} siècles. (French) [Elimination in algebra in the XVII^{th} and XVIII^{th} centuries]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(2):101–117, November 2004. CODEN HISCDU. ISSN 0285-4821.

Patessio:2005:BWD

[PO05] Mara Patessio and Mariko Ogawa. To become a woman doctor in Early Meiji Japan, 1868–1890: Women’s struggles and ambitions. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):159–176, November 2005. CODEN HISCDU. ISSN 0285-4821.

Pyenson:1984:AIJ

[Pye84] Lewis Pyenson. Astronomy and imperialism: J. A. C. Oudemans, the topography of the East Indies, and the rise of the Utrecht Observatory, 1850–1900. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):39–81, March 1984. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837308.

Pyenson:1997:REW

[Pye97] Lewis Pyenson. Review essay: Western Wind: The atomic bomb in American memory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(3): 231–242, March 1997. CODEN HISCDU. ISSN 0285-4821. March.

Pyenson:2000:IIN

[Pye00] Lewis Pyenson. Imperium in Imperio: The natural history of natural knowledge. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):1–15, July 2000. CODEN HISCDU. ISSN 0285-4821.

Pyenson:2005:WHS

[Pye05] Lewis Pyenson. Western historians of science and Oriental science in the Age of Imperialism. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15 (2):97–124, November 2005. CODEN HISCDU. ISSN 0285-4821.

Zhu:2004:RHM

[pZ04] Jian ping Zhu. Research on history of medicine in China in the last five years. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):149–163, March 2004. CODEN HISCDU. ISSN 0285-4821.

Qu:1996:BGM

[Qu96] Anjing Qu. Bian Gang: a mathematician of the 9th Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):17–30, July 1996. CODEN HIS-CDU. ISSN 0285-4821.

Rashed:1994:AYE

[Ras94a] Roshdi Rashed. al-Yāzdī et l’équation ∑_{i=1}^{n} x_{i}^{2} = x^{2}. (French) [al-Yāzdī and the equation ∑_{i=1}^{n} x_{i}^{2} = x^{2}]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):79–101, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Rashed:1994:NVA

[Ras94b] Roshdi Rashed. Notes sur la version arabe des trois premiers livres des Arithmétiques de Diophante, et sur le problème 1.39. (French) [Notes on the Arabic version of the first three books of Arithmetic of Diophantus, and on the problem 1.39]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):39–46, August 1994. CODEN HISCDU. ISSN 0285-4821.

Rashed:1996:ALM

[Ras96] Roshdi Rashed. Archimedean learning in the Middle Ages: the Banū Mūsā. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(1):1–16, July 1996. CODEN HISCDU. ISSN 0285-4821.

Rashed:1997:LSE

[Ras97] Roshdi Rashed. L’histoire des sciences entre Épistémologie et histoire. (French) [The history of sciences between epistemology and history]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):1–10, June 1997. CODEN HISCDU. ISSN 0285-4821.

Rashed:1999:PFD

[Ras99] Roshdi Rashed. Pierre Fermat et les débuts modernes de l’analyse diophantienne. (French) [Pierre Fermat and the modern beginnings of diophantine analysis]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):3– 16, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Rashed:2001:FAG

[Ras01] Roshdi Rashed. Fermat and algebraic geometry. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):1–23, July 2001. CODEN HISCDU. ISSN 0285-4821.

Rie:2012:LKO

[Rie12] Hogetsu Rie. From the laboratory to the kitchen: The oral history of nutrition and its application to the history of science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(2):125–141, December 2012. CODEN HISCDU. ISSN 0285-4821.

Roberts:2007:ODB

[RS07] Gerrylynn K. Roberts and Anna E. Simmons. The overseas dimensions of British chemical societies, c.1890–c.1950. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(3):224–243, March 2007. CODEN HIS-CDU. ISSN 0285-4821.

Rao:2003:GRH

[RV03] K. Srinivasa Rao and G. Vanden Berghe. Gauss, Ramanujan and hypergeometric series revisited. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13 (2):123–133, November 2003. CODEN HISCDU. ISSN 0285-4821.

Saito:1985:BIE

[Sai85] Ken Saito. Book II of Euclid’s Elements in the light of the theory of conic sections. Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):31–60, March 1985. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837327.

Saito:1986:CRE

[Sai86] Ken Saito. Compounded ratio in Euclid and Apollonius. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):25–59, September 1986. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435072;   https://ci.nii.ac.jp/naid/110009837353.

Saito:1993:DRB

[Sai93] Ken Saito. Duplicate ratio in Book VI of Euclid’s Elements. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(2):115–135, December 1993. CODEN HISCDU. ISSN 0285-4821.

Saitoh:1996:SHB

[Sai96] Hikaru Saitoh. Studies on the history of biology in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(2):109–122, December 1996. CODEN HISCDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Saito:2013:IGR

[Sai13] Hirofumi Saito. The Institute of Genetics from 1939 to 1940: Reconsidering Lysenko’s intervention in Soviet genetics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(3):227–236, March 2013. CODEN HIS-CDU. ISSN 0285-4821.

Saito:2016:EMN

[Sai16] Hirofumi Saito. The early movement for normalizing biology education in Russia after Lysenko’s downfall in 1965. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):3–14, March 2016. CODEN HISCDU. ISSN 0285-4821.

Sakanoue:1993:YNH

[Sak93] Masanobu Sakanoue. Yamaoka Nozomu, a ‘humanistic’ historian of chemistry. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):83–94, July 1993. CODEN HISCDU. ISSN 0285-4821.

Sakano:1998:PSL

[Sak98] Tōru Sakano. Population, sex and labor: Medicine and public hygiene in Micronesia under Japanese rule, 1914–1945. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):173–186, December 1998. CODEN HISCDU. ISSN 0285-4821.

Sasaki:1980:BRS

[Sas80] Chikara Sasaki. Book review: Shokichi Iyanaga, Shuntaro Ito and Tohru Sato, Girisha no Sugaku (Greek Mathematics), Kyoritsushuppan, Tokyo, 1979, 280 pp., Yen 3,800. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):112–113, 1980. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837250.

Sasaki:1981:BRK

[Sas81a] Chikara Sasaki. Book review: Koshiro Nakamura, Kinsei Sugaku no Rekishi-Bisekibungaku no Keisei o megutte (The History of Mathematics in the Early Modern Period — Around the Formation of the Differential and Integral Calculus), Nippon Hyoron-sha, Tokyo, 1980, Yen 3,700. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):128–129, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837260.

Sasaki:1981:BRT

[Sas81b] Chikara Sasaki. Book review: Toshisada Endo, Zoshu Nippon Suguakushi (A History of Japanese Mathematics, Enlarged and Revised), Second Definitive Edition, edited by Yoshio Mikami and revised by Akira Hirayama, Koseisha-Koseikaku, Tokyo, 1981, 900 pp., Yen 15,000. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):123–124, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837272.

Sasaki:1981:SSO

[Sas81c] Chikara Sasaki. Scientific studies at Oxford and Cambridge in the Seventeenth Century: A research review. Historia Scientiarum = International journal of the History of Science Society of Japan, 20 (??):57–75, 1981. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837255.

Sasaki:1985:ATP

[Sas85] Chikara Sasaki. The acceptance of the theory of proportion in the Sixteenth and Seventeenth Centuries: Barrow’s reaction to the analytic mathematics. Historia Scientiarum = International journal of the History of Science Society of Japan, 29(??):83–116, September 1985. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii. ac.jp/naid/110009837342. Special Issue, “The Historiography of the History of Mathematics”.

Sasaki:1992:ACE

[Sas92] Chikara Sasaki. Announcement of a change of Editor. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):83, December 1992. CODEN HIS-CDU. ISSN 0285-4821.

Sasaki:1996:HMM

[Sas96] Chikara Sasaki. Historians of mathematics in modern Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(2):67–78, December 1996. CODEN HISCDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Sasaki:1999:FJS

[Sas99a] Chikara Sasaki. The French and Japanese schools of algebra in the seventeenth century: a comparative study. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):17–26, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Pierre de Fermat mathématicien français (Tokyo, 1998).

Sasaki:1999:PFM

[Sas99b] Chikara Sasaki. Pierre de Fermat mathématicien français. (French) [Pierre de Fermat French mathematician]. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):i–ii, 1–71, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français. Invited lectures from the colloquium held in Tokyo, April 27–28, 1998, Historia Sci. (2) 9 (1999), no. 1.

Sasaki:2001:SCP

[Sas01] Chikara Sasaki. Science and chrysanthemum: The paradox of enlightenment in Imperial Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11 (1):24–47, July 2001. CODEN HISCDU. ISSN 0285-4821.

Sasaki:2003:BRR

[Sas03] Chikara Sasaki. Book review: Roshdi Rashed, Les Mathématiques Infinitésimales, Vols. 1–4. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13 (1):59–63, July 2003. CODEN HISCDU. ISSN 0285-4821.

Sasaki:2007:BRR

[Sas07] Chikara Sasaki. Book review: Roshdi Rashed, Geometry and Dioptrics in Classical Islam, 2005, 1176 + xx pp. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(1):56–59, July 2007. CODEN HISCDU. ISSN 0285-4821.

Sasamoto:2009:KVA

[Sas09] Ikuo Sasamoto. Korean victims of the atomic bomb. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):160–169, December 2009. CODEN HIS-CDU. ISSN 0285-4821.

Sato:1981:ALW

[Sat81] Tohru Sato. Archimedes’ lost works on the centers of gravity of solids, plane figures and magnitudes. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??): 1–41, 1981. CODEN HISCDU. ISSN 0285-4821. URL https://ci.

nii.ac.jp/naid/110009837253.

Sato:1985:QSA

[Sat85] Tohru Sato. Quadrature of the surface area of a sphere in the Early Middle Ages — Johannes de Tinemue and Banū Mūsā. Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):61–90, March 1985. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837328.

Sato:1986:RMP

[Sat86] Tohru Sato. A reconstruction of The Method Proposition 17, and the development of Archimedes’ thought on quadrature: Why did Archimedes not notice the internal connection in the problems dealt with in many of his works? Part One. Historia Scientiarum = International journal of the History of Science Society of Japan, 31(??):61–86, September 1986. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435073;
https://ci.nii.ac.jp/naid/110009837354.

Sato:1987:RMP

[Sat87] Tohru Sato. A reconstruction of The Method Proposition 17, and the development of Archimedes’ thought on quadrature: Why did Archimedes not notice the internal connection in the problems dealt with in many of his works? Part Two. Historia Scientiarum = International journal of the History of Science Society of Japan, 32 (??):75–142, 1987. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837361.

Satofuka:1988:BRE

[Sat88] Fumihiko Satofuka. Book review: Erik Baark and Andrew Jamison, eds., Technological Development in China, India and Japan: Crosscultural Perspective, The Macmillan Press Ltd., London, 1986, 255 pp., 10 indices. Historia Scientiarum = International journal of the History of Science Society of Japan, 34(??):81–84, 1988. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837377.

Satofuka:1989:SAS

[Sat89] Fumihiko Satofuka. Some aspects to science studies in Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):65–72, 1989. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837408.

Satofuka:1990:SAD

[Sat90] Fumihiko Satofuka. Some aspects to the debate on scientific tradition in Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):45–52, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837422.

Satofuka:1991:BRT

[Sat91] Fumihiko Satofuka. Book review: Takeshi Hayashi, The Japanese Experience in Technology: From Transfer to Self-Reliance (Tokyo: United Nations University Press, 1990). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):75–76, June 1991. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1992:FSAa

[Sat92a] Fumihiko Satofuka. Forum: Some aspects to the debate on scientific tradition in Japan (II). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1 (3):213–220, March 1992. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1992:FSAb

[Sat92b] Fumihiko Satofuka. Forum: Some aspects to the debate on scientific tradition in Japan (III). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2 (1):61–69, June 1992. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1993:BRS

[Sat93] Fumihiko Satofuka. Book review: Shigeru Nakayama, Science, Technology and Society in Postwar Japan (London: Kegan Paul International, 1991). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):237–238, March 1993. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1994:BRY

[Sat94] Fumihiko Satofuka. Book review: Yukiko Fukasaku, Technology and Industrial Development in Pre-war Japan: Mitsubishi Nagasaki Shipyard 1884–1934 (London: Routledge, 1992). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):57–59, August 1994. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1995:BRS

[Sat95] Fumihiko Satofuka. Book review: Saneh Chamarik and Susantha Goonatilake, eds., Technological Independence: The Asian Experience (Tokyo: United Nations University, 1994). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):251–254, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Satofuka:1996:FSA

[Sat96] Fumihiko Satofuka. Forum: Some aspects to the debate on scientific tradition in Japan (IV). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5 (3):263–268, March 1996. CODEN HISCDU. ISSN 0285-4821.

Sathyamurthy:1998:RES

[Sat98a] T. V. Sathyamurthy. Review essay: SVAYAMBHU: the life and work of Srinivasa Ramanujan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7 (3):235–241, March 1998. CODEN HISCDU. ISSN 0285-4821.

Sato:1998:TRP

[Sat98b] Ken’ichi Sato. On the theory of regular polygons in traditional Japanese mathematics: reconstruction of the process for the calculation of the degree of Kaihoshiki appearing in the Taisei Sankei by Seki and Takebe brothers. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8 (1):71–85, August 1998. CODEN HISCDU. ISSN 0285-4821.

Sato:2001:CGP

[Sat01] Yasushi Sato. A contested gift of power: American assistance to Japan’s space launch vehicle technology, 1965–1975. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(2):176–204, November 2001. CODEN HIS-CDU. ISSN 0285-4821.

Satofuka:2005:BRM

[Sat05] Fumihiko Satofuka. Book review: Maurice Wilkins, The Third Man of the Double Helix: The Autobiography of Maurice Wilkins. Oxford: Oxford University Press, 2003. 352 pp. ISBN 0-19-860665-6. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(1):66–68, July 2005. CODEN HISCDU. ISSN 0285-4821.

Satofuka:2006:BRW

[Sat06a] Fumihiko Satofuka. Book review: Wataru Iijima, Mararia to teikoku: shokuminnchiigaku to Higashi azia no kouiki chitujyo (Hidden History of Malaria: Colonial/Imperial Medicine and an Integrated Regional Order in 20th Century East Asia). 2005. 387 pp., ISBN 4-13-026210-6. 6,800 yen. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2): 210–211, November 2006. CODEN HISCDU. ISSN 0285-4821.

Satofuka:2006:SWC

[Sat06b] Fumihiko Satofuka. Science, war, and colonization in East Asia: An introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):209–212, March 2006. CODEN HISCDU. ISSN 0285-4821.

Sato:2007:SAJ

[Sat07] Yasushi Sato. Systems approach in the Japanese National Railways: How was the systems integration of Shinkansen done? Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(2):65–88, December 2007. CODEN HISCDU. ISSN 0285-4821.

Sato:2009:BRT

[Sat09] Yasushi Sato. Book review: Takehiko Hashimoto, Historical Essays on Japanese Technology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19 (1):75–77, July 2009. CODEN HISCDU. ISSN 0285-4821.

Sato:2012:PPK

[Sat12a] Masahiro Sato. Popularization of psychiatric knowledge in modern Japan at the turn of the twentieth century: Focusing on the newspaper coverage of mental disorders. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22(2):110–124, December 2012. CODEN HISCDU. ISSN 0285-4821.

Sato:2012:MIB

[Sat12b] Yasushi Sato. Managing the interface between politics and technology: Itokawa Hideo, Shima Hideo, and the early Japanese space programs. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 21(3):193–210, March 2012. CODEN HISCDU. ISSN 0285-4821.

Sato:2013:SSC

[Sat13] Kenichi Sato. Surveying in seventeenth-century Japan: Technology transfer from The Netherlands to Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(2):92–112, December 2013. CODEN HISCDU. ISSN 0285-4821.

Sato:2017:PRA

[Sat17] Yasushi Sato. Probabilistic Risk Assessment (PRA) in human spaceflight: Technical, cultural, political, and economic factors behind NASA’s stance to risk. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(1): 63–81, August 2017. CODEN HISCDU. ISSN 0285-4821.

Sesiano:1991:ATC

[Ses91] Jacques Sesiano. An Arabic treatise on the construction of bordered magic squares. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):13–32, ???? 1991. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435113;   https://ci.nii.ac.jp/
naid/110009837430.

Setoguchi:2009:BRP

[Set09] Akihisa Setoguchi. Book review: Paul Callomon and Akira Tada, Yoichiro Hirase and His Role in Japanese Malacology. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(1):72–74, July 2009. CODEN HISCDU. ISSN 0285-4821.

Sheynin:1999:DPL

[She99] Oscar Sheynin. The discovery of the principle of least squares. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(3):249–264, March 1999. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2000:BSR

[She00] Oscar Sheynin. Bessel: some remarks on his work. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):77–83, July 2000. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2001:GBA

[She01a] Oscar Sheynin. Gauss, Bessel and the adjustment of triangulation. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(2):168–175, November 2001. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2001:PS

[She01b] Oscar Sheynin. Pirogov as a statistician. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):213–225, March 2001. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2001:SSP

[She01c] Oscar Sheynin. Social statistics and probability theory in the 19th Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 11(1):86–111, July 2001. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2002:SNS

[She02] Oscar Sheynin. Simon Newcomb as a statistician. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(2):142–167, November 2002. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2003:GPB

[She03a] Oscar Sheynin. Geometric probability and the Bertrand paradox. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):42–53, July 2003. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2003:MMN

[She03b] Oscar Sheynin. Mises on mathematics in Nazi Germany. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(2):134–146, November 2003. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2006:BRJ

[She06a] Oscar Sheynin. Book review: Jacob Bernoulli, The Art of Conjecturing together with Letter to a Friend on Sets in Court Tennis. Translated with an introduction and notes by Edith Dudley Sylla. 2006. xx + 430 pp. ISBN 0-8018-8235-4. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):212–214, November 2006. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2006:BRT

[She06b] Oscar Sheynin. Book review: Theodore M. Porter, Karl Pearson: The Scientific Life in a Statistical Age, 2004. viii + 342 pp. ISBN 0-691-11445-5. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):206–209, November 2006. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2006:MWT

[She06c] Oscar Sheynin. Markov’s work on the treatment of observations. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):80–95, July 2006. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2007:TVM

[She07] Oscar Sheynin. The true value of a measured constant and the theory of errors. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(1):38–48, July 2007. CODEN HISCDU. ISSN 0285-4821.

Sheynin:2008:BAD

[She08] Oscar Sheynin. Bortkiewicz’ alleged discovery: the Law of Small Numbers. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(1):36–48, July 2008. CODEN HISCDU. ISSN 0285-4821. URL http://www.

sheynin.de/download/law_sm_no_Hist_scient.pdf.

Shimodaira:1981:IJ

[Shi81] Kazuo Shimodaira. On Idai of Jingoki. Historia Scientiarum = International journal of the History of Science Society of Japan, 21 (??):87–101, 1981. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837268.

Shimodaira:1992:ODO

[Shi92] Kazuo Shimodaira. Obituary: Dr. Ohya Shinichi’s works. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):73–75, June 1992. CODEN HISCDU. ISSN 0285-4821.

Shinoda:1998:SPN

[Shi98] Mariko Shinoda. Scientists as preservationists: Natural monuments in Japan, 1906–1931. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2): 141–156, December 1998. CODEN HISCDU. ISSN 0285-4821.

Shiba:2000:DCG

[Shi00] Tetsuo Shiba. Dutch chemist Gratama and chemistry in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(3):181–190, March 2000. CO-DEN HISCDU. ISSN 0285-4821.

Shi:2003:NII

[Shi03] Yunli Shi. A note on the Islamic influence on the astronomical instrumentation of the Chosôn dynasty. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(1):33–41, July 2003. CODEN HISCDU. ISSN 0285-4821.

Shin:2004:WMK

[Shi04] Dong-Won Shin. Western medicine, Korean government, and imperialism in late nineteenth-century Korea: The cases of the Choson government hospital and smallpox vaccination. Historia Scientiarum.

Second Series. International Journal of the History of Science Society of Japan, 13(3):164–175, March 2004. CODEN HISCDU. ISSN 0285-4821.

Shin:2008:HFD

[Shi08] Dongwon Shin. How four different political systems have shaped the modernization of traditional Korean medicine between 1900 and 1960. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):225–241, March 2008. CODEN HISCDU. ISSN 0285-4821.

Shortland:1990:CPS

[Sho90] Michael Shortland. ‘On the connexion of the physical sciences’ classification and organization in Early Nineteenth-Century science. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):17–36, 1990. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435111;
https://ci.nii.ac.jp/naid/110009837420.

Sigeko:2000:IJQ

[Sig00] Nisio Sigeko. Ishiwara Jun’s quantum theory, 1911–1915. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):120–129, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Siu:1995:MEA

[Siu95] Man-Keung Siu. Mathematics education in ancient China: what lesson do we learn from it? Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(3):223–232, March 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 2.

Slavutskii:1995:SAP

[Sla95] I. Sh. Slavut·skiĭ. Staudt and arithmetical properties of Bernoulli numbers. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(1):69–74, June 1995. CODEN HISCDU. ISSN 0285-4821.

Smith:2002:KC

[Smi02] Thomas C. Smith. The Kyoto cyclotron. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):74–82, July 2002. CODEN HISCDU. ISSN 0285-4821.

Smith:2018:BID

[Smi18] H. A. Smith. Beyond indulgence: Diet-induced illnesses in Chinese medicine. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):233–253, 2018. CODEN HISCDU. ISSN 0285-4821.

Satofuka:1987:PWS

[SR87] Fumihiko Satofuka and Abdur Rahman. Post war science and technology policy of Japan: A view from a developing country. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):61–76, 1987. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837368.

Steen:1996:TMK

[Ste96] Tomoko Y. Steen. A tribute to Motoo Kimura. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(2):159–169, December 1996. CODEN HIS-CDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Steen:2008:COT

[Ste08] Tomoko Y. Steen. The case of Ohta Tomoko: A woman geneticist in the neutralist–selectionist evolution controversy. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(2):172–184, December 2008. CODEN HIS-CDU. ISSN 0285-4821.

Sugimoto:1986:BZG

[Sug86] Toshio Sugimoto. Berichtigungen der Zahlenrechnungen, die in den Gauss’schen Werken enthalten sind. (German) [Corrections of numerical calculations included in Gauss’s works]. Historia Scientiarum = International journal of the History of Science Society of Japan, 30(??):93–98, March 1986. CODEN HISCDU. ISSN 0285-

4821. URL https://ci.nii.ac.jp/naid/110009837350.

Sugiyama:1992:SPC

[Sug92] Shigeo Sugiyama. The significance of the particulate conception of matter in John Tyndall’s physical researches. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):119–138, December 1992. CODEN HISCDU. ISSN 0285-4821.

Sugiyama:2004:TKM

[Sug04] Shigeo Sugiyama. Traditional Kampo medicine: Unauthenticated in the Meiji era. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(3):209–223, March 2004. CODEN HISCDU. ISSN 0285-4821.

Sugimoto:2013:ECB

[Sug13] Mai Sugimoto. Edmund C. Berkeley as a popularizer and an educator of computers and symbolic logic. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(1):1–23, July 2013. CODEN HISCDU. ISSN 0285-4821.

Suzuki:1981:ESR

[Suz81] Shigeaki Suzuki. An experimental study for the reproduction of ‘so’ (a dairy product) prescribed in the En-gi Shiki in the 9th and 10th Centuries of Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):115–118, 1981. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837270.

Suzuki:1982:BRI

[Suz82] Takanori Suzuki. Book review: Ibn Sina, al-Qanun fi ’l-tibb, Japanese Translation by Hitoshi Igarashi, Tatsuo Sato & Shuntaro Ito, Asahi-shuppan, Tokyo, 1981, 374 pp., Yen 4800. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):131–132, 1982. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837281.

Suzuki:1999:ODH

[Suz99] Takeo Suzuki. Obituary: Dr. Hirayama Akira (1904–1998). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(3):265–272, March 1999. CODEN HIS-CDU. ISSN 0285-4821.

Suzuki:2012:ISI

[Suz12] Akihito Suzuki. Introduction: Special issue: Social history of medicine in modern Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 22 (2):63–67, December 2012. CODEN HISCDU. ISSN 0285-4821.

Siu:1999:OCT

[SV99] Man-Keung Siu and Alexeï Volkov. Official curriculum in traditional Chinese mathematics: how did candidates pass the examinations?

Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):85–99, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français.

Sarina:2015:CSE

[SW15] Sarina and Ying Wang. A comparison of the spread of the English translation of Euclidean geometry in 19th century China and Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(2):88–98, February 2015. CODEN HISCDU. ISSN 0285-4821.

Szabo:1981:AMG

[Sza81] Arpad Szabo. Astronomische Messungen bei den Griechen im 5. Jahrhundert v. Chr. und ihr Instrument. (German) [Astronomical measurements of the Greeks in the 5th Century BCE and its instrument]. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):1–26, 1981. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837263.

Szabo:1985:SPG

[Sza85] Árpád Szabó. Strabon und Pytheas — die geographische Breite von Marseille — Zur Frühgeschichte der mathematischen Geographie. (German) [Straban and Pytheas — the latitude of Marseille — the early history of mathematical geography]. Historia Scientiarum = International journal of the History of Science Society of Japan, 29 (??):3–15, September 1985. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837338. Special Issue,

“The Historiography of the History of Mathematics”.

Szabo:1988:GKM

[Sza88] Arpad Szabo. Der Gnomon als Kalender: In Memoriam Prof. Ryoji Shizuma. (German) [The gnomon as a calendar: In memoriam: Prof. Ryoji Shizuma]. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):31–38, 1988. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435086;   https://ci.nii.ac.jp/naid/110009837380.

Takabayasi:1984:BRC

[Tak84a] Takehiko Takabayasi. Book review: The Collected Works of Sin-itiro Tomonaga, Vol. 1–12, Misuzu Shobo, Tokyo, 1981–83. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):111–116, September 1984. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837323.

Takahashi:1984:MFB

[Tak84b] Ken’ichi Takahashi. The mathematical foundations of Bradwardine’s Rule. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):19–38, March 1984. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837307.

Takase:1985:ONP

[Tak85] Masahito Takase. Origine de la notion de pseudoconvexité dans la théorie des fonctions analytiques de plusieurs variables complexes: Signification du problème inverse de Hartogs. (French) [Origin of the notion of pseudoconvexity in the theory of analytic functions of several complex variables: Meaning of the inverse problem of Hartogs]. Historia Scientiarum = International journal of the History of Science Society of Japan, 28(??):139–151, March 1985. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837330.

Takata:1987:GTP

[Tak87a] Seiji Takata. Ganot’s textbooks of physics introduced into Japan. Historia Scientiarum = International journal of the History of Science Society of Japan, 33(??):25–41, 1987. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435078;
https://ci.nii.ac.jp/naid/110009837366.

Takata:1987:TCA

[Tak87b] Seiji Takata. A textual comparison of all the 4 editions of Mach’s Warmelehre. Historia Scientiarum = International journal of the History of Science Society of Japan, 32(??):63–73, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837360.

Takata:1992:AKS

[Tak92a] Seiji Takata. Amano Kiyoshi the “scientific” historian of science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):1–12, June 1992. CODEN HISCDU. ISSN 0285-4821.

Takata:1992:BRC

[Tak92b] Seiji Takata. Book review: Claudio Egidi, ed., Giovanni Giorgi and his Contribution to Electrical Metrology (1990). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(3):223–224, March 1992. CODEN HISCDU. ISSN 0285-4821.

Takahashi:1993:GLHb

[Tak93a] Ken’ichi Takahashi. Galileo’s labyrinth: his struggle for finding a way out of his erroneous law of natural fall. I. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):169–202, March 1993. CODEN HISCDU. ISSN 0285-4821.

Takahashi:1993:GLHa

[Tak93b] Ken’ichi Takahashi. Galileo’s labyrinth: his struggle for finding a way out of his erroneous law of natural fall. II. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):1–34, July 1993. CODEN HISCDU. ISSN 0285-4821.

Takata:1993:JBF

[Tak93c] Seiji Takata. J. B. Fourier in the history of thermal radiation research. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(3):203–221, March 1993. CODEN HISCDU. ISSN 0285-4821.

Takata:1997:AJP

[Tak97] Seiji Takata. Activity of Japanese physicists in the learned societies from 1877 to 1926. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):81–92, December 1997. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Takuji:1998:DDA

[Tak98] Okamoto Takuji. Demystifying dimensional analysis: the beginning of Bridgman’s scrutiny of theoretical physics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):21–52, August 1998. CODEN HISCDU. ISSN 0285-4821.

Takata:2000:KAP

[Tak00] Seiji Takata. Kiyoshi Amano’s pioneering studies on the history of quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):112–119, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Takeda:2008:FSS

[Tak08] Tokimasa Takeda. The formation of the study of Shushu and its development in the Middle Ages: a tentative study of a field of scientific study peculiar to East Asia. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(3):161–174, March 2008. CODEN HISCDU. ISSN 0285-4821.

Takahashi:2009:OMA

[Tak09] Hiroko Takahashi. One minute after the detonation of the atomic bomb: the erased effects of residual radiation. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):146–159, December 2009. CODEN HISCDU. ISSN 0285-4821.

Takuwa:2013:HTN

[Tak13] Yoshimi Takuwa. The historical transformation of Newton’s experimentum cruris: Pursuit of the demonstration of color immutability. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(2):113–140, December 2013. CODEN HISCDU. ISSN 0285-4821.

Takeda:2015:I

[Tak15] Tokimasa Takeda. Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 24(2):47–49, February 2015. CODEN HISCDU. ISSN 0285-4821.

Tanaka:1983:AOM

[Tan83] Yutaka Tanaka. Aristotelian ontology and modal syllogistic reconstructed. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):87–109, March 1983. CO-DEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837300.

Tanaka:1987:EWP

[Tan87] Yutaka Tanaka. Einstein and Whitehead: The principle of relativity reconsidered. Historia Scientiarum = International journal of the History of Science Society of Japan, 32(??):43–61, 1987. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837359.

Tanimoto:1990:JRA

[Tan90] Tsutomu Tanimoto. The Japanese reception of Alfred Wegener’s theory of the continental drift. Historia Scientiarum = International journal of the History of Science Society of Japan, 39(??): 15–27, 1990. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435104;   https://ci.nii.ac.jp/
naid/110009837412.

Tanaka:1992:KTH

[Tan92] Setsuko Tanaka. Kuwaki turns to history of science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(2):85–108, December 1992. CODEN HIS-CDU. ISSN 0285-4821.

Tanaka:2006:STM

[Tan06] Hiroaki Tanaka. The science and technology mobilization system in Imperial Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):213–221, March 2006. CODEN HISCDU. ISSN 0285-4821.

Editor:1985:PSN

[The85] The Editor. Preface for the special number. Historia Scientiarum = International journal of the History of Science Society of Japan, 29 (??):1, September 1985. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837337. Special Issue, “The

Historiography of the History of Mathematics”.

Tamazawa:2017:AIN

[THI17] Harufumi Tamazawa, Hisashi Hayakawa, and Kiyomi Iwahashi. Astronomy and intellectual networks in the late 18th Century in Japan: A case study of Fushimi in Yamashiro. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(3):172–191, March 2017. CODEN HISCDU. ISSN 0285-4821.

Tian:1999:JTD

[Tia99] Miao Tian. Jiegenfang, tianyuan, and daishu: algebra in Qing China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(1):101–119, July 1999. CODEN HISCDU. ISSN 0285-4821. Colloque: Pierre de Fermat mathematicien français.

Tochinai:2018:JGS

[Toc18] F. Tochinai. Japanese geological scientists and their activities with respect to science communication: With special reference to Professor Seitaro Tsuboi and Chidanken. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(3):319–333, 2018. CODEN HISCDU. ISSN 0285-4821.

Toda:1988:BRH

[Tod88a] Kiyoshi Toda. Book review: Hajime Kashiwagi and Mie Kashiwagi, eds., David M. Knight (original author), Kagakushi Nyumon, Shiryo e no Approach (A Guide to the History of Science: Approach to Sources), 1984, 527, 190 pp., Yen 12,000, Uchida Rokakuho, Tokyo. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):92–93, 1988. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837384.

Toda:1988:BRS

[Tod88b] Kiyoshi Toda. Book review: Shigeru Nakayama and Hiroshi Ishiyama, Kagakushi Kenkyu Nyumon (A Guide to Research in the History of Science), 1987, 352 pp., Yen 2,400, Tokyo Daigaku Shuppan Kai, Tokyo. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):91–92, 1988. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837383.

Tomita:1981:JTJ

[Tom81] Tetsuo Tomita. The Japanese translation of Johann Beckmann’s Beyträge zur Geschichte der Erfindungen, 2 vols., Daiyamondo-sha, Vol. I, 1980, 429 pp., Yen 3,800 / the Japanese translation of Johann Beckmann’s Beyträge zur Geschichte der Erfindungen, 2 vols., Daiyamondo-sha, Vol. II, 1981, 483 pp., Yen 4,200. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):125–126, 1981. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837273.

Toyoda:1982:EDS

[Toy82] Kazuji Toyoda. Early development of stone-working in Greece and its contribution to Achaemenid Persia. Historia Scientiarum = International journal of the History of Science Society of Japan, 23 (??):111–127, 1982. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837291.

Tsuzuki:1981:ASD

[Tsu81] Masanobu Tsuzuki. Art and science at the dawn of modern science. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):43–55, 1981. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837254.

Tsukahara:1994:EQC

[Tsu94] Togo Tsukahara. Elimination of qi by chemical specification: Shift of understanding of Western theory of matter in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):1–24, August 1994. CODEN HIS-CDU. ISSN 0285-4821.

Tsukahara:2000:WCD

[Tsu00] Togo Tsukahara. The Westernization of chemistry from different angles: An examination of three manuscripts by contemporaries of Yoan Udagawa and his Seimi Kaiso. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(3):191–214, March 2000. CODEN HISCDU. ISSN 0285-4821.

Tucker:2014:SIM

[Tuc14] Jennifer Tucker. Science institutions in modern British visual culture: The British Association for the Advancement of Science, 1831– 1931. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(3):191–213, March 2014. CODEN HISCDU. ISSN 0285-4821.

Tucker:2016:OMG

[Tuc16] Jennifer Tucker. “To obtain more general attention for the objects of science”: The depiction of popular science in Victorian illustrated news. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 25(3):190–215, March 2016. CODEN HISCDU. ISSN 0285-4821.

Uchida:1990:RCC

[Uch90] Masao Uchida. The recognition of chemical composition in the Eighteenth Century. Historia Scientiarum = International journal of the History of Science Society of Japan, 40(??):41–57, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435109;   https://ci.nii.ac.jp/naid/110009837417.

Uematsu:1996:HPJ

[Uem96] Eisui Uematsu. Historians of physics in Japan: From the methodological perspective of the history of science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 6(2):79–86, December 1996. CODEN HISCDU. ISSN 0285-4821. Special Issue: Historical Studies of Sciences in Modern Japan.

Unschuld:1984:FCM

[Uns84] Paul U. Unschuld. Frühe christliche Missionshospitäler in China im 19. Jahrhundert. (German) [Early Christian mission hospitals in China in the 19th Century]. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):43–50, September 1984. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837319.

Uematsu:2000:HNF

[UTN00] Eisui Uematsu, Tatsuoki Takeda, and Sigeko Nisio. History of nuclear fusion research in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10 (1):16–45, July 2000. CODEN HISCDU. ISSN 0285-4821.

Vizgin:2005:HSA

[Viz05] Vladimir Pavlovich Vizgin. The history of the Soviet Atomic Project. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(3):182–200, March 2005. CODEN HISCDU. ISSN 0285-4821.

Volkov:1994:CAC

[Vol94] Alexeï Volkov. Calculation of π in ancient China: from Liu Hui to Zu Chongzhi. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(2):139–157, December 1994. CODEN HISCDU. ISSN 0285-4821. Special Issue Mathematics in Asia 1.

Wachtler:1982:FTM

[Wac82] Eberhard Wachtler. Freiberg — ein traditionsreiches montanwissenschaftliches Zentrum. (German) [Freiberg — a traditional mining science center]. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):15–19, 1982. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837285.

Wachtler:1992:BRT

[Wäc92] Eberhard Wächtler. Book review: Tadaaki Kimoto, Die Entwicklung der Technologie und ihre Stellung in der Wissenschaft während der gesellschaftlichen Umwälzung an der Wende vom 18. zum 19. Jahrhundert (1991). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 2(1):70–72, June 1992. CODEN HISCDU. ISSN 0285-4821.

Walker:2005:GWN

[Wal05] Mark Walker. German work on nuclear weapons. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(3):164–181, March 2005. CODEN HIS-CDU. ISSN 0285-4821.

Wang:2006:ASC

[Wan06] Jessica Wang. American scientists and Cold War politics: Current and future research directions. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16 (1):29–34, July 2006. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110006448380.

Watanabe:1990:MOA

[Wat90a] Masao Watanabe. Makyo: An object of art and science. Historia Scientiarum = International journal of the History of Science Society of Japan, 40(??):1–7, 1990. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435106;
https://ci.nii.ac.jp/naid/110009837414.

Watanabe:1990:THS

[Wat90b] Masao Watanabe. Teaching history of science in Japanese universities. Historia Scientiarum = International journal of the History of Science Society of Japan, 40(??):59–69, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837418.

Watanabe:1993:OMY

[Wat93] Masao Watanabe. Obituary: Masakazu Yoshinaka (1944–1992). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 3(1):99–101, July 1993. CODEN HISCDU. ISSN 0285-4821.

Wen:2016:AAB

[Wen16] Heng Wen. “Autonomy”, “alienation”, and beyond: The relationship between Chinese physicists and the Kuomintang government, 1928–1945. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 26(1):42–64, March 2016. CODEN HISCDU. ISSN 0285-4821.

Wittner:2008:PMP

[Wit08] David G. Wittner. Practice makes perfect: Foreign knowledge and the Kamaishi experiment. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18 (2):88–99, December 2008. CODEN HISCDU. ISSN 0285-4821.

Xiaoli:1999:WSA

[Xia99] Sun Xiaoli. A wrong statement about Leibniz and his interpretation of Chinese I Ching figure. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(3): 239–247, March 1999. CODEN HISCDU. ISSN 0285-4821.

Xu:1999:TKR

[Xu99] Zelin Xu. Takebe Katahiro and Romberg algorithm. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(2):155–164, November 1999. CODEN HISCDU. ISSN 0285-4821.

Yabuuti:1980:AEW

[Yab80] Kiyosi Yabuuti. Astronomy in the East and the West: The development and characteristics. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??):1–9, 1980. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837243.

Yabuuti:1997:IAC

[Yab97] Kiyosi Yabuuti. Islamic astronomy in China during the Yuan and Ming dynasties. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):11–43, June 1997. CODEN HISCDU. ISSN 0285-4821. Translated, revised and with an appendix by Benno van Dalen, with an epilogue by Michio Yano and van Dalen.

Yadav:2000:AWI

[Yad00] B. S. Yadav. André Weil’s India in the early thirties. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(1):84–91, July 2000. CODEN HISCDU. ISSN 0285-4821.

Yagi:1981:AAC

[Yag81] Eri Yagi. Analytical approach to Clausius’s first memoir on mechanical theory of heat (1850). Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??): 77–94, 1981. CODEN HISCDU. ISSN 0285-4821. URL https://

ci.nii.ac.jp/naid/110009837256.

Yagi:1984:DJS

[Yag84] Eri Yagi. Derek J. De Solla Price (1922–1983), personal reminiscences. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):115–117, March 1984. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837311.

Yagi:2010:MLB

[Yag10] Eri Yagi. To the memory of Lawrence Badash (1934–2010). Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(2):147–149, December 2010. CODEN HISCDU. ISSN 0285-4821.

Yajima:1998:HPM

[Yaj98] Michiko Yajima. Hilgendorf predated Morse in bringing Charles Darwin’s theory of evolution to Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(2):133–140, December 1998. CODEN HISCDU. ISSN 0285-4821.

Yajima:2006:JWG

[Yaj06] Michiko Yajima. Japanese wartime geology: A case study in Northeast China. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(3):222–232, March 2006. CODEN HISCDU. ISSN 0285-4821.

Yajima:2007:BRG

[Yaj07] Michiko Yajima. Book review: Gregory Clancey, Earthquake Nation: The Cultural Politics of Japanese Seismicity, 1868–1930. Berkeley: University of California Press, 2006. 331 pp. ISBN: 0-520-24607-1. $49.95. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17(1):54–55, July 2007. CODEN HISCDU. ISSN 0285-4821.

Yamazaki:1982:EDN

[Yam82] Masakatsu Yamazaki. An early development of nuclear physics in Japan: In the case at Osaka Imperial University. Historia Scientiarum = International journal of the History of Science Society of Japan, 22(??):69–80, 1982. CODEN HISCDU. ISSN 0285-4821.

URL https://ci.nii.ac.jp/naid/110009837276.

Yamaguchi:1983:FHV

[Yam83a] Chuhei Yamaguchi. On the formation of Helmholtz’ view of life processes in his studies of fermentation and muscle action: In relation to his discovery of the Law of Conservation of Energy. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):29–37, September 1983. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110009837440.

Yamazaki:1983:BRJ

[Yam83b] Masakatsu Yamazaki. Book review: T. J. Fraser, The Genesis and Evolution of Time, The University of Massachusetts Press, Cambridge,1982, 205 pp. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):120, March 1983. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837304.

Yamazaki:1989:OIA

[Yam89] Masakatsu Yamazaki. Origin of the idea of the atomic bomb: Part 1: Physical research on the atomic bomb and the decision to make the bomb during World War II. Historia Scientiarum = International journal of the History of Science Society of Japan, 38(??):1–14, 1989. CODEN HISCDU. ISSN

0285-4821. URL https://ci.nii.ac.jp/naid/110006435100;
https://ci.nii.ac.jp/naid/110009837404.

Yamazaki:1994:KOH

[Yam94] Toshio Yamazaki. Kunio Oka, his life and work. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 4(1):25–38, August 1994. CODEN HISCDU. ISSN 0285-4821.

Yamazaki:1995:MST

[Yam95] Masakatsu Yamazaki. The mobilization of science and technology during the Second World War in Japan — a historical study of the activities of the Technology Board based upon the files of Tadashiro Inoue. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 5(2):167–182, November 1995. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Technology in Japan.

Yamada:2003:SRL

[Yam03] Toshihiro Yamada. Stenian revolution or Leibnizian revival?: Constructing geo-history in the Seventeenth Century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 13(2):75–100, November 2003. CODEN HIS-CDU. ISSN 0285-4821.

Yamazaki:2009:NEP

[Yam09] Masakatsu Yamazaki. Nuclear energy in postwar Japan and antinuclear movements in the 1950s. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19(2):132–145, December 2009. CODEN HISCDU. ISSN 0285-4821.

Yamaguchi:2010:SAS

[Yam10] Mari Yamaguchi. The strategy for acceptance of the scanning tunneling microscope: Observations of the Si(111) 7 × 7 reconstructed surface, 1959–1986. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(2): 123–146, December 2010. CODEN HISCDU. ISSN 0285-4821.

Yamaguchi:2014:VAA

[Yam14] Mari Yamaguchi. Visualizing the atomic arrangements in crystals: Bragg’s bubble raft model of dislocations in metals. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(3):241–260, March 2014. CODEN HIS-CDU. ISSN 0285-4821.

Yano:1980:APR

[Yan80] Michio Yano. Aryabhata’s possible rebuttal to objections to his theory of the rotation of the Earth. Historia Scientiarum = International journal of the History of Science Society of Japan, 19(??): 101–105, 1980. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110009837248.

Yan:1990:REN

[Yan90] Kangnian Yan. A re-examination into Newton’s definition of mass and Mach’s criticism. Historia Scientiarum = International journal of the History of Science Society of Japan, 40(??):29– 40, 1990. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110006435108;   https://ci.nii.ac.jp/
naid/110009837416.

Yano:1991:P

[Yan91] Michio Yano. Preface. Historia Scientiarum = International journal of the History of Science Society of Japan, 42(??):i, 1991. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837428.

Yano:2008:I

[Yan08] Michio Yano. Introduction. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 17 (3):159–160, March 2008. CODEN HISCDU. ISSN 0285-4821.

Yatsumimi:1983:BRY

[Yat83] Toshifumi Yatsumimi. Book reviews: Yōjirō Kimura, Nihon Shizenshi no Seiritsu: Rangaku to Honzōgaku (Formation of Natural History in Japan: Dutch Learning and Chinese Herbal Material Medica), Chūōkōron Sha, Tokyo, 1974, 386 pp., Yen 1,800; Yōjirō Kimura, Shiboruto to Nihon no Shokubutsu: Tōzai Bunka Kōryū no Gensen (Siebold and the Flora of Japan: Origin of the Cultural Exchange between the East and the West), Kōwa Shuppan, Tokyo, 1981, 235 pp., Yen 1,400. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):101–102, September 1983. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837446.

Ying:2013:SGD

[Yin13] Jia-Ming Ying. A survey of geometrical diagrams in Korean mathematical texts from the 17th to the 19th century. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 23(1):38–58, July 2013. CODEN HISCDU. ISSN 0285-4821.

Yoshihara:2005:OFT

[YKK05] H. Kenji Yoshihara, Teiji Kobayashi, and Masanori Kaji. Ogawa family and their “nipponium” research: Successful separation of the element 75 before its discovery by Noddacks. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 15(2):177–190, November 2005. CODEN HISCDU. ISSN 0285-4821.

Yagi:1997:TYN

[YMN97] Eri Yagi, Hisako Matsuda, and Kyomi Narita. Toshiko Yuasa (1909– 1980), and the nature of her archives at Ochanomizu Univ. in Tokyo. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):153–164, December 1997. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Yagi:2010:HET

[YO10] Eri Yagi and Tadokoro Rika Okamoto. A history of entropy through various methods: Specially focused on technical term analysis. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 20(1):47–56, July 2010. CODEN HISCDU. ISSN 0285-4821.

Yokoyama:2018:EPM

[Yok18] T. Yokoyama. On eugenic policy and the movement of the National Temperance League in prewar Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(3):353–376, 2018. CODEN HISCDU. ISSN 0285-4821.

Yoshida:1981:DDT

[Yos81a] Akira Yoshida. Davy et Dulong sur la théorie des acides au debut du XIX^{e} siècle. (French) [Davy and Dulong on the theory of acids at the beginning of the Nineteenth Century]. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):103–114, 1981. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837269.

Yoshida:1981:BRC

[Yos81b] Tadashi Yoshida. Book review: Chugoku Tenmongaku Sugaku Shu (A Collection of Chinese Astronomical and Mathematical Works), Japanese Translation with Introduction and Notes by Kiyosi Yabuuti, Keizo Hashimoto and Hideki Kawahara, Asahi-Shuppan, Tokyo, 1980, 429 pp. + xl., Yen 4,800. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??): 119–122, 1981. CODEN HISCDU. ISSN 0285-4821. URL https:/

/ci.nii.ac.jp/naid/110009837271.

Yoshioka:1981:BRC

[Yos81c] Hitoshi Yoshioka. Book review: C. Werskey, The Visible College, Allen Lane, London, 1978, 376 pp., £10. Historia Scientiarum = International journal of the History of Science Society of Japan, 20(??):125–127, 1981. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110009837259.

Yoshida:1982:BRP

[Yos82] Akira Yoshida. Book review: Pietro Redondi, L’accueil des idées de Sadi Carnot et la technologie française de 1820 à 1860 — De la légende à l’histoire, Librairie philosophique J. Vrin, Paris, 1980, 234 p., 99,00FF. Historia Scientiarum = International journal of the History of Science Society of Japan, 23(??):129–130, 1982. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837292.

Yoshida:1983:INM

[Yos83a] Fumikazu Yoshida. The industry of nations and Marx’s Das Capital. Historia Scientiarum = International journal of the History of Science Society of Japan, 24(??):77–85, March 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837299.

Yoshida:1983:RWT

[Yos83b] Fumikazu Yoshida. Robert Willis’ theory of mechanism and Karl Marx. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):87–92, September 1983. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837444.

Yoshida:1984:BRM

[Yos84a] Akira Yoshida. Book review: Michel Blay, La conceptualisation newtonienne des phénomènes de la couleur, Librairie philosophique J. Vrin, Paris, 1983, 293 p., 99F. Historia Scientiarum = International journal of the History of Science Society of Japan, 27(??):117–118, September 1984. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837324.

Yoshida:1984:NIG

[Yos84b] Akira Yoshida. Une nouvelle interprétation de la genèse des expériences de Lavoisier sur la combustion en 1772. (French) [A new interpretation of the genesis of Lavoisier’s experiments on combustion in 1772]. Historia Scientiarum = International journal of the History of Science Society of Japan, 26(??):83–94, March 1984. CODEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/

naid/110009837309.

Yoshida:1998:BRM

[Yos98] Akira Yoshida. Book review: Michel-Pierre Lerner, Le monde des sphères. (Paris: Les belles lettres) T. I, 1996, 403 pp. 215F. ISBN 2-251-42005-3; T. II, 1997, 418 pp. 190F. ISBN 2-251-42006-1. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):97–98, August 1998. CODEN HISCDU. ISSN 0285-4821.

Yoshihara:2000:ODN

[Yos00] H. Kenji Yoshihara. Ogawa’s discovery of nipponium and its reevaluation. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(3):257–??, March 2000. CODEN HISCDU. ISSN 0285-4821.

Yoshida:2001:BRK

[Yos01] Tadashi Yoshida. Book review: Kume Kunitake Monjo 2: Kagakugijutsushi Kankei (Documents by Kume Kunitake, Vol. 2: Materials relevant to the History of Science and Technology) edited by Kume Bijutsukan, texts 340 pp. + introduction 12 pp. + appendix 24 pp. + vii (index). Publisher: Yoshikawa Kobunkan (Tokyo), published in January 2000. Price: 16,000 Yen. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):281–285, March 2001. CODEN HISCDU. ISSN 0285-4821.

Young:2018:DMP

[You18] W. E. Young. Domesticating medicine: The production of familial knowledge in Nineteenth-Century Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(2):127–149, 2018. CODEN HISCDU. ISSN 0285-4821.

Yang:2006:INWa

[YR06] Jing Yang and Jean W. Richard. The influence of Norbert Wiener on Hua-Loo-Keng. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(1):96–103, July 2006. CODEN HISCDU. ISSN 0285-4821.

Yoshida:1997:ATB

[YS97] Haruyo Yoshida and Shigeo Sugiyama. Aikitu Tanakadate and the beginning of the physical researches in Japan. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(2):93–106, December 1997. CODEN HISCDU. ISSN 0285-4821. Special Issue: History of Physical Research in Modern Japan.

Yoshida:1991:GFT

[YT91] Haruyo Yoshida and Seiji Takata. The growth of Fourier’s theory of heat conduction and his experimental study. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 1(1):1–26, June 1991. CODEN HISCDU. ISSN 0285-4821. See errata [Ano91i].

Yuste:2004:TPO

[Yus04] Piedad Yuste. Trapezoidal partition in Old Babylonian geometry: a note about tablet TMS XXIII. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14 (1):67–76, July 2004. CODEN HISCDU. ISSN 0285-4821.

Yuste:2006:GTO

[Yus06] Piedad Yuste. Geometry for trapezoids in the Old Babylonian period: an observation about Tablet IM 52301. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16(2):128–143, November 2006. CODEN HISCDU. ISSN 0285-4821.

Yuste:2008:GMG

[Yus08] Piedad Yuste. Geometry in Mesopotamia and genesis of algorithms. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 18(1):24–35, July 2008. CODEN HISCDU. ISSN 0285-4821.

Yuste:2009:GOB

[Yus09] Piedad Yuste. Geometry in the Old Babylonian Period: A note on the Problem Text VAT 8393. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 19 (1):19–28, July 2009. CODEN HISCDU. ISSN 0285-4821.

Yano:1990:TCK

[YV90] Michio Yano and Merce Viladrich. Taysir computation of Kushyar ibn Labban. Historia Scientiarum = International journal of the History of Science Society of Japan, 41(??):1–16, 1990. CODEN

HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110006435110;   https://ci.nii.ac.jp/naid/110009837419.

Yamazaki:2005:PSI

[YW05] Masakatsu Yamazaki and Mark Walker. Preface to special issue “Comparative History of Nuclear Weapons Projects in Japan, Germany, and Russia in the 1940s”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 14(3):157–159, March 2005. CODEN HISCDU. ISSN 0285-4821.

Yamada:2018:SIH

[YY18] T. Yamada and M. Yajima. Special issue the history of geological sciences in east asia: Geosciences in transition. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(3):279–282, 2018. CODEN HISCDU. ISSN 0285-4821.

Zhang:2007:SUU

[Zha07] Li Zhang. From the Soviet Union to the U.S.: International exchange in China’s polymer science. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 16 (3):294–309, March 2007. CODEN HISCDU. ISSN 0285-4821. URL

https://ci.nii.ac.jp/naid/110006473350.

Zhang:2018:HMG

[Zha18] J. Zhang. How modern geology was published: A case study of Chinese geological journals, 1919 to 1949. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 27(3):283–299, 2018. CODEN HISCDU. ISSN 0285-4821.